### 29 Mar 2026 v1

# Contraherent cosheaves of contramodules on Noetherian formal schemes

Abstract.
We define the exact category of contraherent cosheaves of contramodules on a locally Noetherian formal scheme, as well as the exact categories of locally contraherent cosheaves of contramodules (with respect to a given open covering). The exposition in the section of preliminaries in adic commutative algebra is worked out in the greater generality of arbitrary commutative rings with adic topologies (of finitely generated ideals).
1. Introduction
One of the main original goals of the theory of contraherent cosheaves, developed in [17, 29] and discussed in [28], was to globalize contramodules over formal schemes and ind-schemes. In the present paper, we accomplish this goal in the modest generality of (locally) Noetherian formal schemes with adic topologies on the rings of functions on affine open formal subschemes. These are the formal schemes in the sense of Hartshorne’s textbook [12, Section II.9], as distinguished from the more general approaches of Grothendieck’s EGAI [11, Sections 0.7 and I.10] and The Stacks project [13, Chapter Tag 0AHW].
Why contraherent cosheaves of contramodules? Why not quasi-coherent sheaves of contramodules? This question was raised in [28, Section 2] and answered in [32, Section 1.10], with a definitive counterexample in [32, Remark 10.7 and Example 10.8]. Simply put, the localization functors in contramodule categories do not have the requisite exactness properties (which they have in the usual module categories). The exactness properties of the colocalization functors in the module categories are better. Another explanation is that one wants to work with infinite direct products of contramodules, and the localization functors do not commute with direct products. The main results of the paper [32] establish the existence, an explicit construction, and the exactness properties of the colocalization functors in contramodule categories in a generality far surpassing the needs of the present paper.
This paper starts with a section “Preliminaries in adic commutative algebra”, where we discuss the affine geometry of contramodules and quotseparated contramodules in the generality of commutative rings with the adic topology of a finitely generated ideal. Then follows a section where we discuss cosheaves of contramodules and define the exact category of locally contraherent cosheaves of contramodules on a locally Noetherian formal scheme $\mathfrak{X}$ with respect to its open covering $\mathbf{W}$. We also consider the dual-analogous abelian category of quasi-coherent torsion sheaves on $\mathfrak{X}$, which is well-known in the literature [1, 2].
Acknowledgement
I am grateful to Jan Št’ovíček and Michal Hrbek for helpful discussions. The author is supported by the GAČR project 26-22734S and the Institute of Mathematics, Czech Academy of Sciences (research plan RVO: 67985840).
2. Preliminaries in Adic Commutative Algebra
2.1. Adic topological rings and their maps
Let $R$ be a commutative ring. By an adic topology on $R$ we mean a ring topology for which there exists a finitely generated ideal $I\subset R$ such that the ideal $I^{n}\subset R$ is open for every $n\geq 1$ and the descending sequence $I^{n}$,  $n\geq 1$, of powers of the ideal $I$ is a base of neighborhoods of zero in $R$. If this is the case, then the topology on $R$ is said to be $I$-adic, the ring $R$ is said to be an adic topological ring, and the ideal $I$ is said to be the ideal of definition of the adic topological ring $R$.
Let $R$ be an adic topological ring and $I\subset R$ be an ideal of definition of $R$. Then, for every $n\geq 1$, the ideal $I^{n}\subset R$ is an ideal of definition, too. Another given ideal $J\subset R$ is an ideal of definition if and only if there exist integers $n$, $m\geq 1$ such that $I^{n}\subset J$ and $J^{m}\subset I$. For any two ideals of definition $I$, $J\subset R$, the ideals $I+J$,  $I\cap J$, and $IJ\subset R$ are also ideals of definition in $R$.
Examples 2.1.1.
The discrete topology on any commutative ring $R$ is adic. The zero ideal $(0)\subset R$ is an ideal of definition of $R$ in the discrete topology. The indiscrete topology on $R$ is adic as well. The unit ideal $(1)\subset R$ is an ideal of definition of $R$ in the indiscrete topology.
Lemma 2.1.2.
Let $I$ and $J$ be two ideals of definition of an adic topological ring $R$. Then the quotient ring $R/I$ is Noetherian if and only if the quotient ring $R/J$ is Noetherian.
Proof.
The assertion holds due to the assumption that the ideals $I$ and $J$ are finitely generated. If so, then the ring $R/I$ is Noetherian if and only if the ring $R/I^{n}$ is (for all, or equivalently, any given $n\geq 1$). ∎
We will say that an adic topological ring $R$ is adically Noetherian (cf. [31, Section 9]) if it satisfies the equivalent conditions of Lemma 2.1.2.
If the adic topological ring $R$ is adically Noetherian, then the nilradical of $R/I$ is a nilpotent ideal. In other words, the ideal $I$ contains a suitable power of its radical: there exists $n\geq 1$ such that $(\sqrt{I})^{n}\subset I$. Consequently, there exists a unique maximal ideal of definition $I_{\mathrm{mx}}\subset R$, characterized by the properties that $I_{\mathrm{mx}}$ is an ideal of definition of $R$ and the quotient ring $R/I_{\mathrm{mx}}$ is reduced (i. e., contains no nonzero nilpotent elements). The maximal ideal of definition $I_{\mathrm{mx}}$ can be constructed as the radical of any other ideal of definition, $I_{\mathrm{mx}}=\sqrt{I}\subset R$.
Lemma 2.1.3.
Let $R$ and $S$ be two commutative rings with adic topologies, and let $f\colon R\longrightarrow S$ be a ring homomorphism. Then the following four conditions are equivalent:
  1. (1)
    the ring homomorphism $f$ is continuous;
  2. (2)
    there exist an ideal of definition $I\subset R$ and an ideal of definition $J\subset S$ such that $f(I)\subset J$;
  3. (3)
    for every ideal of definition $J\subset S$ there exists an ideal of definition $I\subset R$ such that $f(I)\subset J$;
  4. (4)
    for every ideal of definition $I\subset R$ there exists an ideal of definition $J\subset S$ such that $f(I)\subset J$.
If the topological rings $R$ and $S$ are adically Noetherian, then conditions (1–4) are also equivalent to the following condition:
  1. (5)
    denoting by $I_{\mathrm{mx}}\subset R$ and $J_{\mathrm{mx}}\subset S$ the maximal ideals of definition of $R$ and $S$, one has $f(I_{\mathrm{mx}})\subset J_{\mathrm{mx}}$.
Proof.
(1) $\Longrightarrow$ (3) The ideal of definition $J\subset S$ is open, so there exists an open ideal $I^{\prime}\subset R$ such that $f(I^{\prime})\subset J$. Now any open ideal $I^{\prime}\subset R$ contains some ideal of definition, i. e., there exists an ideal of definition $I$ such that $I\subset I^{\prime}$.
(3) $\Longrightarrow$ (2) Obvious.
(2) $\Longrightarrow$ (1) Let $I^{\prime}\subset R$ and $J^{\prime}\subset S$ be ideals of definition for which $f(I^{\prime})\subset J^{\prime}$, and let $J\subset S$ be an open ideal. Then there exists $n\geq 1$ such that $(J^{\prime})^{n}\subset J$. Hence $f((I^{\prime})^{n})\subset(J^{\prime})^{n}\subset J$, so $(I^{\prime})^{n}\subset f^{-1}(J)$. Thus $f^{-1}(J)$ is an open ideal in $R$.
(2) $\Longrightarrow$ (4) Let $I^{\prime}\subset R$ and $J^{\prime}\subset S$ be ideals of definition for which $f(I^{\prime})\subset J^{\prime}$, and let $I\subset R$ be another given ideal of definition. Then there exists $n\geq 1$ such that $I^{n}\subset I^{\prime}$, hence $f(I^{n})\subset J^{\prime}$. Put $J=Sf(I)+J^{\prime}\subset S$, where $Sf(I)$ is the extension of the ideal $I\subset R$ in the ring $S$. Then we have $J^{\prime}\subset J$ and $J^{n}\subset J^{\prime}$, hence $J$ is an ideal of definition of $S$. By construction, $f(I)\subset J$.
(4) $\Longrightarrow$ (2) Obvious.
(2) $\Longrightarrow$ (5) We have $I_{\mathrm{mx}}=\sqrt{I}$ and $J_{\mathrm{mx}}=\sqrt{J}$. Thus $f(I)\subset J$ implies $f(I_{\mathrm{mx}})\subset J_{\mathrm{mx}}$.
(5) $\Longrightarrow$ (2) Take $I=I_{\mathrm{mx}}$ and $J=J_{\mathrm{mx}}$. ∎
Lemma 2.1.4.
Let $R$ and $S$ be two commutative rings with adic topologies, and let $f\colon R\longrightarrow S$ be a ring homomorphism. Then the following three conditions are equivalent:
  1. (1)
    there exist an ideal of definition $I\subset R$ and an ideal of definition $J\subset S$ such that $J\subset Sf(I)$;
  2. (2)
    for every ideal of definition $J\subset S$ there exist an ideal of definition $I\subset R$ and and integer $n\geq 1$ such that $J^{n}\subset Sf(I)$;
  3. (3)
    for any two ideals of definition $I\subset R$ and $J\subset S$ there exists an integer $n\geq 1$ such that $J^{n}\subset Sf(I)$;
  4. (4)
    for every ideal of definition $I\subset R$ there exists an ideal of definition $J\subset S$ such that $J\subset Sf(I)$.
Proof.
(1) $\Longrightarrow$ (2) Let $I^{\prime}\subset R$ and $J^{\prime}\subset S$ be ideals of definition for which $J^{\prime}\subset Sf(I^{\prime})$, and let $J\subset S$ be another given ideal of definition. Then there exists $n\geq 1$ such that $J^{n}\subset J^{\prime}$, hence $J^{n}\subset Sf(I^{\prime})$. It remains to put $I=I^{\prime}$.
(2) $\Longrightarrow$ (3) Let $I^{\prime}\subset R$ be an ideal of definition and $n^{\prime}\geq 1$ be an integer such that $J^{n^{\prime}}\subset Sf(I^{\prime})$. Pick an integer $m\geq 1$ such that $(I^{\prime})^{m}\subset I$. Then $J^{n^{\prime}m}\subset Sf((I^{\prime})^{m})\subset Sf(I)$. It remains to put $n=n^{\prime}m$.
(3) $\Longrightarrow$ (4) $\Longrightarrow$ (1) Obvious. ∎
Remark 2.1.5.
Notice that the equivalent conditions of Lemma 2.1.4 do not imply that for every ideal of definition $J\subset S$ there exists an ideal of definition $I\subset R$ such that $J\subset Sf(I)$. For a counterexample, it suffices to take $R=k$ to be a field and $S=k[\epsilon]/(\epsilon^{2})$ to be the $k$-algebra of dual numbers. Endow both $R$ and $S$ with the discrete topologies, and let $f\colon R\longrightarrow S$ be the identity inclusion. Then $I=0$ is the only ideal of definition in $R$, and $J=(\epsilon)$ is an ideal of definition in $S$, but $J\not\subset Sf(I)$. On the other hand, $J^{\prime}=(0)$ is also an ideal of definition in $S$, and $J^{\prime}=Sf(I)$.
We will say that a ring homomorphism $f\colon R\longrightarrow S$ acting between two rings with adic topologies is a tight ring map if the equivalent conditions of Lemma 2.1.4 hold. Notice that a tight ring map need not be continuous in our definition. Using the criterion of Lemma 2.1.4(4), one can easily check that the composition of any two tight ring maps of adic topological rings is tight.
Given an $I$-adic topological ring $R$ and an $R$-module $M$, the $I$-adic topology on $M$ is defined by the property that the submodules $I^{n}M\subset M$,  $n\geq 1$, form a base of open neighborhoods of zero. Clearly, if $I^{\prime}$ and $I^{\prime\prime}$ are two ideals of definition of $R$, then the $I^{\prime}$-adic and the $I^{\prime\prime}$-adic topologies on $M$ coincide.
Lemma 2.1.6.
Let $R$ and $S$ be two commutative rings with adic topologies, and let $f\colon R\longrightarrow S$ be a ring homomorphism. Then the following five conditions are equivalent:
  1. (1)
    the ring homomorphism $f$ is continuous and tight;
  2. (2)
    there exists an ideal of definition $I\subset R$ such that the topology on $S$ is the $I$-adic topology;
  3. (3)
    for every ideal of definition $I\subset R$, the topology on $S$ is the $I$-adic topology;
  4. (4)
    there exists an ideal of definition $I\subset R$ such that $J=Sf(I)$ is an ideal of definition in $S$;
  5. (5)
    for every ideal of definition $I\subset R$, the ideal $J=Sf(I)\subset S$ is an ideal of definition in $S$.
Proof.
(2) $\Longleftrightarrow$ (3) Follows from the discussion in the paragraph preceding the lemma.
(1) $\Longrightarrow$ (3) Since $f$ is continuous, Lemma 2.1.3(3) says that for every ideal of definition $J\subset S$ there exists an ideal of definition $I\subset R$ such that $Sf(I)\subset J$. Since $f$ is tight, Lemma 2.1.4(4) says that for every ideal of definition $I\subset R$ there exists an ideal of definition $J\subset S$ such that $J\subset Sf(I)$. Thus the $J$-adic and the $I$-adic topologies on $S$ coincide.
(3) $\Longrightarrow$ (1) Under (3), both the topologies on $R$ and $S$ are the $I$-adic topologies. As every $R$-module map $M\longrightarrow N$ is continuous with respect to the $I$-adic topologies on $M$ and $N$, it follows that $f$ is continuous.
To check that $f$ is tight, let $I\subset R$ and $J\subset S$ be any two ideals of definition. In the $I$-adic topology, $Sf(I)$ is an open subgroup in $S$. Thus (2) implies that there exists $n\geq 1$ for which $J^{n}\subset Sf(I)$. So the condition of Lemma 2.1.4(3) is satisfied.
(1) $\Longrightarrow$ (5) By Lemma 2.1.4(4), there exists an ideal of definition $J^{\prime}\subset S$ such that $J^{\prime}\subset Sf(I)$. By Lemma 2.1.3(3), there exists an ideal of definition $I^{\prime}\subset R$ such that $Sf(I^{\prime})\subset J^{\prime}$. Pick $n\geq 1$ such that $I^{n}\subset I^{\prime}$; then we have $(Sf(I))^{n}=Sf(I^{n})\subset J^{\prime}\subset Sf(I)$. Since $J^{\prime}$ is an ideal of defintion in $S$, it follows that $Sf(I)$ is an ideal of definition in $S$, too.
(5) $\Longrightarrow$ (4) $\Longrightarrow$ (2) Obvious. ∎
It is clear from Lemma 2.1.6 that, for any homomorphism of commutative rings $f\colon R\longrightarrow S$ and any adic topology on the ring $R$, there exists a unique adic topology on the ring $S$ making the ring map $f$ continuous and tight. It also follows that any tight continuous ring map that is an isomorphism of abstract commutative rings is an isomorphism of topological rings.
Examples 2.1.7.
Let $k$ be a field and $R=S=k[x]$ be the ring of polynomials in one variable $x$ over $k$. Let $f\colon R\longrightarrow S$ be the identity map. Consider the ideals $I=(x)\subset R$, $J=(x)\subset S$, $I_{0}=(0)\subset R$, and $J_{0}=(0)\subset S$. Then the map $f$ is continuous and tight with respect to the $I$-adic topology on $R$ and the $J$-adic topology on $S$. The map $f$ is also continuous and tight with respect to the $I_{0}$-adic topology on $R$ and the $J_{0}$-adic topology on $S$. With respect to the $I_{0}$-adic topology on $R$ and the $J$-adic topology on $S$, the map $f$ is continuous but not tight. With respect to the $I$-adic topology on $R$ and the $J_{0}$-adic topology on $S$, the map $f$ is tight but not continuous (in our terminology).
To give another example, let $R=k[x]$ be the ring of polynomials in one variable and $S=k[x,y]$ be the ring of polynomials in two variables. Consider the ideals $I=(x)\subset R$ and $J=(x,y)\subset S$. Let $g\colon R\longrightarrow S$ be the identity inclusion map and $h\colon S\longrightarrow R$ be the $k$-algebra homomorphism taking $x$ to $x$ and $y$ to 0. Then the map $g$ is continuous but not tight, while the map $h$ is continuous and tight. For a perhaps more complicated example of a tight continuous map, see Remark 2.1.5 (in fact, any morphism of discrete commutative rings is continuous and tight).
Lemma 2.1.8.
Let $R$ and $S$ be two commutative rings with adic topologies, and let $f\colon R\longrightarrow S$ be a tight continuous ring homomorphism. Then the following two conditions are equivalent:
  1. (1)
    there exists an ideal of definition $I\subset R$ such that the $R/I^{n}$-module $S/Sf(I^{n})$ is flat for every $n\geq 1$;
  2. (2)
    for every open ideal $I\subset R$, the $R/I$-module $S/Sf(I)$ is flat.
Proof.
The point is that if $I\subset I^{\prime}\subset R$ are two ideals in $R$ such that the $R/I$-module $S/Sf(I)$ is flat, then the $R/I^{\prime}$-module $S/Sf(I^{\prime})=R/I^{\prime}\otimes_{R/I}S/Sf(I)$ is flat, too. ∎
We will say that a tight continuous homomorphism of adic topological rings $f\colon R\longrightarrow S$ is flat if it satisfies the equivalent conditions of Lemma 2.1.8. Clearly, if a tight continuous map of adic topological rings $f$ is flat as a map of abstract rings (i. e., the $R$-module $S$ is flat), then $f$ is also flat as as a tight continuous map of adic topological rings. But the converse is not true in general.
Using the criterion of Lemma 2.1.8(2), one can easily check that the composition of any two flat tight continuous ring maps of adic topological rings is flat.
2.2. Torsion modules
Let $R$ be a commutative ring and $s\in R$ be an element. Following the notation in [17, Section 1.1], [20, Section 2.1], etc., we denote by $R[s^{-1}]$ the ring $R$ with the element $s$ formally inverted. In other words, $R[s^{-1}]=S^{-1}R$ is the localization of $R$ at the multiplicative subset $S=\{1,s,s^{2},s^{3},\dotsc\}\subset R$.
Let $M$ be an $R$-module. An element $x\in M$ is said to be $s$-torsion if there exists $n\geq 1$ such that $s^{n}x=0$ in $M$. The $R$-module $M$ is said to be $s$-torsion if all the elements of $M$ are $s$-torsion. Equivalently, an $R$-module $M$ is $s$-torsion if and only if $R[s^{-1}]\otimes_{R}M=0$.
Let $I\subset R$ be an ideal. We say that an $R$-module $\mathcal{M}$ is $I$-torsion if $\mathcal{M}$ is $s$-torsion for all $s\in R$. One can easily see that it suffices to check this condition for the element $s$ ranging over any chosen set of generators of the ideal II[30, Lemma 1.2].
Other definitions of the term “$I$-torsion $R$-module” exist in the literature (see, e. g., [14, Section 3]), but the difference between them does not manifest itself for finitely generated ideals $I$, which we are really interested in. So, for a finitely generated ideal $I\subset R$, an $R$-module $\mathcal{M}$ is $I$-torsion if and only if, for every element $x\in\mathcal{M}$, there exists an integer $n\geq 1$ such that $I^{n}x=0$ in $\mathcal{M}$[30, Lemma 1.3].
For any $R$-module $M$ and any ideal $J\subset R$, we denote by $_{J}M\subset M$ the submodule of all elements annihilated by $J$ in $M$. So $_{J}M\simeq\operatorname{Hom}_{R}(R/J,M)$. According to the previous paragraph, an $R$-module $\mathcal{M}$ is $I$-torsion if and only if $\mathcal{M}=\bigcup_{n\geq 1}{}_{I^{n}}\mathcal{M}$.
Lemma 2.2.1.
Let $R$ be a commutative ring, $I\subset R$ be an ideal, $\mathcal{M}$ be an $I$-torsion $R$-module, and $N$ be an $R$-module. Then the $R$-module $\mathcal{M}\otimes_{R}N$ is $I$-torsion.
Proof.
The assertions is obvious; see [20, Lemma 6.1(a)] for some details. ∎
Now let $R$ be a commutative ring endowed with an adic topology in the sense of Section 2.1. Let $I$, $J\subset R$ be two ideals of definition of the ring $R$. Then an $R$-module $\mathcal{M}$ is $I$-torsion if and only if it is $J$-torsion. If this is the case, we say that $\mathcal{M}$ is a torsion $R$-module (with respect to the given adic topology of $R$).
An $R$-module $\mathcal{M}$ is torsion if and only if, for every element $x\in\mathcal{M}$, the annihilator of $x$ in $R$ is an open ideal. In other words, an $R$-module $\mathcal{M}$ is torsion if and only if it is discrete (as a module over the topological ring $R$) in the sense of [39, Section VI.4], [3, Section 1.4], [15, Section D.5.1], [17, Section E.2], [18, Sections 1.4 and 2.1], [22, Section 2.3], [23, Section 2.4], [35, Section 7.2], [36, Section 1], etc.
For any ring $R$, we denote by $R{\operatorname{\mathsf{--Mod}}}$ the abelian category of (left) $R$-modules. Given an ideal $I$ in a commutative ring $R$, the notation $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}\subset R{\operatorname{\mathsf{--Mod}}}$ stands for the full subcategory of $I$-torsion $R$-modules. For an adic topological ring $R$, we will write simply $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$ instead of $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}$, presuming any ideal of definition $I$ of the topological ring $R$. We will also write $R{\operatorname{\mathsf{--Tors}}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$.
For any ideal $I$ in a commutative ring $R$, the full subcategory $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}$ is closed under submodules, quotients, extensions, and infinite direct sums in $R{\operatorname{\mathsf{--Mod}}}$. In other words, $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}$ is a localizing subcategory, or in a different terminology, a hereditary torsion class [39, Sections VI.2–3] in $R{\operatorname{\mathsf{--Mod}}}$.
It follows that the category $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}$ is a Grothendieck abelian category. The identity inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is exact and preserves infinite direct sums (hence also all colimits), but not infinite products. The infinite direct product functors in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}$ are usually not exact.
2.3. Adic completions
Let $R$ be a commutative ring and $I\subset R$ be a finitely generated ideal. Then the construction of the $I$-adic completion
[…]
There is a natural completion map $\lambda_{I,M}\colon M\longrightarrow\Lambda_{I}(M)$. We say that an $R$-module $M$ is $I$-adically separated if the map $\lambda_{I,M}$ is injective, and that $M$ is $I$-adically complete if $\lambda_{I,M}$ is surjective. For any $R$-module $M$, the $R$-module $\Lambda_{I}(M)$ is $I$-adically separated and complete [40, Corollary 3.6]. The latter assertion depends on the assumption that the ideal $I\subset R$ is finitely generated [40, Example 1.8]. The functor $\Lambda_{I}$ is left adjoint to the inclusion of the full subcategory of $I$-adically separated and complete $R$-modules into $R{\operatorname{\mathsf{--Mod}}}$, and the map $\lambda_{I,M}$ is the adjunction unit [20, Theorem 5.8].
[…]
Given an adic topological ring $R$, the completion functor $\Lambda_{I}\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}$ and the natural map $\lambda_{I,M}$ depend on the adic topology on the ring $R$ only, and not on the choice of a particular ideal of definition $I\subset R$. Consequently, the properties of an $R$-module to be $I$-adically separated or $I$-adically complete are also determined by the $I$-adic topology on $R$ and do not depend on the choice of an ideal of definition $I$ for the given adic topology. So we will write $\Lambda=\Lambda_{I}$ and $\lambda_{M}=\lambda_{I,M}$, presuming any ideal of definition $I$ of the topological ring $R$.
Following the terminology above, an adic topological ring $R$ is said to be separated if the map $\lambda_{R}\colon R\longrightarrow\mathfrak{R}$ is injective and complete if the map $\lambda_{R}$ is surjective. For any adic topological ring $R$, the adic topological ring $\mathfrak{R}=\Lambda(R)$ is separated and complete, and the completion homomorphism of topological rings $\lambda_{R}\colon R\longrightarrow\mathfrak{R}$ is continuous, tight, and flat (in the sense of Section 2.1).
Let $R$ be an adic topological ring and $\mathfrak{R}=\Lambda(R)$ be its completion. Then the rules $\mathfrak{I}=\mathfrak{R}\lambda_{R}(I)$ and $I=\lambda_{R}^{-1}(\mathfrak{I})$ establish a natural bijection between the ideals of definition $I\subset R$ and the ideals of definition $\mathfrak{I}\subset\mathfrak{R}$. More generally, the same rules provide a bijection between open ideals $I\subset R$ and open ideals $\mathfrak{I}\subset\mathfrak{R}$. The map $\lambda_{R}$ induces ring isomorphisms $R/I\simeq\mathfrak{R}/\mathfrak{I}$ for open ideals $I\subset R$ and $\mathfrak{I}\subset\mathfrak{R}$ corresponding to each other under this bijection.
Any continuous homomorphism of adic topological rings $f\colon R\longrightarrow S$ induces a continuous homomorphism of the completions $\Lambda(f)\colon\Lambda(R)\longrightarrow\Lambda(S)$.
Lemma 2.3.1.
Let $f\colon R\longrightarrow S$ be a tight continuous homomorphism of adic topological rings. Then
(a) the induced continuous ring map of the completions $\Lambda(f)\colon\Lambda(R)\longrightarrow\Lambda(S)$ is also tight;
(b) the tight continuous ring map $f$ is flat if and only if the tight continuous ring map $\Lambda(f)$ is flat.
Proof.
Put $\mathfrak{R}=\Lambda(R)$,  $\mathfrak{S}=\Lambda(S)$, and $\mathfrak{f}=\Lambda(f)$.
Part (a): let $I\subset R$ be an ideal of definition in $R$. Then, by Lemma 2.1.6(5), the ideal $J=Sf(I)\subset S$ is an ideal of definition in $S$. Following the discussion above, the ideal $\mathfrak{I}=\mathfrak{R}\lambda_{R}(I)\subset\mathfrak{R}$ is an ideal of definition in $\mathfrak{R}$ and the ideal $\mathfrak{J}=\mathfrak{S}\lambda_{S}(J)$ is an ideal of definition in $\mathfrak{S}$. Clearly, we have $\mathfrak{J}=\mathfrak{S}\mathfrak{f}(\mathfrak{I})$. By Lemma 2.1.6(4), it follows that the ring map $\mathfrak{f}$ is tight.
Part (b): in the notation of the previous paragraph, we have $R/I\simeq\mathfrak{R}/\mathfrak{I}$ and $S/J\simeq\mathfrak{S}/\mathfrak{J}$. Hence $S/J$ is a flat module over $R/I$ if and only if $\mathfrak{S}/\mathfrak{J}$ is a flat module over $\mathfrak{R}/\mathfrak{I}$. ∎
Remark 2.3.2.
The converse assertion to Lemma 2.3.1(a) is not true: There exist continuous maps of adic topological rings $f\colon R\longrightarrow S$ such that the continuous ring map $\Lambda(f)$ is tight but the map $f$ is not. For a counterexample, it suffices to take $R=S=k[x]$ and $f=\mathrm{id}_{k[x]}$ as in Examples 2.1.7, with the discrete ($(0)$-adic) topology on $R$ and the indiscrete ($(1)$-adic) topology on $S$. Then one has $\Lambda(R)=R$ and $\Lambda(S)=0$; so the ring map $\mathfrak{f}=\Lambda(f)$ is tight; but the ring map $f$ is not.
Let $R$ be an adic topological ring and $\mathfrak{R}=\Lambda(R)$ be its completion. The adic completion functor $\Lambda\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ or $\Lambda\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}$ is neither left, nor right exact [40, Example 3.20]. This fact is closely related to the fact that the category of separated and complete $R$-modules is not abelian [20, Example 2.7(1)]. To overcome these difficulties, one considers contramodule $R$-modules instead of the separated and complete $R$-modules; see the next section.
2.4. Contramodules
Let $R$ be a commutative ring and $s\in R$ be an element. An $R$-module $P$ is said to be an $s$-contramodule if $\operatorname{Hom}_{R}(R[s^{-1}],P)=0=\operatorname{Ext}_{R}^{1}(R[s^{-1}],P)$. The ring $R[s^{-1}]$ is considered as an $R$-module here. The definition above does not mention any $\operatorname{Ext}^{i}$ with $i\geq 2$ because the projective dimension of the $R$-module $R[s^{-1}]$ never exceeds 11[17, Section 1.1], [16, proof of Lemma B.7.1], [20, proof of Lemma 2.1]. A detailed discussion of this definition, which goes back to [15, Remark A.1.1], [16, Theorem B.1.1], [17, Section D.1], and [19, Section 2], can be found in [20, Sections 1–3].
Let $I\subset R$ be an ideal. An $R$-module $\mathfrak{P}$ is said to be an $I$-contramodule (or an $I$-contramodule $R$-module) if $\mathfrak{P}$ is an $s$-contramodule for every $s\in I$. It suffices to check this condition for the element $s$ ranging over any chosen set of generators of the ideal II[16, Theorem B.1.1], [19, Section 2]. A detailed discussion with two proofs can be found in [20, Theorem 5.1].
Lemma 2.4.1.
Let $R$ be a commutative ring and $I\subset R$ be an ideal. Then
(a) for any $I$-torsion $R$-module $\mathcal{M}$ and any $R$-module $P$, the $R$-module $\operatorname{Hom}_{R}(\mathcal{M},P)$ is an $I$-contramodule;
(b) for any $R$-module $M$ and any $I$-contramodule $R$-module $\mathfrak{P}$, the $R$-module $\operatorname{Hom}_{R}(M,\mathfrak{P})$ is an $I$-contramodule.
Proof.
This is [20, Lemma 6.1(b)]. ∎
Assume that the ideal $I\subset R$ is finitely generated. Then any $I$-contramodule $R$-module is $I$-adically complete [20, Theorem 5.6], but it need not be $I$-adically separated [37, Example 2.5], [40, Example 3.20], [14, Example 4.33], [15, Section A.1.1], [18, Section 1.5], [20, Example 2.7(1)]. Any $I$-adically separated and complete $R$-module is an $I$-contramodule [20, Lemma 5.7].
An $I$-contramodule $R$-module is said to be quotseparated [34, Section 5.5], [24, Section 1] if it is a quotient $R$-module of an $I$-adically separated and complete $R$-module. The references in the previous paragraph provide examples of quotseparated $I$-contramodule $R$-modules that are not $I$-adically separated. For an example of an $I$-contramodule $R$-module that is not quotseparated, see [19, Example 2.6], [24, Examples 1.8]. Over a Noetherian commutative ring $R$, all $I$-contramodule $R$-modules are quotseparated [16, Theorem B.1.1], [24, Corollary 3.7 and Remark 3.8]. A perhaps more detailed discussion can be found in [31, Section 1].
Now let $R$ be a commutative ring endowed with an adic topology in the sense of Section 2.1. Let $I$, $J\subset R$ be two ideals of definition of the ring $R$. Then an $R$-module $\mathfrak{P}$ is an $I$-contramodule if and only if it is a $J$-contramodule [20, Remark 5.5]. If this is the case, we will say that $\mathfrak{P}$ is a contramodule $R$-module (with respect to the given adic topology of $R$). Moreover, a contramodule $R$-module is quotseparated as an $I$-contramodule $R$-module if and only if it is quotseparated as a $J$-contramodule $R$-module; so one can speak of quotseparated contramodule $R$-modules.
Given an ideal $I$ in a commutative ring $R$, we denote by $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}\subset R{\operatorname{\mathsf{--Mod}}}$ the full subcategory of $I$-contramodule $R$-modules. When the ideal $I\subset R$ is finitely generated, the notation $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}\subset R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ stands for the full subcategory of quotseparated $I$-contramodule $R$-modules. For an adic topological ring $R$, we will write simply $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\subset R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ instead of $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}\subset R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$, presuming any ideal of definition $I$ of the topological ring $R$.
For any ideal $I$ in a commutative ring $R$, the full subcategory $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ is closed under kernels, cokernels, extensions, and infinite direct products in $R{\operatorname{\mathsf{--Mod}}}$[7, Proposition 1.1], [19, Section 2], [20, Theorem 1.2(a)]. It follows that $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ is an abelian category. In fact, $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ is a locally $\aleph_{1}$-presentable abelian category with enough projective objects [33, Example 4.1(3)] (see also [21, Example 1.3(4)]). The identity inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is exact and preserves infinite products (hence also all limits) and $\aleph_{1}$-directed colimits, but not infinite direct sums. The infinite direct sum functors in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ are usually not exact.
In the case of a finitely generated ideal $I\subset R$, the full subcategory $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}$ is closed under kernels, cokernels, and infinite direct products in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ and in $R{\operatorname{\mathsf{--Mod}}}$, and also under subobjects and quotient objects in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$[24, Lemma 1.2]. (However, the full subcategory of quotseparated contramodules is not closed under extensions; in fact, any $I$-contramodule $R$-module is an extension of two quotseparated $I$-contramodule $R$-modules [24, Proposition 1.6].) It follows that $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}$ is an abelian category. In fact, $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ is also a locally $\aleph_{1}$-presentable abelian category with enough projective objects [24, end of Section 1]. The identity inclusion functors $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ and $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ are exact and preserve infinite products (hence also all limits) and $\aleph_{1}$-directed colimits. The inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ does not preserve infinite direct sums. The infinite direct sum functors in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}$ are usually not exact (the global dimension $1$ case is one major exception [16, Remark 1.2.1]).
To any complete, separated topological ring $\mathfrak{R}$ where open right ideals form a base of neighborhoods of zero, the abelian category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ of left $\mathfrak{R}$-contramodules is naturally assigned [15, Remark A.3], [16, Section 1.2], [18, Section 2.1], [33, Sections 1.2 and 5], [22, Section 2.7], [23, Sections 2.6–2.7], [35, Section 6], [36, Section 1]. The category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ comes equipped with a natural exact, faithful forgetful functor $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}$, preserving infinite products (hence all limits).
In the case of an adic topological ring $R$, the composition of forgetful functors $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is fully faithful, and its essential image coincides with the full subcategory of quotseparated contramodule $R$-modules [24, Proposition 1.5]. So there is a natural equivalence (in fact, isomorphism) of categories $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\simeq R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$.
Lemma 2.4.2.
Let $R$ be an adic topological ring. Then
(a) for any torsion $R$-module $\mathcal{M}$ and any $R$-module $P$, the $R$-module $\operatorname{Hom}_{R}(\mathcal{M},P)$ is a quotseparated contramodule $R$-module (in fact, a separated and complete $R$-module);
(b) for any $R$-module $M$ and any complete, separated $R$-module $\mathfrak{P}$, the $R$-module $\operatorname{Hom}_{R}(M,\mathfrak{P})$ is complete and separated;
(c) for any $R$-module $M$ and any quotseparated contramodule $R$-module $\mathfrak{P}$, the $R$-module $\operatorname{Hom}_{R}(M,\mathfrak{P})$ is a quotseparated contramodule $R$-module.
Proof.
Part (a) is the assertion in parentheses in [31, Lemma 1.1(b)]. To prove parts (b) and (c), represent $M$ as a quotient $R$-module of a free $R$-module $F$. Then $\operatorname{Hom}_{R}(M,\mathfrak{P})$ is an $R$-submodule of the $R$-module $\operatorname{Hom}_{R}(F,\mathfrak{P})$, which an infinite product of copies of $\mathfrak{P}$. Furthermore, $\operatorname{Hom}_{R}(M,\mathfrak{P})$ is a contramodule $R$-module by Lemma 2.4.1(b). Now assertion (b) follows from the fact that the full subcategory of separated $R$-modules is closed under subobjects and infinite products in $R{\operatorname{\mathsf{--Mod}}}$. Since the full subcategory $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ is closed under subobjects, quotient objects, and infinite products in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$, part (c) follows as well. For an argument based on the notion of a contramodule over a topological ring, see [16, Section B.2]. ∎
2.5. Injective torsion modules, projective and flat contramodules
For any abelian category $\mathsf{A}$, we denote by $\mathsf{A}^{\mathsf{inj}}\subset\mathsf{A}$ the full subcategory of injective objects in $\mathsf{A}$. More generally, the same notation applies to any exact category $\mathsf{A}$ (in the sense of Quillen). Dually, the notation $\mathsf{B}_{\mathsf{prj}}\subset\mathsf{B}$ stands for the full subcategory of projective objects in an abelian or exact category $\mathsf{B}$.
Let $I$ be an ideal in a commutative ring $R$. Given an $R$-module $M$, we denote by $\Gamma_{I}(M)\subset M$ the submodule of all $I$-torsion elements in $M$. The functor $M\longmapsto\Gamma_{I}(M)\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}$ is right adjoint to the identity inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$.
As any Grothendieck category, the abelian category $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}$ has enough injective objects. For every injective $R$-module $K$, the torsion $R$-module $\Gamma_{I}(K)$ is injective as an object of $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}$. Indeed, the maximal torsion submodule functor $\Gamma_{I}$ is right adjoint to an exact functor, so it takes injective objects to injective objects. There are enough injective objects of the form $\Gamma_{I}(K)$ in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}$, where $K\in R{\operatorname{\mathsf{--Mod}}}^{\mathsf{inj}}$. Hence an $I$-torsion $R$-module $\mathcal{K}$ is injective in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}$ if and only if it is a direct summand of the $I$-torsion $R$-module $\Gamma_{I}(K)$ for some injective $R$-module $K$.
Let $R$ be an adic topological ring with an ideal of definition $I\subset R$. Clearly, the submodule $\Gamma(M)=\Gamma_{I}(M)\subset M$ only depends on the adic topology on $R$, and not on the choice of a particular ideal of definition $I$. Using Baer’s criterion of injectivity, one can easily check that a torsion $R$-module $\mathcal{K}$ is an injective object of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$ if and only if the $R/I^{n}$-module $_{I^{n}}\mathcal{K}$ is injective for every $n\geq 1$. Equivalently, a torsion $R$-module $\mathcal{K}$ is injective (as a torsion $R$-module) if and only if the $R/J$-module $_{J}\mathcal{K}$ is injective for every ideal of definition $J\subset R$. If this is the case, then the $R/J$-module $_{J}\mathcal{K}$ is also injective for every open ideal $J\subset R$.
When the ring $R$ is Noetherian (as an abstract ring), one can easily prove using the Artin–Rees lemma that any injective object of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$ is also injective in $R{\operatorname{\mathsf{--Mod}}}$. Thus one has $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}^{\mathsf{inj}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\cap R{\operatorname{\mathsf{--Mod}}}^{\mathsf{inj}}$ in this case. The following lemma summarizes the observations above.
Lemma 2.5.1.
Let $R$ be a Noetherian commutative ring endowed with an adic topology, let $I\subset R$ be an ideal of definition of the topological ring $R$, and let $\mathcal{K}$ be a torsion $R$-module. Then the following six conditions are equivalent:
  1. (1)
    $\mathcal{K}$ is an injective object of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$;
  2. (2)
    there exists an injective $R$-module $K$ such that $\mathcal{K}$ is a direct summand of the torsion $R$-module $\Gamma_{I}(K)$;
  3. (3)
    for every $n\geq 1$, the $R/I^{n}$-module $_{I^{n}}\mathcal{K}$ is injective;
  4. (4)
    for every ideal of definition $J\subset R$, the $R/J$-module $_{J}\mathcal{K}$ is injective;
  5. (5)
    for every open ideal $J\subset R$, the $R/J$-module $_{J}\mathcal{K}$ is injective;
  6. (6)
    $\mathcal{K}$ is injective as an object of $R{\operatorname{\mathsf{--Mod}}}$.
Corollary 2.5.2.
Let $R$ be a Noetherian commutative ring endowed with an adic topology. Then the fully faithful inclusion of abelian categories $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ induces isomorphisms on all the Ext groups/modules.
Proof.
The assertion follows from Lemma 2.5.1 (1) $\Leftrightarrow$ (6). For a stronger result concerning the full-and-faithfulness of the induced functor between the unbounded derived categories, see [19, Theorem 1.3]. The Noetherianity assumption can be weakened, but it cannot be simply dropped; see [24, Theorem 4.1]. ∎
For any ideal $I$ in a commutative ring $R$, the identity inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}\allowbreak\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ has a left adjoint functor $\Delta_{I}\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$[33, Examples 4.1(2–3)], [21, Example 1.3(4)]. An explicit construction of the functor $\Delta_{I}$ in the case of a finitely generated ideal $I\subset R$ can be found in [19, Proposition 2.1]; see [20, Sections 6–7] for a more detailed discussion.
For every projective $R$-module $P$, the $I$-contramodule $R$-module $\Delta_{I}(P)$ is projective as an object of $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$. Indeed, the functor $\Delta_{I}$ is left adjoint to an exact functor, so it takes projective objects to projective objects. The $R$-modules $\Delta_{I}(F)$, where $F$ ranges over the free $R$-modules, are called the free $I$-contramodule $R$-modules. There are enough free $I$-contramodule $R$-modules in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$. Hence an $I$-contramodule $R$-module $\mathfrak{P}$ is projective in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ if and only if it is a direct summand of some free $I$-contramodule $R$-module.
Let $R$ be an adic topological ring. Then the full subcategory $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}\subset R{\operatorname{\mathsf{--Mod}}}$ does not depend on the choice of an ideal of definition $I\subset R$ of the topological ring $R$. So we will write $\Delta=\Delta_{I}$, and speak of free contramodule $R$-modules presuming any ideal of definition $I$ of the adic topology on $R$.
The 0-th left derived functor $\mathbb{L}_{0}\Lambda$ of the (neither left nor right exact) adic completion functor $\Lambda$ is the left adjoint functor $\mathbb{L}_{0}\Lambda\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ to the identity inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$[24, Proposition 1.3]. Similarly to the discussion above, for every projective $R$-module $P$, the quotseparated contramodule $R$-module $\mathbb{L}_{0}\Lambda(P)$ is projective as an object of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$. Notice that, by the definition of the 0-th derived functor, one has $\mathbb{L}_{0}\Lambda(P)=\Lambda(P)$ for all projective $R$-modules $P$.
The $R$-modules $\Lambda(F)=\mathbb{L}_{0}\Lambda(F)$, where $F$ ranges over the free $R$-modules, are called the free quotseparated contramodule $R$-modules. There are enough free quotseparated contramodule $R$-modules in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$. Hence quotseparated contramodule $R$-module $\mathfrak{P}$ is projective in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ if and only if it is a direct summand of some free quotseparated contramodule $R$-module.
[…]
For any complete, separated topological ring $\mathfrak{R}$ where open right ideals form a base of neighborhoods of zero, and for any set $X$, the set/abelian group $\mathfrak{R}[[X]]$ has a natural structure of left $\mathfrak{R}$-contramodule, called the free $\mathfrak{R}$-contramodule spanned by $X$. The free $\mathfrak{R}$-contramodules are projective objects in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$, there are enough of them, and all projective $\mathfrak{R}$-contramodules are direct summands of free ones.
[…]
When the ring $R$ is Noetherian (as an abstract ring), one has $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ (see Section 2.4). Hence a natural isomorphism $\Delta\simeq\mathbb{L}_{0}\Lambda$ of functors $R{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$.
Lemma 2.5.3.
Let $R$ be an adic topological ring, let $I\subset R$ be an open ideal, and let $P$ be an $R$-module. Then the natural adjunction maps $P\longrightarrow\Delta(P)\longrightarrow\mathbb{L}_{0}\Lambda(P)\longrightarrow\Lambda(P)$ induce isomorphisms of $R/I$-modules $R/I\otimes_{R}P\simeq R/I\otimes_{R}\Delta(P)\simeq R/I\otimes_{R}\mathbb{L}_{0}\Lambda(P)\simeq R/I\otimes_{R}\Lambda(P)$.
Proof.
Follows from the facts that the functor $P\longmapsto R/I\otimes_{R}P$ is left adjoint to the inclusion of full subcategory $R/I{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$, and that all $R/I$-modules are separated and complete $R$-modules. ∎
Corollary 2.5.4.
Let $R$ be an adic topological ring, let $M$ be a torsion $R$-module, and let $P$ be an $R$-module. Then the natural adjunction maps $P\longrightarrow\Delta(P)\longrightarrow\mathbb{L}_{0}\Lambda(P)\longrightarrow\Lambda(P)$ induce isomorphisms of $R$-modules $M\otimes_{R}P\simeq M\otimes_{R}\Delta(P)\simeq M\otimes_{R}\mathbb{L}_{0}\Lambda(P)\simeq M\otimes_{R}\Lambda(P)$.
Proof.
Any torsion $R$-module $M$ is a direct union of $R/I$-modules, where $I$ ranges over the ideals of definition of $R$. Since the tensor product functor $\otimes_{R}$ preserves direct limits, it suffices to consider the case of an $R/I$-module $M$. In that case, the assertion follows from Lemma 2.5.3 in view of the natural isomorphism $M\otimes_{R}Q\simeq M\otimes_{R/I}(R/I\otimes_{R}Q)$ for any $R$-module $Q$. ∎
Lemma 2.5.5.
Let $R$ be a Noetherian commutative ring endowed with an adic topology, let $I\subset R$ be an ideal of definition of the topological ring $R$, and let $\mathfrak{P}$ be a contramodule $R$-module. Then the following seven conditions are equivalent:
  1. (1)
    $\mathfrak{P}$ is a projective object of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$;
  2. (2)
    there exists a free $R$-module $F$ such that $\mathfrak{P}$ is a direct summand of the contramodule $R$-module $\Delta(F)=\mathbb{L}_{0}\Lambda(F)=\Lambda(F)$;
  3. (3)
    for every $n\geq 1$, the $R/I^{n}$-module $\mathfrak{P}/I^{n}\mathfrak{P}$ is projective;
  4. (4)
    for every ideal of definition $J\subset R$, the $R/J$-module $\mathfrak{P}/J\mathfrak{P}$ is projective;
  5. (5)
    for every open ideal $J\subset R$, the $R/J$-module $\mathfrak{P}/J\mathfrak{P}$ is projective;
  6. (6)
    for every $n\geq 1$, the $R/I^{n}$-module $\mathfrak{P}/I^{n}\mathfrak{P}$ is flat, and the $R/I$-module $\mathfrak{P}/I\mathfrak{P}$ is projective;
  7. (7)
    $\mathfrak{P}$ is a flat $R$-module and the $R/I$-module $\mathfrak{P}/I\mathfrak{P}$ is projective.
Proof.
(1) $\Longleftrightarrow$ (2) Explained above. No Noetherianity assumption on $R$ is needed here. (In the context of $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ and $\Delta_{I}$, the assertion even holds for infinitely generated ideals $I\subset R$.)
(3) $\Longleftrightarrow$ (4) $\Longleftrightarrow$ (5) This is obvious and holds for any adic topological ring $R$ (does not need the Noetherianity assumption).
(3) $\Longleftrightarrow$ (6) This is a particular case of a quite general lemma about an associative ring with a nilpotent ideal [16, Lemma B.10.2].
(2) $\Longrightarrow$ (3) Both in the contexts of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ and $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$, this holds for any adic topological ring $R$ and does not need the Noetherianity assumption. The fact that the $R/I$-module $R/I\otimes_{R}\mathfrak{F}$ is free whenever $\mathfrak{F}$ is either a free contramodule $R$-module or a free quotseparated contramodule $R$-module follows from Lemma 2.5.3 (applied to a free $R$-module $M$).
(1) $\Longleftrightarrow$ (3) In the context of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$, this also holds for any adic topological ring $R$ and does not need the Noetherianity assumption. See [17, Corollary E.1.10(a)] for an even more general result.
(7) $\Longrightarrow$ (6) Obvious.
(1), (3), or (6) $\Longrightarrow$ (7) The assumption that $R$ is Noetherian is essential here. See [16, Corollary B.8.2] or [20, Theorem 10.5] for a complete proof of (1) $\Longleftrightarrow$ (3) $\Longleftrightarrow$ (6) $\Longleftrightarrow$ (7) for Noetherian rings $R$. ∎
A discussion of the functor of contratensor product $\odot_{\mathfrak{R}}$ of discrete right $\mathfrak{R}$-modules and left $\mathfrak{R}$-contramodules, for a complete, separated topological ring $\mathfrak{R}$ where open right ideals form a base of neighborhoods of zero, can be found in [18, Section 3.3], [33, Section 5], [22, Section 2.8], [23, Sections 2.8], [35, Section 7.2], [36, Section 1]. In the context of an adic topological ring $R$ and its adic completion $\mathfrak{R}$, the natural maps
[…]
are isomorphisms for any torsion $R$-module $\mathcal{M}$ and any quotseparated contramodule $R$-module $\mathfrak{P}$. This assertion, which is easy to check directly, follows also from the fact that the forgetful functor $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is fully faithful [35, Lemma 7.11]. For this reason, we will not dwell on the contratensor product functor $\odot_{\mathfrak{R}}$, working with the tensor product $\otimes_{R}$ or $\otimes_{\mathfrak{R}}$ instead.
So, for an adic topological ring $R$, we are interested in the tensor product functor
$\otimes_{R}\colon R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\times R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muR{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}},$
which is well-defined by Lemma 2.2.1. A contramodule $R$-module $\mathfrak{F}$ is said to be (contra)flat (as a contramodule $R$-module) if the tensor product functor $\otimes_{R}\mathfrak{F}\colon\allowbreak R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$ is exact. Let $I\subset R$ be an ideal of definition. Since any short exact sequence of torsion $R$-modules is a direct limit of short exact sequences of $R/I^{n}$-modules, $n\geq 1$, a contramodule $R$-module $\mathfrak{F}$ is flat if and only if the $R/I^{n}$-module $\mathfrak{F}/I^{n}\mathfrak{F}$ is flat for all $n\geq 1$.
For any set $X$ and any torsion $R$-module $\mathcal{M}$, we have a natural isomorphism of torsion $R$-modules
$\mathcal{M}\otimes_{R}\mathfrak{R}[[X]]\simeq\mathcal{M}[X]=\mathcal{M}^{(X)}.$
Therefore, the free quotseparated contramodule $R$-modules are flat (as contramodule $R$-modules), and it follows that the projective quotseparated contramodule $R$-modules are flat, too.
The projective contramodule $R$-modules, i. e., the projective objects of the abelian category $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ are flat contramodule $R$-modules as well [31, Section 3] (see also the proof of Lemma 2.5.5 (2) $\Rightarrow$ (3) above). So flat contramodule $R$-modules need not be separated if they are not quotseparated. However, the following assertion holds.
Lemma 2.5.6.
For any adic topological ring $R$, any flat quotseparated contramodule $R$-module is (complete and) separated. In particular, if the ring $R$ is Noetherian (as an abstract ring), then all flat contramodule $R$-modules are separated.
Proof.
This is a special case of the much more general result of [17, Corollary E.1.7], or of the even more general [33, Corollary 6.15]. The Noetherian adic case can be found in [16, proof of Lemma B.9.2] or [20, Corollary 10.3(b)]. ∎
The following corollary is a partial generalization of [14, Lemma 3.5 and Proposition 3.6].
Corollary 2.5.7.
Let $R$ be an adic topological ring with an ideal of definition $I\subset R$, and let $F$ be an $R$-module. Then the following four conditions are equivalent:
  1. (1)
    the $R/I^{n}$-module $F/I^{n}F$ is flat for every $n\geq 1$;
  2. (2)
    the contramodule $R$-module $\Delta(F)$ is flat;
  3. (3)
    the contramodule $R$-module $\mathbb{L}_{0}\Lambda(F)$ is flat;
  4. (4)
    the contramodule $R$-module $\Lambda(F)$ is flat.
If any one of the conditions (1–4) holds, then the $R$-module $\mathbb{L}_{0}\Lambda(F)$ is complete and separated, and the natural $R$-module map $\mathbb{L}_{0}\Lambda(F)\longrightarrow\Lambda(F)$ is an isomorphism. Consequently, the completion map $F\longrightarrow\Lambda(F)$ induces an isomorphism of $R$-modules $\operatorname{Hom}_{R}(\Lambda(F),\mathfrak{P})\simeq\operatorname{Hom}_{R}(F,\mathfrak{P})$ for any quotseparated contramodule $R$-module $\mathfrak{P}$ in this case.
Proof.
The equivalence of the four conditions (1–4) follows from Lemma 2.5.3. If condition (3) holds, then the $R$-module $\mathbb{L}_{0}\Lambda(F)$ is complete and separated by Lemma 2.5.6. Passing to the projective limit of the isomorphisms from Lemma 2.5.3 over the integers $n\geq 1$, we conclude that $\mathbb{L}_{0}\Lambda(F)\simeq\Lambda(F)$. Since the functor $\mathbb{L}_{0}\Lambda\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ is left adjoint to the inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$, the last assertion of the corollary follows. ∎
Corollary 2.5.8.
Let $R$ be an adic topological ring with an ideal of definition $I\subset R$, and let $\mathfrak{F}$ be a quotseparated contramodule $R$-module. Assume that the separated contramodule $R$-module $\Lambda(\mathfrak{F})$ is flat. Then the $R$-module $\mathfrak{F}$ is separated, so $\mathfrak{F}\simeq\Lambda(\mathfrak{F})$.
Proof.
This can be deduced from the next-to-last assertion of Corollary 2.5.7 by noticing that $\mathfrak{F}\simeq\mathbb{L}_{0}\Lambda(\mathfrak{F})$ for any quotseparated contramodule $R$-module $\mathfrak{F}$. Alternatively, one can say that the quotseparated contramodule $R$-module $\mathfrak{F}$ is flat by Corollary 2.5.7 (4) $\Rightarrow$ (1), and refer to Lemma 2.5.6. ∎
Lemma 2.5.9.
Let $R$ be an adic topological ring and $\mathcal{M}$ be a torsion $R$-module. Then
(a) the full subcategory of flat quotseparated contramodule $R$-modules is closed under extensions and kernels of epimorphisms in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$;
(b) for any short exact sequence of quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{H}\longrightarrow\mathfrak{G}\longrightarrow\mathfrak{F}\longrightarrow 0$ with a flat quotseparated contramodule $R$-module $\mathfrak{F}$, the short sequence of torsion $R$-modules $0\longrightarrow\mathcal{M}\otimes_{R}\mathfrak{H}\longrightarrow\mathcal{M}\otimes_{R}\mathfrak{G}\longrightarrow\mathcal{M}\otimes_{R}\mathfrak{F}\longrightarrow 0$ is exact.
Proof.
The argument is based on the category equivalence/isomorphism $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\simeq\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. Part (a) is a special case of [17, Lemmas E.1.4 and E.1.5]. A further generalization can be found in [33, Corollaries 6.8 and 6.13]. Part (b) is a special case of [33, Lemma 6.10]. ∎
When the ring $R$ is Noetherian (as an abstract ring), one can show that a contramodule $R$-module is flat as a contramodule $R$-module if and only if it is flat as $R$-module. This is the main assertion of the following lemma.
Lemma 2.5.10.
Let $R$ be a Noetherian commutative ring endowed with an adic topology, let $I\subset R$ be an ideal of definition of the topological ring $R$, let $\mathfrak{R}=\Lambda(R)$ be the adic completion of $R$, and let $\mathfrak{F}$ be a contramodule $R$-module. Then the following seven conditions are equivalent:
  1. (1)
    $\mathfrak{F}$ is a flat contramodule $R$-module;
  2. (2)
    $\mathfrak{F}$ is a flat contramodule $\mathfrak{R}$-module;
  3. (3)
    for every $n\geq 1$, the $R/I^{n}$-module $\mathfrak{F}/I^{n}\mathfrak{F}$ is flat;
  4. (4)
    for every ideal of definition $J\subset R$, the $R/J$-module $\mathfrak{F}/J\mathfrak{F}$ is flat;
  5. (5)
    for every open ideal $J\subset R$, the $R/J$-module $\mathfrak{F}/J\mathfrak{F}$ is flat;
  6. (6)
    $\mathfrak{F}$ is a flat $R$-module;
  7. (7)
    $\mathfrak{F}$ is a flat $\mathfrak{R}$-module.
Proof.
(1) $\Longleftrightarrow$ (3) Explained above. No Noetherianity assumption on $R$ is needed here.
(3) $\Longleftrightarrow$ (4) $\Longleftrightarrow$ (5) This is obvious and holds for any adic topological ring $R$ (does not need the Noetherianity assumption).
(1) $\Longleftrightarrow$ (2) Without the Noetherianity assumption, one has to be careful. There is a natural isomorphism of the categories of quotseparated contramodules $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\simeq\mathfrak{R}{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$, as both the categories are isomorphic to $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. Simply put, any quotseparated contramodule $R$-module admits a unique extension of its $R$-module structure to an $\mathfrak{R}$-module structure, making it a quotseparated contramodule $\mathfrak{R}$-module; and any $R$-module morphism of quotseparated contramodule $R$-modules is an $\mathfrak{R}$-module morphism. However, a nonquotseparated contramodule $R$-module does not admit an extension of its $R$-module structure to an $\mathfrak{R}$-module structure, in general. Instead, the nonquotseparated contramodule $R$-modules have natural module structures over the ring $\Delta(R)$[21, Example 5.2(3)].
Nevertheless, in the context of quotseparated contramodules, the equivalence (1) $\Longleftrightarrow$ (2) holds for any adic topological ring $R$ and follows immediately from the equivalence (1) $\Longleftrightarrow$ (3).
(6) $\Longrightarrow$ (5) Obvious.
(3) $\Longrightarrow$ (6) The assumption that $R$ is Noetherian is essential here. A proof can be found in [16, Lemma B.9.2] or in [20, Corollary 10.3(a)].
(2) $\Longleftrightarrow$ (7) This is a particular case of (1) $\Longleftrightarrow$ (6). ∎
Corollary 2.5.11.
Let $R$ and $S$ be Noetherian commutative rings endowed with adic topologies, and let $f\colon R\longrightarrow S$ be a tight continuous ring map. Then the following conditions are equivalent:
  1. (1)
    $f$ is flat as a tight continuous map of adic topological rings;
  2. (2)
    $\Lambda(S)$ is a flat $R$-module;
  3. (3)
    $\Lambda(S)$ is a flat $\Lambda(R)$-module.
Proof.
Notice that $\mathfrak{S}=\Lambda(S)$ is a separated and complete module over the adic topological ring $R$ (as the adic/completion topology on $\mathfrak{S}$ coincides with the topology induced on $\mathfrak{S}$ by the adic topology on $R$, due to the assumptions that $f$ is tight and continuous). Therefore, $\mathfrak{S}$ is a contramodule $R$-module. Furthermore, for every open ideal $I\subset R$, the ideal $Sf(I)\subset S$ is open in $S$, hence $S/Sf(I)\simeq\mathfrak{S}/\mathfrak{S}\lambda_{S}(f(I))$. The assertion of the Corollary can be obtained by applying Lemma 2.5.10 (5) $\Leftrightarrow$ (6) $\Leftrightarrow$ (7) to the contramodule $R$-module $\mathfrak{F}=\mathfrak{S}$. ∎
Proposition 2.5.12.
Let $R$ be a Noetherian commutative ring endowed with an adic topology. Then the fully faithful inclusion of abelian categories $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ induces isomorphisms on all the Ext groups/modules.
Proof.
This is [16, Theorem B.8.1]. For a stronger result concerning the full-and-faithfulness of the induced functor between the unbounded derived categories, see [19, Theorem 2.9]. The Noetherianity assumption can be weakened, but it cannot be simply dropped; see [24, Theorems 4.2 and 4.3]. ∎
2.6. Change of scalars
Given a homomorphism of associative rings $f\colon R\longrightarrow S$, we denote by $f_{*}\colon S{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ the functor of restriction of scalars, taking every left $S$-module NN to its underlying left $R$-module NN. The left adjoint functor to $f_{*}$, taking every left $R$-module $M$ to the left $S$-module $S\otimes_{R}M$, will be denoted by $f^{*}\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow S{\operatorname{\mathsf{--Mod}}}$; it is called the functor of extension of scalars with respect to $f$. The right adjoint functor to $f_{*}$, taking every left $R$-module $P$ to the left $S$-module $\operatorname{Hom}_{R}(S,P)$, will be denoted by $f^{!}\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow S{\operatorname{\mathsf{--Mod}}}$ and called the functor of coextension of scalars.
Let $R$ and $S$ be adic topological rings, and let $f\colon R\longrightarrow S$ be a ring homomorphism. Parts (b) and (d) of the following lemma are generalizations of [17, Lemma D.4.1(a)].
Lemma 2.6.1.
Assume that the ring homomorphism $f$ is continuous. Consider the functor of restriction of scalars $f_{*}\colon S{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$. Then
(a) the forgetful functor $f_{*}$ takes torsion $S$-modules to torsion $R$-modules;
(b) the forgetful functor $f_{*}$ takes contramodule $S$-modules to contramodule $R$-modules;
(c) the forgetful functor $f_{*}$ takes separated $S$-modules to separated $R$-modules;
(d) the forgetful functor $f_{*}$ takes quotseparated contramodule $S$-modules to quotseparated contramodule $R$-modules.
Proof.
Recall that, by Lemma 2.1.3(2), there exist an ideal of definition $I\subset R$ and an ideal of definition $J\subset S$ such that $f(I)\subset J$.
Part (a): more generally, for any homomorphism of commutative rings $f\colon R\longrightarrow S$ and any two ideals $I\subset R$ and $J\subset S$ such that $f(I)\subset J$, the functor $S{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ takes $J$-torsion $S$-modules to $I$-torsion $R$-modules. Indeed, for any $S$-module NN and any fixed element $r\in R$, an element $y\in N$ is $f(r)$-torsion (as an element of the $S$-module NN) if and only if $y$ is $r$-torsion (as an element of the underlying $R$-module of $N$).
Part (b): once again, for any homomorphism of commutative rings $f\colon R\longrightarrow S$ and any two ideals $I\subset R$ and $J\subset S$ such that $f(I)\subset J$, the functor $S{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ takes $J$-contramodule $S$-modules to $I$-contramodule $R$-modules. The point is that, for any $S$-module $Q$ and any fixed element $r\in R$, there is a natural isomorphism of the Ext groups $\operatorname{Ext}^{*}_{S}(S[f(r)^{-1}],Q)\simeq\operatorname{Ext}^{*}_{R}(R[r^{-1}],Q)$, because $R[r^{-1}]$ is a flat $R$-module and $S\otimes_{R}R[r^{-1}]\simeq S[f(r)^{-1}]$. See, e. g., [34, Lemma 4.1(a)]; cf. [20, Lemma 2.1]. So the $S$-module $Q$ is an $f(r)$-contramodule if and only if the underlying $R$-module of $Q$ is an $r$-contramodule.
Part (c): for every $S$-module $C$ and every $n\geq 1$, one has $I^{n}C\subset J^{n}C\subset C$. Hence $\bigcap_{n\geq 1}J^{n}C=0$ implies $\bigcap_{n\geq 1}I^{n}C=0$.
Part (d) follows from parts (b) and (c). Recall that all contramodule $S$-modules are complete, and all separate complete $S$-modules are contramodules, as mentioned in Section 2.4. So the class of separated and complete $S$-modules coincides with the class of separated contramodule $S$-modules.
Let $\mathfrak{Q}$ be a quotseparated contramodule $S$-module. Then $\mathfrak{Q}$ is a contramodule $S$-module and a quotient $S$-module of a separated contramodule $S$-module $\mathfrak{C}$. Then, by (b) and (c), the underlying $R$-module of $\mathfrak{Q}$ is a contramodule $R$-module and a quotient $R$-module of a separated contramodule $R$-module $\mathfrak{C}$. ∎
Let $f\colon R\longrightarrow S$ be a continuous homomorphism of adic topological rings. Then Lemma 2.6.1(a) provides an exact, faithful functor of restriction of scalars
$f_{\diamond}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muR{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}.$
The functor $f_{\diamond}$ has a right adjoint functor
$f^{\diamond}\colon R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muS{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}},$
which can be computed by the formula $f^{\diamond}(\mathcal{M})=\Gamma_{J}(\operatorname{Hom}_{R}(S,\mathcal{M}))$ for any $\mathcal{M}\in R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$, where $J$ is an ideal of definition in $S$. Moreover, for any $R$-module $M$ one has a natural isomorphism of torsion $S$-modules
$f^{\diamond}(\Gamma_{I}(M))\simeq\Gamma_{J}(f^{!}(M)),$
where $I$ is an ideal of definition in $R$. As a right adjoint functor to an exact functor, the functor $f^{\diamond}$ takes injective torsion $R$-modules to injective torsion $S$-modules. The same assertions hold in the greater generality of a commutative ring homomorphism $f\colon R\longrightarrow S$ and any two ideals $I\subset R$,  $J\subset S$ such that $f(I)\subset J$[31, proof of Lemma 15.1(a)]. See [23, Section 2.9] and [32, Section 6] for a generalization to the categories of discrete modules over topological associative rings.
We call $f^{\diamond}$ the functor of coextension of scalars (for torsion modules, with respect to a continuous homomorphism of adic topological rings).
Furthermore, Lemma 2.6.1(b) provides an exact, faithful functor of restriction of scalars
$f_{\#}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muR{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}.$
The functor $f_{\#}$ has a left adjoint functor
$f^{\#}\colon R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muS{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}},$
which can be computed by the formula $f^{\#}(\mathfrak{P})=\Delta_{J}(S\otimes_{R}\mathfrak{P})$ for any $\mathfrak{P}\in R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$, where $J$ is an ideal of definition in $S$. Moreover, for any $R$-module $P$ one has a natural isomorphism of contramodule $S$-modules
$f^{\#}(\Delta_{I}(P))\simeq\Delta_{J}(f^{*}(P)),$
where $I$ is an ideal of definition in $R$. As a left adjoint functor to an exact functor, the functor $f^{\#}$ takes projective contramodule $R$-modules to projective contramodule $S$-modules. Specifically, one has $f^{\#}(\Delta_{I}(R[X]))=\Delta_{J}(S[X])$ for any set $X$ (so the functor $f^{\#}$ also takes free contramodule $R$-modules to free contramodule $S$-modules). The same assertions hold in the greater generality of a commutative ring homomorphism $f\colon R\longrightarrow S$ and any two ideals $I\subset R$,  $J\subset S$ such that $f(I)\subset J$[31, proof of Lemma 15.1(b)].
We call $f^{\#}$ the functor of contraextension of scalars (for contramodules, with respect with respect to a continuous homomorphism of adic topological rings).
Finally, Lemma 2.6.1(d) provides an exact, faithful functor of restriction of scalars
$f_{\sharp}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muR{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}.$
The functor $f_{\sharp}$ has a left adjoint functor
$f^{\sharp}\colon R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muS{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}},$
which can be computed by the formula $f^{\sharp}(\mathfrak{P})=\mathbb{L}_{0}\Lambda_{J}(S\otimes_{R}\mathfrak{P})$ for any $\mathfrak{P}\in R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$, where $J$ is an ideal of definition in SS[31, proof of Lemma 15.1(c)]. Moreover, for any $R$-module $P$ one has a natural isomorphism of quotseparated contramodule $S$-modules
$f^{\sharp}(\mathbb{L}_{0}\Lambda_{I}(P))\simeq\mathbb{L}_{0}\Lambda_{J}(f^{*}(P)),$
where $I$ is an ideal of definition in $R$. As a left adjoint functor to an exact functor, the functor $f^{\sharp}$ takes projective quotseparated contramodule $R$-modules to projective quotseparated contramodule $S$-modules. Specifically, one has $f^{\sharp}(\Lambda_{I}(R[X]))=\Lambda_{J}(S[X])$ for any set $X$ (so the functor $f^{\sharp}$ also takes free quotseparated contramodule $R$-modules to free quotseparated contramodule $S$-modules). See [23, Section 2.9] and [32, Section 6] for a generalization to the categories of contramodules over topological associative rings.
The functor $f^{\sharp}$ is called the functor of contraextension of scalars (for quotseparated contramodules, with respect with respect to a continuous homomorphism of adic topological rings).
Lemma 2.6.2.
Let $R$ be an adic topological ring, $S$ be a discrete (i. e., $(0)$-adic) commutative ring, and $f\colon R\longrightarrow S$ be a continuous ring homomorphism. Then
(a) for every torsion $R$-module $\mathcal{M}$, there is a natural isomorphism of $S$-modules
$f^{\diamond}(\mathcal{M})\simeq f^{!}(\mathcal{M})=\operatorname{Hom}_{R}(S,\mathcal{M});$
(b) for every contramodule $R$-module $\mathfrak{P}$, there is a natural isomorphism of $S$-modules
$f^{\#}(\mathfrak{P})\simeq f^{*}(\mathfrak{P})=S\otimes_{R}\mathfrak{P};$
(c) for every quotseparated contramodule $R$-module $\mathfrak{P}$, there is a natural isomorphism of $S$-modules
$f^{\sharp}(\mathfrak{P})\simeq f^{*}(\mathfrak{P})=S\otimes_{R}\mathfrak{P}.$
Proof.
Part (a): more generally, given a continuous homomorphism of adic topological rings $f\colon R\longrightarrow S$ and a torsion $R$-module $\mathcal{M}$, one has $f^{\diamond}(\mathcal{M})=f^{!}(\mathcal{M})$ whenever $f^{!}(\mathcal{M})$ is a torsion $S$-module. When the ring $S$ is discrete, all $S$-modules are torsion.
Part (b): more generally, given a continuous homomorphism of adic topological rings $f\colon R\longrightarrow S$ and a contramodule $R$-module $\mathfrak{P}$, one has $f^{\#}(\mathfrak{P})=f^{*}(\mathfrak{P})$ whenever $f^{*}(\mathfrak{P})$ is a contramodule $S$-module. When the ring $S$ is discrete, all $S$-modules are contramodules.
Part (c): more generally, given a continuous homomorphism of adic topological rings $f\colon R\longrightarrow S$ and a quotseparated contramodule $R$-module $\mathfrak{P}$, one has $f^{\sharp}(\mathfrak{P})=f^{*}(\mathfrak{P})$ whenever $f^{*}(\mathfrak{P})$ is a quotseparated contramodule $S$-module. When the ring $S$ is discrete, all $S$-modules are quotseparated contramodules. ∎
Lemma 2.6.3.
Let $f\colon R\longrightarrow S$ be a continuous homomorphism of adic topological rings, and let $I\subset R$ and $J\subset S$ be two ideals of definition such that $f(I)\subset J$. Then
(a) for every torsion $R$-module $\mathcal{M}$, there is a natural isomorphism of $S/J$-modules
$_{J}f^{\diamond}(\mathcal{M})\simeq\operatorname{Hom}_{R/I}(S/J,{}_{I}\mathcal{M});$
(b) for every contramodule $R$-module $\mathfrak{P}$, there is a natural isomorphism of $S/J$-modules
$f^{\#}(\mathfrak{P})/Jf^{\#}(\mathfrak{P})\simeq S/J\otimes_{R/I}(\mathfrak{P}/I\mathfrak{P});$
(c) for every quotseparated contramodule $R$-module $\mathfrak{P}$, there is a natural isomorphism of $S/J$-modules
$f^{\sharp}(\mathfrak{P})/Jf^{\sharp}(\mathfrak{P})\simeq S/J\otimes_{R/I}(\mathfrak{P}/I\mathfrak{P}).$
Proof.
Let us view $R/I$ and $S/J$ as adic topological rings with the discrete topologies. Denote the related continuous homomorphisms of adic topological rings by $p\colon R\longrightarrow R/I$, $p^{\prime}\colon S\longrightarrow S/J$, and $f^{\prime}\colon R/I\longrightarrow S/J$. Then we have a commutative square diagram of continuous homomorphisms of adic topological rings $p^{\prime}f=f^{\prime}p$. Hence the commutative diagrams of the related functors of coextension and contraextension of scalars $p^{\prime}{}^{\diamond}f^{\diamond}\simeq f^{\prime}{}^{\diamond}p^{\diamond}$, $p^{\prime}{}^{\#}f^{\#}\simeq f^{\prime}{}^{\#}p^{\#}$, and $p^{\prime}{}^{\sharp}f^{\sharp}\simeq f^{\prime}{}^{\sharp}p^{\sharp}$. It remains to use Lemma 2.6.2 for the computation of the functors of co/contraextension of scalars with respect to $p$, $p^{\prime}$, and $f^{\prime}$ in order to obtain the desired natural isomorphisms. ∎
The following corollary is a generalization of [17, Lemma D.4.3].
Corollary 2.6.4.
Let $f\colon R\longrightarrow S$ be a continuous homomorphism of adic topological rings, and let $I\subset R$ and $J\subset S$ be two ideals of definition such that $f(I)\subset J$. Then, for any flat quotseparated contramodule $R$-module $\mathfrak{F}$, there is a natural isomorphism of quotseparated contramodule $S$-modules
[…]
In particular, the quoseparated contramodule $S$-module $f^{\sharp}(\mathfrak{F})$ is flat. The functor $f^{\sharp}$ takes short exact sequences of flat quotseparated contramodule $R$-modules to short exact sequences of flat quotseparated contramodule $S$-modules.
Furthermore, the contramodule $S$-module $f^{\#}(\mathfrak{F})$ is flat for any flat contramodule $S$-module $\mathfrak{F}$.
Proof.
All the assertions are based on Lemma 2.6.3(b–c). The flatness of $f^{\sharp}(\mathfrak{F})$ and $f^{\#}(\mathfrak{F})$ can be established, under the respective assumptions, using the facts that a contramodule $R$-module $\mathfrak{F}$ is flat if and only if the $R/I^{n}$-module $\mathfrak{F}/I^{n}\mathfrak{F}$ is flat for every $n\geq 1$, and similarly for contramodule $S$-modules (see Section 2.5). The formula for $f^{\sharp}(\mathfrak{F})$ follows by virtue of Lemma 2.5.6. The exactness assertion holds in view of Lemma 2.5.9(b), as the functor of projective limit of sequences of surjective maps of abelian groups is exact. Alternatively, one can also apply the argument from [17, proof of Lemma D.4.3] using Lemma 2.5.9. The result of [41, Theorem 1.2 or 2.8] or [17, Lemma E.1.3] is relevant here. ∎
Corollary 2.6.5.
Let $f\colon R\longrightarrow S$ be a continuous homomorphism of adic topological rings, and let $0\longrightarrow\mathfrak{Q}\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{F}\longrightarrow 0$ be a short exact sequence of quotseparated contramodule $R$-modules with a flat quotseparated contramodule $R$-module $\mathfrak{F}$. Then the short sequence of quotseparated contramodule $S$-modules $0\longrightarrow f^{\sharp}(\mathfrak{Q})\longrightarrow f^{\sharp}(\mathfrak{P})\longrightarrow f^{\sharp}(\mathfrak{F})\longrightarrow 0$ is exact.
Proof.
This is a special case of the second assertion of [32, Lemma 6.2]. The claim in question follows purely formally from the exactness assertion of Corollary 2.6.4. One needs to use the facts that the functor $f^{\sharp}$ is right exact (as a left adjoint functor), there are enough projective quotseparated contramodule $R$-modules in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$, the projective quotseparated contramodule $R$-modules are flat (see Section 2.5), and the kernels of surjective morphisms of flat quotseparated contramodule $R$-modules are flat (see Lemma 2.5.9(a)). ∎
Now we pass from the continuous maps of adic topological rings to the tight ones. Part (b) of the following lemma is a generalization of [17, Lemma D.4.1(b)].
Lemma 2.6.6.
Let $R$ and $S$ be adic topological rings, and let $f\colon R\longrightarrow S$ be a tight ring map. In this context:
(a) if $\mathcal{N}$ is an $S$-module such that the $R$-module $f_{*}(\mathcal{N})$ is torsion, then the $S$-module $\mathcal{N}$ is torsion;
(b) if $\mathfrak{Q}$ is an $S$-module such that the $R$-module $f_{*}(\mathfrak{Q})$ is a contramodule, then the $S$-module $\mathfrak{Q}$ is a contramodule.
Proof.
Recall that, by Lemma 2.1.4(1), there exist an ideal of definition $I\subset R$ and an ideal of definition $J\subset S$ such that $J\subset Sf(I)$. Put $J^{\prime}=Sf(I)$; so we have $J\subset J^{\prime}$.
Part (a): as mentioned in Section 2.2, an $S$-module $\mathcal{N}$ is $J^{\prime}$-torsion whenever it is $s$-torsion for all $s$ ranging over some chosen set of generators of the ideal $J^{\prime}\subset S$. The ideal $J^{\prime}$ is generated by its subset $f(I)\subset J^{\prime}$. For any element $r\in I$, the underlying $R$-module of $\mathcal{N}$ is $r$-torsion by assumption. According to the proof of Lemma 2.6.1(a), it follows that the $S$-module $\mathcal{N}$ is $f(r)$-torsion. Thus $\mathcal{N}$ is $J^{\prime}$-torsion, and therefore also $J$-torsion.
Part (b): as mentioned in Section 2.4, an $S$-module $\mathfrak{Q}$ is a $J^{\prime}$-contramodule whenever it is an $s$-contramodule for all $s$ ranging over some chosen set of generators of the ideal $J^{\prime}\subset S$. The ideal $J^{\prime}$ is generated by its subset $f(I)\subset J^{\prime}$. For any element $r\in I$, the underlying $R$-module of $\mathfrak{Q}$ is an $r$-contramodule by assumption. According to the proof of Lemma 2.6.1(b), it follows that the $S$-module $\mathfrak{Q}$ is an $f(r)$-contramodule. Thus $\mathfrak{Q}$ is a $J^{\prime}$-contramodule, and therefore also a $J$-contramodule. ∎
A quotseparated version of Lemma 2.6.6(b) holds under an additional flatness assumption on the morphism $f$. See Theorem 2.11.2 and Remark 2.11.3 below.
Part (b) of the following lemma is a generalization of [17, Lemma D.4.2(a)].
Lemma 2.6.7.
Let $R$ and $S$ be adic topological rings, and let $f\colon R\longrightarrow S$ be a tight ring map. Then
(a) for any torsion $R$-module $\mathcal{M}$, the $S$-module $f^{*}(\mathcal{M})=S\otimes_{R}\mathcal{M}$ is also torsion;
(b) for any contramodule $R$-module $\mathfrak{P}$, the $S$-module $f^{!}(\mathfrak{P})=\operatorname{Hom}_{R}(S,\mathfrak{P})$ is also a contramodule.
Proof.
Part (a): by Lemma 2.2.1, the $R$-module $S\otimes_{R}\mathcal{M}$ is torsion. By Lemma 2.6.6(a), it follows that the $S$-module $S\otimes_{R}\mathcal{M}$ is torsion.
Part (b): by Lemma 2.4.1(b), the $R$-module $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is a contramodule. By Lemma 2.6.6(b), it follows that the $S$-module $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is a contramodule. ∎
Let $f\colon R\longrightarrow S$ be a tight map of adic topological rings. Then Lemma 2.6.7(a) claims that the functor of extension of scalars $f^{*}\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow S{\operatorname{\mathsf{--Mod}}}$ restricts to a functor
$f^{*}\colon R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muS{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}},$
which we will also call the extension of scalars. When the map $f$ is both continuous and tight, the functor $f^{*}$ is left adjoint to the functor of restriction of scalars $f_{\diamond}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$.
Dual-analogously, for any tight map of adic topological rings $f\colon R\longrightarrow S$, Lemma 2.6.7(b) claims that the functor of coextension of scalars $f^{!}\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow S{\operatorname{\mathsf{--Mod}}}$ restricts to a functor
$f^{!}\colon R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muS{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}},$
which we will also call the coextension of scalars. When the map $f$ is both continuous and tight, the functor $f^{!}$ is right adjoint to the functor of restriction of scalars $f_{\#}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$.
For a quotseparated version of Lemma 2.6.7(b), which we can only prove under a flatness assumption on the map $f$, see Proposition 2.11.5 below.
Remark 2.6.8.
For a continuous map of adic topological rings $f\colon R\longrightarrow S$ that is not tight, the functor of restriction of scalars $f_{\diamond}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$ does not have a left adjoint functor, generally speaking. It suffices to consider the case of the ring of polynomials in one variable $R=S=k[x]$ over a field $k$, with the identity map $f=\mathrm{id}_{k[x]}\colon R\longrightarrow S$, the $(0)$-adic (i. e., discrete) topology on $R$, and the $(x)$-adic topology on $S$. Then the forgetful functor $f_{\diamond}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$ is fully faithful, but it does not preserve infinite direct products (because the full subcategory of $x$-torsion modules is not closed under infinite products in $k[x]{\operatorname{\mathsf{--Mod}}}$). Accordingly, the functor $f_{\diamond}$ is not a right adjoint.
Similarly, for a continuous map of adic topological rings $f\colon R\longrightarrow S$ that is not tight, the functors of restriction of scalars $f_{\#}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ and $f_{\sharp}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ do not have right adjoint functors, generally speaking. In the same example of adic topological rings $R$, $S$ and the map $f$ as above, the forgetful functor $f_{\#}=f_{\sharp}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ is fully faithful, but it does not preserve infinite direct sums (because the full subcategory of $x$-contramodule modules is not closed under infinite direct sums in $k[x]{\operatorname{\mathsf{--Mod}}}$). Accordingly, the functor $f_{\#}=f_{\sharp}$ is not a left adjoint. (See the discussion in [23, Section 2.9] and the details in [32, Example 10.1].)
The definition of a flat map of adic topological rings was given at the end of Section 2.1.
Lemma 2.6.9.
Let $f\colon R\longrightarrow S$ be a flat tight continuous map of adic topological rings. Then
(a) the functor $f^{*}\colon R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\longrightarrow S{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$ is exact;
(b) the functor $f_{\diamond}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$ takes injective torsion $S$-modules to injective torsion $R$-modules;
(c) the functor $f_{\#}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ takes flat contramodule $S$-modules to flat contramodule $R$-modules;
(d) the functor $f_{\sharp}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ takes flat quotseparated contramodule $S$-modules to flat quotseparated contramodule $R$-modules.
Proof.
The assertions of parts (a) and (b) are equivalent restatements of each other, because there are enough injective objecs in $S{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$. Nevertheless, we prove them separately below. Let $I\subset R$ be an ideal of definition in $R$; then, by Lemma 2.1.6(5),  $Sf(I)\subset S$ is an ideal of definition in $S$.
Part (a): every short exact sequence of $R$-modules is a direct limit of short exact sequences of $R/I^{n}$-modules, $n\geq 1$. For any $R/I^{n}$-module $E$, one has $f^{*}(E)=S\otimes_{R}E\simeq S/Sf(I^{n})\otimes_{R/I^{n}}E$. Since $S/Sf(I^{n})$ is a flat module over $R/I^{n}$, the functor $f^{*}$ takes short exact sequences of $R/I^{n}$-modules to short exact sequences of $S/Sf(I^{n})$-modules.
Part (b): for any $S$-module $K$, we have $f_{*}({}_{Sf(I)}K)={}_{I}f_{*}(K)$. For an injective torsion $S$-module $\mathcal{K}$, the $S/Sf(I)$-module $_{Sf(I)}\mathcal{K}$ is injective. Since $S/Sf(I)$ is a flat $R/I$-module, it follows that $_{I}f_{\diamond}(\mathcal{K})$ is injective as an $R/I$-module, as desired.
Part (c): for any $S$-module $P$, we have $f_{*}(P/Sf(I)P)=f_{*}(P)/If_{*}(P)$. For a flat contramodule $S$-module $\mathfrak{F}$, the $S/Sf(I)$-module $\mathfrak{F}/Sf(I)\mathfrak{F}$ is flat. Since $S/Sf(I)$ is a flat $R/I$-module, it follows that $f_{\#}(\mathfrak{F})/If_{\#}(\mathfrak{F})$ is flat as an $R/I$-module, as desired.
Part (d) is the particular case of part (c) for quotseparated contramodules. ∎
Let $R$ be an adic topological ring. Then the completion map $\lambda_{R}\colon R\longrightarrow\mathfrak{R}$ is a flat tight continuous map of adic topological rings (as mentioned in Section 2.3).
Clearly, any torsion $R$-module admits a unique extension of its $R$-module structure to an $\mathfrak{R}$-module structure. The functor $\lambda_{R}\mskip 1.5mu{}_{\diamond}\colon\mathfrak{R}{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$ is an equivalence (in fact, isomorphism) of abelian categories. Consequently, the functors $\lambda_{R}^{*}\colon R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$ and $\lambda_{R}^{\diamond}\colon R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$ adjoint to $\lambda_{R}\mskip 1.5mu{}_{\diamond}$ on the left and on the right are also equivalences of categories, isomorphic to each other and inverse to $\lambda_{R}\mskip 1.5mu{}_{\diamond}$.
Furthermore, it is clear from the discussion of the category of contramodules over the topological ring $\mathfrak{R}$ in Section 2.4 that the functor $\lambda_{R}\mskip 1.5mu{}_{\sharp}\colon\mathfrak{R}{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ is an equivalence (in fact, isomorphism) of abelian categories. Consequently, the functors $\lambda_{R}^{\sharp}\colon R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ and $\lambda_{R}^{!}\colon R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ adjoint to $\lambda_{R}\mskip 1.5mu{}_{\sharp}$ on the left and on the right are also equivalences of categories, isomorphic to each other and inverse to $\lambda_{R}\mskip 1.5mu{}_{\sharp}$.
Notice that the functor $\lambda_{R}\mskip 1.5mu{}_{\#}\colon\mathfrak{R}{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ is always fully faithful, but it need not be an equivalence of categories in general. See the discussion in the proof of Lemma 2.5.10.
2.7. Tensor products of adic topological rings
Let $R$ be a commutative ring, let $S$ and $T$ be two adic topological rings, and let $f\colon R\longrightarrow S$ and $g\colon R\longrightarrow T$ be two ring homomorphisms. Consider the ring $W=T\otimes_{R}S$, and denote by $f^{\prime}\colon T\longrightarrow T\otimes_{R}S$ and $g^{\prime}\colon S\longrightarrow T\otimes_{R}S$ the two induced ring homomorphisms.
Lemma 2.7.1.
(a) There exists a unique finest ring topology on the ring $W$ with a base of neighborhoods of zero formed by open ideals such that the ring maps $f^{\prime}$ and $g^{\prime}$ are continuous.
(b) There exists a unique coarsest ring topology on the ring $W$ with a base of neighborhoods of zero formed by open ideals such that, for every topological commutative ring $V$ with a base of neighborhoods of zero formed by open ideals, and for every ring homomorphism $h\colon W\longrightarrow V$ for which the compositions $hf^{\prime}\colon T\longrightarrow V$ and $hg^{\prime}\colon S\longrightarrow V$ are continuous, the ring homomorphism $h$ is continuous as well.
(c) The ring topologies on $W$ defined in parts (a) and (b) coincide, and it is an adic topology.
(d) Let $J\subset S$ be an ideal of definition in $S$ and $K\subset T$ be an ideal of definition in $T$. Then $L=W(f^{\prime}(K)+g^{\prime}(J))\subset W$ is an ideal of definition in the adic topology on $W$ defined in parts (a–c).
Proof.
Consider the adic topology on $W$ with the ideal of definition constructed in part (d). Let us check that this topology on $W$ satisfies the conditions stated in parts (a) and (b).
Part (a): the ring maps $f^{\prime}$ and $g^{\prime}$ are continuous by Lemma 2.1.3(2). Conversely, let $\tau$ be a ring topology on $W$, with a base of neighborhoods of zero formed by open right ideals, such that the maps $f^{\prime}$ and $g^{\prime}$ are continuous with respect to $\tau$. To show that $\tau$ is equal to or coarser than the adic topology on $W$ defined above, we need to check that all the ideals in $W$ that are open in $\tau$ are also open in our adic topology. In other words, we need to check that for every ideal $E\subset W$ that is open in $\tau$ there exists an integer $n\geq 1$ such $L^{n}\subset E$.
Indeed, since the map $f^{\prime}\colon T\longrightarrow W$ is continuous with respect to $\tau$, there exists an integer $n_{1}\geq 1$ such that $f^{\prime}(K^{n_{1}})\subset E$. Similarly, since the map $g^{\prime}\colon S\longrightarrow W$ is continuous with respect to $\tau$, there exists an integer $n_{2}\geq 1$ such that $g^{\prime}(J^{n_{2}})\subset E$. Put $n=n_{1}+n_{2}-1$. Then we have $(f^{\prime}(K)+g^{\prime}(J))^{n}\subset E$, hence $L^{n}\subset E$.
Part (b): firstly we assume that the compositions $hf^{\prime}$ and $hg^{\prime}$ are continuous, and prove that the map $h$ is continuous with respect to our adic topology on $W$. Let $E\subset V$ be an open ideal. We need to show that there exists an integer $n\geq 1$ such that $h(L^{n})\subset E$. Indeed, since the map $hf^{\prime}$ is continuous, there exists an integer $n_{1}\geq 1$ such that $hf^{\prime}(K^{n_{1}})\subset E$. Similarly, since the map $hg^{\prime}$ is continuous, there exists an integer $n_{2}\geq 1$ such that $hg^{\prime}(J^{n_{2}})\subset E$. Put $n=n_{1}+n_{2}-1$. Then we have $(hf^{\prime}(K)+hg^{\prime}(J))^{n}\subset E$, hence $h(L^{n})\subset E$.
Now let $\tau$ be a ring topology on $W$ such that, for every topological commutative ring $V$ with a base of neighborhoods of zero formed by open ideals, and for every ring homomorphism $h\colon W\longrightarrow V$ for which the compositions $hf^{\prime}\colon T\longrightarrow V$ and $hg^{\prime}\colon S\longrightarrow V$ are continuous, the map $h$ is continuous with respect to $\tau$. To show that $\tau$ is equal to or finer than the adic topology on $W$ defined above, we need to check that all the ideals in $W$ that are open in our adic topology are also open in $\tau$.
Put $V=W$, and endow the ring $V$ with our adic topology on $W$. Let $h\colon W\longrightarrow V$ be the identity map. Then the maps $hf^{\prime}$ and $hg^{\prime}$ are continuous, as we have shown in the very beginning of the proof of part (a) above. So the map $h$ is continuous with respect to the topology $\tau$ on $W$ and our adic topology on $V$. This means precisely that $\tau$ is equal to or finer than our adic topology on $W$. ∎
The adic topology on the ring $W=T\otimes_{R}S$ constructed in Lemma 2.7.1 will be called the tensor product topology.
In the sequel, we will sometimes abuse the notation and write $L=K\otimes_{R}S+T\otimes_{R}J\subset T\otimes_{R}S=W$ instead of $L=W(f^{\prime}(K)+g^{\prime}(J))\subset W$ in the context of Lemma 2.7.1(d). Notice that ideals of the form $L=K\otimes_{R}S+T\otimes_{R}J$, where $J$ ranges over the ideals of definition in $S$ and $K$ ranges over the ideals of definition in $T$, form a base of neighborhoods of zero in the tensor product topology on $T\otimes_{R}S$.
Lemma 2.7.2.
(a) Let $f\colon R\longrightarrow S$ be a continuous homomorphism of adic topological rings. Endow the tensor product $R\otimes_{R}S$ with the tensor product topology. Then the natural ring isomorphisms $S\longrightarrow R\otimes_{R}S\longrightarrow S$ are isomorphisms of topological rings.
(b) Let $R$ and $S$ be commutative rings, $T$, $U$, and $V$ be adic topological rings, and $R\longrightarrow T$,  $R\longrightarrow U$,  $S\longrightarrow U$, and $S\longrightarrow V$ be ring homomorphisms. Endow all the relevant tensor products with the tensor product topologies. Then the natural ring isomorphism $(T\otimes_{R}U)\otimes_{S}V\simeq T\otimes_{R}(U\otimes_{S}V)$ is an isomorphism of topological rings.
Proof.
Part (a) Let $J\subset S$ be an ideal of definition. By Lemma 2.1.3(3), there exists an ideal of definition $I\subset R$ such that $f(I)\subset J$. Now the isomorphism $R\otimes_{R}S\simeq S$ transforms the ideal $I\otimes_{R}S+R\otimes_{R}J=R\otimes_{R}J\subset R\otimes_{R}S$ into the ideal $J\subset S$. By Lemma 2.7.1(d),  $I\otimes_{R}S+R\otimes_{R}J$ is an ideal of definition in $R\otimes_{R}S$. Thus $R\otimes_{R}S\simeq S$ is an isomorphism of topological rings. Part (b) is obvious. ∎
Lemma 2.7.3.
In the notation above, endow the ring $W=T\otimes_{R}S$ with the tensor product topology. Assume that the ring $R$ is endowed with an adic topology such that $f\colon R\longrightarrow S$ is a tight continuous ring map and $g\colon R\longrightarrow T$ is a continuous ring map. Then
(a) the continuous ring map $f^{\prime}\colon T\longrightarrow W$ is also tight;
(b) if the tight continuous ring map $f$ is flat, then the tight continuous ring map $f^{\prime}$ is flat, too;
(c) if the ring map $\Lambda(f)\colon\Lambda(R)\longrightarrow\Lambda(S)$ is a (topological) isomorphism, then the ring map $\Lambda(f^{\prime})\colon\Lambda(T)\longrightarrow\Lambda(W)$ is an isomorphism as well.
Proof.
Part (a): by Lemma 2.1.3(2), there exist an ideal of definition $I\subset R$ and an ideal of definition $K\subset T$ such that $g(I)\subset K$. By Lemma 2.1.6(5), the ideal $J=Sf(I)\subset S$ is an ideal of definition in $S$. By Lemma 2.7.1(4), the ideal $L=W(f^{\prime}(K)+g^{\prime}(J))\subset W$ is an ideal of definition in $W$. Now we have $W(f^{\prime}(K)+g^{\prime}(J))=W(f^{\prime}(K)+g^{\prime}f(I))=W(f^{\prime}(K)+f^{\prime}g(I))=Wf^{\prime}(K)$. By Lemma 2.1.4(1) or 2.1.6(4), it follows that the map $f^{\prime}$ is tight.
Part (b): now we apply Lemma 2.1.3(3) to the effect that, for every ideal of definition $K\subset T$, there exists an ideal of definition $I\subset R$ such that $f(I)\subset K$. Then, in the notation of the previous paragraph, we have $W/L\simeq T/K\otimes_{R/I}S/J$. The ring $S/J$ is flat as an $R/I$-module by Lemma 2.1.8(2), hence the ring $W/L$ is flat as a $T/K$-module.
Part (c): first of all, for a tight continuous map $f$, the map $\Lambda(f)$ is also tight (and continuous) by Lemma 2.3.1(a). Hence $\Lambda(f)$ is an isomorphism of abstract rings if and only if $\Lambda(f)$ is an isomorphism of topological rings.
Now let $K\subset T$ be an ideal of definition. Once again, in the notation above, we have $W/L\simeq T/K\otimes_{R/I}S/J$. If $\Lambda(f)$ is an isomorphism, then so is the map $R/I\longrightarrow S/J$. Hence the map $T/K\longrightarrow W/L$ is an isomorphism, too. As this holds for all ideals of definition $K\subset T$, we can pass to the projective limit and conclude that the map $\Lambda(f^{\prime})$ is an isomorphism. ∎
Now consider the following commutative diagram of homomorphisms of commutative rings:
S~T~ShSR~f~g~ThTRfhRg
[…]
Assume that adic topologies are given on the rings $S$, $\widetilde{S}$, $T$, and $\widetilde{T}$ such that $h_{S}$ and $h_{T}$ are continuous ring maps. Consider the tensor product rings $T\otimes_{R}S$ and $\widetilde{T}\otimes_{\widetilde{R}}\widetilde{S}$, and endow them with the tensor product topologies. Then the homomorphism of tensor products $h\colon T\otimes_{R}S\longrightarrow\widetilde{T}\otimes_{\widetilde{R}}\widetilde{S}$ induced by the ring homomorphisms $h_{S}$, $h_{T}$, and $h_{R}$ is a continuous ring map.
The following lemma is a generalization of Lemma 2.7.3(c).
Lemma 2.7.4.
Let $S$, $T$, and $R$ be adic topological rings, and let $f\colon R\longrightarrow S$ and $g\colon R\longrightarrow T$ be continuous ring maps. Consider the induced continuous maps of the completed rings $\Lambda(f)\colon\Lambda(R)\longrightarrow\Lambda(S)$ and $\Lambda(g)\colon\Lambda(R)\longrightarrow\Lambda(T)$. Furthermore, consider the induced continuous homomorphism of the tensor product rings
$h\colon T\otimes_{R}S\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\Lambda(T)\otimes_{\Lambda(R)}\Lambda(S).$
Then the induced homomorphism of the completions of the tensor products
$\Lambda(h)\colon\Lambda(T\otimes_{R}S)\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\Lambda(\Lambda(T)\otimes_{\Lambda(R)}\Lambda(S)).$
is an isomorphism of topological rings.
Proof.
Put $W=T\otimes_{R}S$. Let $J\subset S$ be an ideal of definition and $K\subset T$ be an ideal of definition. Since the intersection of any two ideals of definition in $R$ is also an ideal of definition in $R$,  Lemma 2.1.3(3) implies existence of an ideal of definition $I\subset R$ such that $f(I)\subset J$ and $g(I)\subset K$. Then, according to Lemma 2.7.1(d),  $L=W(f^{\prime}(K)+g^{\prime}(J))\subset W$ is an ideal of definition in $W$. We have $W/L\simeq T/K\otimes_{R/I}S/J$.
Put $\mathfrak{R}=\Lambda(R)$,  $\mathfrak{S}=\Lambda(S)$,  $\mathfrak{T}=\Lambda(T)$, and $\widetilde{W}=\mathfrak{T}\otimes_{\mathfrak{R}}\mathfrak{S}$ (the latter notation is intended to emphasize the fact that the topological ring $\mathfrak{T}\otimes_{\mathfrak{R}}\mathfrak{S}$ need not be separated or complete). Consider the ideals $\mathfrak{I}=\mathfrak{R}\lambda_{R}(I)\subset\mathfrak{R}$, $\mathfrak{J}=\mathfrak{S}\lambda_{S}(J)\subset\mathfrak{S}$, and $\mathfrak{K}=\mathfrak{T}\lambda_{S}(K)\subset\mathfrak{T}$ corresponding to the ideals $I$, $J$, and $K$ under the bijection described in Section 2.3. Put $\mathfrak{f}=\Lambda(f)$ and $\mathfrak{g}=\Lambda(g)$, and denote by $\mathfrak{f}^{\prime}\colon\mathfrak{T}\longrightarrow\widetilde{W}$ and $\mathfrak{g}^{\prime}\colon\mathfrak{S}\longrightarrow\widetilde{W}$ the induced homomorphisms of topological rings. Then $\widetilde{L}=\widetilde{W}(\mathfrak{f}^{\prime}(\mathfrak{K})+\mathfrak{g}^{\prime}(\mathfrak{J}))\subset\widetilde{W}$ is an ideal of definition in $\widetilde{W}$, and we have $\widetilde{W}/\widetilde{L}\simeq\mathfrak{T}/\mathfrak{K}\otimes_{\mathfrak{R}/\mathfrak{I}}\mathfrak{S}/\mathfrak{J}$.
[…]
Corollary 2.7.5.
[…]
is also an isomorphism of topological rings.
Proof.
Follows from Lemma 2.7.4. ∎
2.8. Formal open immersions and formal open coverings
Let us say that a homomorphism of (discrete) commutative rings $f\colon R\longrightarrow S$ is an open immersion if the induced map of the spectra $\operatorname{Spec}S\longrightarrow\operatorname{Spec}R$ is an open immersion of affine schemes. Equivalently, $f$ is an open immersion if and only if it is a flat ring epimorphism of finite presentation [39, Section XI.2], [10, Theoreme IV.17.9.1]. In particular, we will say that a homomorphism of commutative rings $R\longrightarrow S$ is a principal open immersion if there exists an element $r\in R$ such that $S$ is isomorphic to $R[r^{-1}]$ as a commutative $R$-algebra.
Similarly, let us say that a collection of homomorphisms of commutative rings $f_{\alpha}\colon R\longrightarrow S_{\alpha}$ is a (principal) open covering if every map $f_{\alpha}$ is a (principal) open immersion and the collection of the induced maps of the spectra $\operatorname{Spec}S_{\alpha}\longrightarrow\operatorname{Spec}R$ is an affine open covering of the affine scheme $\operatorname{Spec}R$. Equivalently, the latter condition means that, for any field $k$ and any ring homomorphism $R\longrightarrow k$, there exists an index $\alpha$ such that $S_{\alpha}\otimes_{R}k\neq 0$. The following criterion is a basic result concerning affine open subschemes in affine schemes.
Lemma 2.8.1.
A homomorphism of commutative rings $f\colon R\longrightarrow S$ is an open immersion if and only if there exists a finite collection of elements $r_{1}$, …, $r_{m}\in R$ such that the induced ring maps $R[r_{j}^{-1}]\longrightarrow S[f(r_{j})^{-1}]$ are isomorphisms for all $1\leq j\leq m$ and the collection of principal open immersions $S\longrightarrow S[f(r_{j})^{-1}]$,  $1\leq j\leq m$, is an open covering. ∎
Lemma 2.8.2.
Let $f\colon R\longrightarrow S$ be a homomorphism of commutative rings such that $S$ is a flat $R$-module, and let $I\subset R$ be a nilpotent ideal. Then
(a) $f\colon R\longrightarrow S$ is an isomorphism if and only if $\bar{f}\colon R/I\longrightarrow S/Sf(I)$ is an isomorphism;
(b) $f\colon R\longrightarrow S$ is a principal open immersion if and only if $\bar{f}\colon R/I\longrightarrow S/Sf(I)$ is a principal open immersion;
(c) $f\colon R\longrightarrow S$ is an open immersion if and only if $\bar{f}\colon R/I\longrightarrow S/Sf(I)$ is an open immersion.
Proof.
Part (a): more generally, for any nilpotent ideal $I$ in a commutative ring $R$, a homomorphism of flat $R$-modules $f\colon F\longrightarrow G$ is an isomorphism if and only if the induced map $\bar{f}\colon F/IF\longrightarrow G/IG$ is an isomorphism. This is provable using the Nakayama lemma for nilpotent ideals.
Part (b): the “only if” assertion is obvious. To prove the “if”, let $\bar{r}\in R/I$ be an element such that the $R/I$-algebra $S/Sf(I)$ is isomorphic to $(R/I)[\bar{r}^{-1}]$. Let $r\in R$ be any element such that $\bar{r}=r+I$. Applying part (a) to the $R$-algebra morphisms $R[r^{-1}]\longrightarrow S[f(r)^{-1}]\longleftarrow S$ and the nilpotent ideals $I[r^{-1}]\subset R[r^{-1}]$ and $SI\subset S$, one concludes that both of these $R$-algebra morphisms are isomorphisms.
Part (c): once again, the “only if” assertion is obvious. To prove the “if”, let $\bar{r}_{1}$, …, $\bar{r}_{m}\in R/I$ be a finite collection of elements satisfying the conditions of Lemma 2.8.1 for the ring homomorphism $R/I\longrightarrow S/Sf(I)$. Let $r_{j}\in R$,  $1\leq j\leq m$, be any element such that $\bar{r}_{j}=r_{j}+I$. Applying part (a) to the ring homomorphisms $R[r_{j}^{-1}]\longrightarrow S[f(r_{j})^{-1}]$ and the nilpotent ideals $I[r_{j}^{-1}]\subset R[r_{j}^{-1}]$, one concludes that these ring homomorphisms are isomorphisms. To prove that the collection of principal open immersions $S\longrightarrow S[f(r_{j})^{-1}]$,  $1\leq j\leq m$, is an open covering, it suffices to observe that any homomorphism from $S$ to a field $k$ factorizes through $S/Sf(I)$. ∎
Lemma 2.8.3.
Let $f_{\alpha}\colon R\longrightarrow S_{\alpha}$ be a collection of homomorphisms of commutative rings such that the rings $S_{\alpha}$ are flat $R$-modules, and let $I\subset R$ be a nilpotent ideal. Consider the induced ring homomorphisms $\bar{f}_{\alpha}\colon R/I\longrightarrow S_{\alpha}/S_{\alpha}f_{\alpha}(I)$. Then the collection $(f_{\alpha})$ is an open covering if and only if the collection $(\bar{f}_{\alpha})$ is an open covering.
Proof.
One needs to use Lemma 2.8.2(c). On top of that, the final argument from the proof of the same lemma works: Any ring homomorphism from $R$ to a field $k$ factorizes through $R/I$. ∎
Lemma 2.8.4.
A homomorphism of commutative rings $f\colon R\longrightarrow S$ is an open immersion if and only if there exists an open covering $g_{\alpha}\colon S\longrightarrow T_{\alpha}$ of the ring $S$ such that the compositions $g_{\alpha}f\colon R\longrightarrow T_{\alpha}$ are open immersions. If this is the case, then one can choose $(g_{\alpha})$ to be a principal open covering for which all the compositions $g_{\alpha}f$ are also principal open immersions.
Proof.
The first assertion expresses locality of the notion of an open immersion of schemes. It is immediate from the definitions. The second assertion is a restatement of the fact that principal affine open subschemes form a topology base in $\operatorname{Spec}R$. ∎
Theorem 2.8.5.
Let $R$ be a commutative ring and $I\subset R$ be a nilpotent ideal. Then the functor $S\longmapsto S/Sf(I)$ provides an equivalence between the category of commutative $R$-algebras $f\colon R\longrightarrow S$ such that $f$ is an open immersion and the category of commutative $R/I$-algebras $\bar{f}\colon R/I\longrightarrow\overline{S}$ such that $\bar{f}$ is an open immersion. In other words, the map $\operatorname{Spec}S\longmapsto\operatorname{Spec}S/Sf(I)$ is an isomorphism between the poset of affine open subschemes in $\operatorname{Spec}R$ and the poset of affine open subschemes in $\operatorname{Spec}R/I$ (with respect to the inclusion order). An open immersion $f$ is principal if and only if the open immersion $\bar{f}$ is.
Proof.
We are not aware of any argument within the realm of commutative ring theory. The following proof is, basically, a collection of references. By [10, Theorem IV.18.1.2], the functor $X\longmapsto\operatorname{Spec}R/I\times_{\operatorname{Spec}R}X$ is an equivalence between the category of étale schemes over $\operatorname{Spec}R$ and the category of étale schemes over $\operatorname{Spec}R/I$. By [10, Theorem IV.17.9.1], an étale morphism is an open immersion if and only if it is radicial (in the sense of [11, Section I.3.7]). It follows easily that an étale morphism $X\longrightarrow\operatorname{Spec}R$ is an open immersion if and only if the corresponding morphism $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}X\longrightarrow\operatorname{Spec}R/I$ is an open immersion. So the map $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}{-}$ is an equivalence between the category of open subschemes in $\operatorname{Spec}R$ and the category of open subschemes in $\operatorname{Spec}R/I$. The same conclusion can be simply obtained from the facts that open subschemes in schemes correspond bijectively to open subsets in the underlying topology, and the underlying topological spaces of $\operatorname{Spec}R$ and $\operatorname{Spec}R/I$ coincide.
The key step is to prove that an open subscheme $Y\subset\operatorname{Spec}R$ is affine if and only if the scheme $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}Y$ is affine. The “only if” assertion is obvious. To prove the “if”, notice that the underlying topological spaces of $Y$ and $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}Y$ coincide. Quasi-compactness and quasi-separatedness are properties of the underlying topological space of a scheme; and all affine schemes are quasi-compact and quasi-separated. So the scheme $Y$ is quasi-compact and quasi-separated (in fact, separated, as an open subscheme of an affine scheme). By [8, Théorème II.5.2.1(d)], [9, IV.1.7.17], [13, Lemmas Tags 01XB and 01XF or 01XG], a quasi-compact scheme $Y$ is affine if and only if $H^{1}(Y,\mathcal{F})=0$ for every quasi-coherent sheaf $\mathcal{F}$ on $Y$. Here $H^{*}(Y,{-})$ denotes the cohomology groups of sheaves of abelian groups on $Y$, which only depend on the underlying topological space of $Y$. It remains to point out that every quasi-coherent sheaf $\mathcal{F}$ on $Y$ has a finite decreasing filtration by quasi-coherent subsheaves $\mathcal{F}\supset I\mathcal{F}\supset I^{2}\mathcal{F}\supset\dotsb\supset I^{n}\mathcal{F}=0$, where $n\geq 1$ is an integer such that $I^{n}=0$ in $R$, and the successive quotient sheaves $I^{m}\mathcal{F}/I^{m+1}\mathcal{F}$ are quasi-coherent sheaves on $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}Y$.
Finally, if $f\colon R\longrightarrow S$ is an open immersion and the $R/I$-algebra $S/Sf(I)$ is isomorphic to $(R/I)[\bar{r}^{-1}]$ for some element $\bar{r}\in R/I$, then the $R$-algebra $S$ is isomorphic to $R[r^{-1}]$ for any element $r\in R$ such that $\bar{r}=r+I$. This assertion follows immediately from the bijection estalished above. ∎
Lemma 2.8.6.
Let $f\colon R\longrightarrow S$ be a tight continuous map of adic topological rings. Then the following three conditions are equivalent:
  1. (1)
    for every ideal of definition $I\subset R$, the induced ring map $\bar{f}\colon R/I\longrightarrow S/Sf(I)$ is an open immersion;
  2. (2)
    there exists an ideal of definition $I\subset R$ such that the induced ring map $\bar{f}\colon R/I^{n}\longrightarrow S/Sf(I^{n})$ is an open immersion for every $n\geq 1$;
  3. (3)
    $f$ is flat (as a map of adic topological rings) and there exists an ideal of definition $I\subset R$ such that the induced ring map $\bar{f}\colon R/I\longrightarrow S/Sf(I)$ is an open immersion.
Proof.
(1) $\Longrightarrow$ (2) Holds because $I^{n}\subset R$ is an ideal of definition for any ideal of definition $I\subset R$.
(2) $\Longrightarrow$ (1) The point is that if $I\subset I^{\prime}\subset R$ are two ideals and the ring map $R/I\longrightarrow S/Sf(I)$ is an open immersion, then the ring map $R/I^{\prime}\longrightarrow S/Sf(I^{\prime})$ is an open immersion, too.
(1) $\Longrightarrow$ (3) Holds because, for all open immersions of commutative rings $\bar{f}\colon\overline{R}\longrightarrow\overline{S}$, the ring $\overline{S}$ is a flat $\overline{R}$-module.
(3) $\Longrightarrow$ (2) Follows from Lemma 2.8.2(c). ∎
We will say that a tight continuous map of adic topological rings $f\colon R\longrightarrow S$ is a formal open immersion if it satisfies the equivalent conditions of Lemma 2.8.6. Using the criterion of Lemma 2.8.6(1), one can easily check that the composition of any two formal open immersions is a formal open immersion.
Lemma 2.8.7.
Let $f\colon R\longrightarrow S$ be a continuous map of adic topological rings. Consider the ring $S\otimes_{R}S$, and endow it with the tensor product topology. In this context:
(a) the induced map of topological rings $S\otimes_{R}S\longrightarrow S$ is continuous and tight;
(b) the induced maps of topological rings $S\rightrightarrows S\otimes_{R}S$ are continuous;
(c) if the map $f$ is tight, then so are the maps $S\rightrightarrows S\otimes_{R}S$;
(d) if the map $f$ is a formal open immersion, then all the three induced maps of the completions
$\Lambda(S)\rightrightarrows\Lambda(S\otimes_{R}S)\longrightarrow\Lambda(S)$
are isomorphisms of topological rings.
Proof.
Parts (a) and (b) do not depend on the topology on $R$ (one can use the discrete topology on $R$, for example). Part (a): denote the map in question by $p\colon S\otimes_{R}S\longrightarrow S$. Let $J\subset S$ be an ideal of definition. Then, by Lemma 2.7.1(d),  $J\otimes_{R}S+S\otimes_{R}J$ is an ideal of definition in $S\otimes_{R}S$. Now we have $J=p(J\otimes_{R}S+S\otimes_{R}J)$, so $p$ is a tight continuous map by Lemma 2.1.6(4).
Part (b) follows from the discussion in Section 2.7 (including Lemma 2.7.2(a)).
Part (c): denote the maps in question by $i_{1}$ and $i_{2}\colon S\rightrightarrows S\otimes_{R}S$. Let $I\subset R$ be an ideal of definition. By Lemma 2.1.6(5), the ideal $J=Sf(I)\subset S$ is an ideal of definition in $S$. By Lemma 2.7.1(d), the ideal $J\otimes_{R}S+S\otimes_{R}J\subset S\otimes_{R}S$ is an ideal of definition in $S\otimes_{R}S$. Now we have $Sf(I)\otimes_{R}S=S\otimes_{R}Sf(I)$, hence $J\otimes_{R}S+S\otimes_{R}J=J\otimes_{R}S=S\otimes_{R}J\subset S\otimes_{R}S$. By Lemma 2.1.6(4), it follows that $i_{1}$ and $i_{2}$ are tight ring maps.
Part (d): in the notation of the previous paragraph, let $I$ range over all the ideals of definition in $R$. Then the ideals $J=Sf(I)$ form a base of neighborhoods of zero in $S$, while the ideals $J\otimes_{R}S=S\otimes_{R}J$ form a base of neighborhoods of zero in $S\otimes_{R}S$. Now we have $S/(J\otimes_{R}S+S\otimes_{R}J)\simeq S/J\otimes_{R/I}S/J$. The natural maps $S/J\rightrightarrows S/J\otimes_{R/I}S/J\longrightarrow S/J$ are ring isomorphisms, since $R/I\longrightarrow S/J$ is an open immersion (of abstract commutative rings) by Lemma 2.8.6(1). Passing to the projective limit over $I$, we obtain the desired isomorphisms of complete adic topological rings. ∎
Lemma 2.8.8.
Let $f\colon R\longrightarrow S$ be a tight continuous map of adic topological rings. Then the following four conditions are equivalent:
  1. (1)
    for every ideal of definition $I\subset R$, the induced ring map $\bar{f}\colon R/I\longrightarrow S/Sf(I)$ is a principal open immersion;
  2. (2)
    there exists an ideal of definition $I\subset R$ such that the induced ring map $\bar{f}\colon R/I^{n}\longrightarrow S/Sf(I^{n})$ is a principal open immersion for every $n\geq 1$;
  3. (3)
    $f$ is flat (as a map of adic topological rings) and there exists an ideal of definition $I\subset R$ such that the induced ring map $\bar{f}\colon R/I\longrightarrow S/Sf(I)$ is a principal open immersion;
  4. (4)
    there exists an element $r\in R$ such that, endowing the ring $R[r^{-1}]$ with the adic topology for which the natural ring map $l\colon R\longrightarrow R[r^{-1}]$ is continuous and tight, there exists a unique isomorphism of topological rings $\Lambda(S)\simeq\Lambda(R[r^{-1}])$ making the following diagram commutative:
[…]
Proof.
(1) $\Longrightarrow$ (2) and (1) $\Longrightarrow$ (3) Similar to the proof of Lemma 2.8.6.
(2) $\Longrightarrow$ (1) The point is that if $I\subset I^{\prime}\subset R$ are two ideals and the ring map $R/I\longrightarrow S/Sf(I)$ is a principal open immersion, then the ring map $R/I^{\prime}\longrightarrow S/Sf(I^{\prime})$ is a principal open immersion, too.
(3) $\Longrightarrow$ (2) Follows from Lemma 2.8.2(b).
(4) $\Longrightarrow$ (1) Obvious.
(4) $\Longrightarrow$ (3) Follows from Lemma 2.3.1(b).
(3) $\Longrightarrow$ (4) Let $\bar{r}\in R/I$ be an element such that the $R/I$-algebra $S/Sf(I)$ is isomorphic to $(R/I)[\bar{r}^{-1}]$. Pick any element $r\in R$ for which $\bar{r}=r+I$. Let $n\geq 1$ be an integer; put $\tilde{r}=r+I^{n}\in R/I^{n}$. By the proof of Lemma 2.8.2(b), the $R/I^{n}$-algebra $S/Sf(I^{n})$ is isomorphic to $(R/I^{n})[\tilde{r}^{-1}]$. As the category of open subschemes in $\operatorname{Spec}R$ is a poset, an isomorphism of $R/I^{n}$-algebras $(R/I^{n})[\tilde{r}^{-1}]\simeq S/Sf(I^{n})$ is unique. Hence, as $n$ varies, such isomorphisms form a projective system. Passing to the projective limit over $n\geq 1$, we obtain the desired isomorphism of adic topological rings and $\Lambda(R)$-algebras $\Lambda(R[r^{-1}])\simeq\Lambda(S)$. ∎
We will say that a tight continuous map of adic topological rings $f\colon R\longrightarrow S$ is a principal formal open immersion if it satisfies the equivalent conditions of Lemma 2.8.8. Using the criterion of Lemma 2.8.8(1), one can easily check that the composition of any two principal formal open immersions is a principal formal open immersion.
Corollary 2.8.9.
Let $R$ be an adic topological ring and $I\subset R$ be an ideal of definition. Then the functor $S\longmapsto S/Sf(I)$ provides an equivalence between the category of commutative $R$-algebras $f\colon R\longrightarrow S$ with an adic topology on $S$ such that $f$ is a formal open immersion and the category of commutative $R/I$-algebras $\bar{f}\colon R/I\longrightarrow\overline{S}$ such that $\bar{f}$ is an open immersion. A formal open immersion $f$ is principal if and only if the open immersion $\bar{f}$ is principal.
Proof.
[…]
Lemma 2.8.10.
Let $f_{\alpha}\colon R\longrightarrow S_{\alpha}$ be a collection of tight continuous maps of adic topological rings. Then the following three conditions are equivalent:
  1. (1)
    for every ideal of definition $I\subset R$, the collection of induced ring maps $\bar{f}_{\alpha}\colon R/I\longrightarrow S_{\alpha}/S_{\alpha}f_{\alpha}(I)$ is an open covering;
  2. (2)
    there exists an ideal of definition $I\subset R$ such that the collection of induced ring maps $\bar{f}_{\alpha}\colon R/I^{n}\longrightarrow S_{\alpha}/S_{\alpha}f_{\alpha}(I^{n})$ is an open covering for every $n\geq 1$;
  3. (3)
    the maps $f_{\alpha}$ are flat (as maps of adic topological rings) and there exists an ideal of definition $I\subset R$ such that the collection of induced ring maps $\bar{f}_{\alpha}\colon R/I\longrightarrow S_{\alpha}/S_{\alpha}f_{\alpha}(I)$ is an open covering.
Proof.
Follows from Lemmas 2.8.3 and 2.8.6. ∎
We will say that a collection of tight continuous maps of adic topological rings $f_{\alpha}\colon R\longrightarrow S_{\alpha}$ is a formal open covering if it satisfies the equivalent conditions of Lemma 2.8.10. Clearly, any formal open covering of an adic topological ring has a finite subcovering. A principal formal open covering is a formal open covering $(f_{\alpha})$ such that every map $f_{\alpha}$ is a principal formal open immersion.
Lemma 2.8.11.
A tight continuous map of adic topological rings $f\colon R\longrightarrow S$ is an formal open immersion if and only if there exists a formal open covering $g_{\alpha}\colon S\longrightarrow T_{\alpha}$ of the adic topological ring $S$ such that the compositions $g_{\alpha}f\colon R\longrightarrow T_{\alpha}$ are formal open immersions. If this is the case, then one can choose $(g_{\alpha})$ to be a principal formal open covering for which all the compositions $g_{\alpha}f$ are also principal formal open immersions.
Proof.
First assertion: to prove the “only if”, it suffices to take the set of indices $\{\alpha\}$ to be the singleton $\{0\}$, and the trivial formal open covering with $T_{0}=S$ and $g_{0}=\mathrm{id}_{S}$. To prove the “if”, let $I\subset R$ be an ideal of definition. By Lemma 2.1.6(5), the ideal $Sf(I)\subset S$ is an ideal of definition in $S$, while the ideals $T_{\alpha}g_{\alpha}f(I)\subset T_{\alpha}$ are ideals of definition in $T_{\alpha}$. By Lemma 2.8.10(1), the collection of induced ring maps $\bar{g}_{\alpha}\colon S/Sf(I)\longrightarrow T_{\alpha}/T_{\alpha}g_{\alpha}f(I)$ is an open covering. By Lemma 2.8.6(1), the compositions $\bar{g}_{\alpha}\bar{f}\colon R/I\longrightarrow T_{\alpha}/T_{\alpha}g_{\alpha}f(I)$ are open immersions. By the first assertion of Lemma 2.8.4, it follows that the induced ring map $\bar{f}\colon R/I\longrightarrow S/Sf(I)$ is an open immersion. Using Lemma 2.8.6(1) again, we conclude that $f\colon R\longrightarrow S$ is a formal open immersion.
Second assertion: let $I\subset R$ be an ideal of definition. By Lemma 2.8.6(1), the induced ring map $\bar{f}\colon R/I\longrightarrow S/Sf(I)$ is an open immersion. By the sectond assertion of Lemma 2.8.4, there exists a principal open covering $\bar{g}_{\alpha}\colon S/Sf(I)\longrightarrow\overline{T}_{\alpha}$ such that the compositions $\bar{g}_{\alpha}\bar{f}\colon R/I\longrightarrow\overline{T}_{\alpha}$ are principal open immersions. By Corollary 2.8.9, there exist principal formal open immersions $g_{\alpha}\colon S\longrightarrow T_{\alpha}$ together with $S/Sf(I)$-algebra isomorphisms $T_{\alpha}/T_{\alpha}g_{\alpha}f(I)\simeq\overline{T}_{\alpha}$. By Lemma 2.8.10(1), the collection $(g_{\alpha})$ is a formal open covering. ∎
Lemma 2.8.12.
Let $f_{\alpha}\colon R\longrightarrow S_{\alpha}$ be a formal open covering of an adic topological ring $R$. Then a quotseparated contramodule $R$-module $\mathfrak{F}$ is flat if and only if the quotseparated contramodule $S_{\alpha}$-module $f_{\alpha}^{\sharp}(\mathfrak{F})$ is flat for every index $\alpha$. A contramodule $R$-module $\mathfrak{F}$ is flat if and only if the contramodule $S_{\alpha}$-module $f_{\alpha}^{\#}(\mathfrak{F})$ is flat for every index $\alpha$.
Proof.
Let us prove the quotseparated version. Let $I\subset R$ be an ideal of definition in $R$. Then $J_{\alpha}=S_{\alpha}f_{\alpha}(I)\subset S_{\alpha}$ is an ideal of definition in $S_{\alpha}$ for every index $\alpha$. By Lemma 2.8.10(1), the collection of induced ring maps $\bar{f}_{\alpha}\colon R/I\longrightarrow S_{\alpha}/J_{\alpha}$ is an open covering. By Lemma 2.6.3(c), we have $f_{\alpha}^{\sharp}(\mathfrak{F})/Jf_{\alpha}^{\sharp}(\mathfrak{F})\simeq S_{\alpha}/J_{\alpha}\otimes_{R/I}\mathfrak{F}/I\mathfrak{F}$. Now it remains to point out that the $R/I$-module $\mathfrak{F}/I\mathfrak{F}$ is flat if and only if the $S_{\alpha}/J_{\alpha}$-module $S_{\alpha}/J_{\alpha}\otimes_{R/I}\mathfrak{F}/I\mathfrak{F}$ is flat for every index $\alpha$. The proof of the nonquotseparated case is similar and based on Lemma 2.6.3(b). ∎
Lemma 2.8.13.
Let $R$, $S$, and $T$ be adic topological rings, let $f\colon R\longrightarrow S$ be a tight continuous ring map, and let $g\colon R\longrightarrow T$ be a continuous ring map. Put $W=T\otimes_{R}S$, and denote by $f^{\prime}\colon T\longrightarrow W$ and $g^{\prime}\colon S\longrightarrow W$ the induced ring maps. Endow the commutative ring $W$ with the tensor product topology, as in Section 2.7. In this setting:
(a) if the tight continuous ring map $f$ is a formal open immersion, then the tight continuous ring map $f^{\prime}$ is a formal open immersion, too;
(b) if the tight continuous ring map $f$ is a principal formal open immersion, then the tight continuous ring map $f^{\prime}$ is a principal formal open immersion, too.
Proof.
Similar to the proof of Lemma 2.7.3(b). ∎
2.9. Very flat and contraadjusted contramodules
In the sequel, we will use a shorthand notation $R{\operatorname{\mathsf{--Contra}}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ for any adic topological ring $R$. In the case of a complete, separated adic topological ring $\mathfrak{R}$, a contramodule over $\mathfrak{R}$ as a topological ring is essentially the same thing as a quotseparated contramodule $\mathfrak{R}$-module (see the end of Section 2.4), so this notation is unambiguous. Let us also recall a similar notation $R{\operatorname{\mathsf{--Tors}}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$ from Section 2.2.
We start with a recollection of the case of an abstract commutative ring $R$ (without any adic or other topology). An $R$-module $P$ is said to be contraadjusted [17, Section 1.1], [38, Section 5], [20, Sections 2 and 8], [34, Section 0.5], [26, Section 2 and Example 3.2], [28, Section 4.3] if $\operatorname{Ext}^{1}_{R}(R[s^{-1}],P)=0$ for all elements $s\in R$. An $R$-module $F$ is said to be very flat [17, Section 1.1], [38, Section 2], [20, Section 8], [34, Section 0.5], [26, Section 2 and Example 2.5], [28, Section 4.3] if $\operatorname{Ext}^{1}_{R}(F,C)=0$ for all contraadjusted $R$-modules $C$.
The full subcategory of contraadjusted $R$-modules is closed under extensions, quotients, and infinite products in $R{\operatorname{\mathsf{--Mod}}}$. The full subcategory of very flat $R$-modules is closed under extensions, kernels of epimorphisms, infinite direct sums, and tensor products in $R{\operatorname{\mathsf{--Mod}}}$. All very flat $R$-modules have projective dimensions at most $1$.
A homomorphism of commutative rings $f\colon R\longrightarrow S$ is said to be very flat if, for every element $s\in S$, the $R$-module $S[s^{-1}]$ is very flat. If this is the case, the ring $S$ is also said to be a very flat commutative $R$-algebra. Very flatness of the $R$-module $S$ is not sufficient for the ring homomorphism $R\longrightarrow S$ to be very flat [34, Example 9.7], [28, Section 6.1]. The composition of any two very flat homomorphisms of commutative rings is very flat [17, Lemma 1.2.3(b)]. Every open immersion of commutative rings is very flat [17, Lemma 1.2.4].
Lemma 2.9.1.
Let $R$ be an adic topological ring and $\mathfrak{F}$ be a contramodule $R$-module. Then the following two conditions are equivalent:
  1. (1)
    $\mathfrak{F}$ is a flat contramodule $R$-module and there exists an ideal of definition $I\subset R$ for which $\mathfrak{F}/I\mathfrak{F}$ is a very flat $R/IR$-module;
  2. (2)
    for every open ideal $I\subset R$, the $R/I$-module $\mathfrak{F}/I\mathfrak{F}$ is very flat.
Proof.
(2) $\Longrightarrow$ (1) holds because all very flat $R/I$-modules are flat.
(1) $\Longrightarrow$ (2) Let $I\subset I^{\prime}\subset R$ be two ideals of definition in $R$. Then $I^{\prime}/I$ is a finitely generated nilpotent ideal in the ring $R/I$. Therefore, assuming that the $R/I$-module $\mathfrak{F}/I\mathfrak{F}$ is flat, the $R/I$-module $\mathfrak{F}/I\mathfrak{F}$ is very flat if and only if the $R/I^{\prime}$-module $\mathfrak{F}/I^{\prime}\mathfrak{F}$ is very flat [17, Lemma 1.7.10(b)]. Now if $I\subset I^{\prime}\subset R$ are two ideals such that $I$ is an ideal of definition, and the $R/I$-module $\mathfrak{F}/I\mathfrak{F}$ is very flat, then the $R/I^{\prime}$-module $\mathfrak{F}/I^{\prime}\mathfrak{F}$ is very flat by [17, Lemma 1.2.2(b)]. ∎
Let $R$ be an adic topological ring. A contramodule $R$-module $\mathfrak{F}$ is said to be very flat if it satisfies the equivalent conditions of Lemma 2.9.1. All projective contramodule $R$-modules and all projective quotseparated contramodule $R$-modules are very flat (see the proof of Lemma 2.5.5 (2) $\Rightarrow$ (3)). It follows from Lemma 2.5.9 that the full subcategory of very flat quotseparated contramodule $R$-modules is closed under extensions and kernels of epimorphisms in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$. Furthermore, in view of [17, Corollary E.1.10(a)], all very flat quotseparated contramodule $R$-modules have projective dimensions at most $1$ as objects of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$. These assertions are special cases of [17, Corollary E.4.1].
Lemma 2.9.2.
Let $R$ be an adic topological ring and $F$ be a flat $R$-module. Then, for any quotseparated contramodule $R$-module $\mathfrak{P}$, there is a natural $\operatorname{Ext}^{1}$-adjunction isomorphism of abelian groups/$R$-modules
$\operatorname{Ext}^{1}_{R}(F,\mathfrak{P})\simeq\operatorname{Ext}^{1}_{R{\operatorname{\mathsf{--Contra}}}}(\Lambda(F),\mathfrak{P}).$
Proof.
By [14, Lemma 3.5 or Proposition 3.6], we have $\mathbb{L}_{i}\Lambda(F)=0$ for all $i>0$. Therefore, for any short exact sequence of $R$-modules $0\longrightarrow M\longrightarrow N\longrightarrow F\longrightarrow 0$, the short sequence of quotseparated contramodule $R$-modules $0\longrightarrow\mathbb{L}_{0}\Lambda(M)\longrightarrow\mathbb{L}_{0}\Lambda(N)\longrightarrow\mathbb{L}_{0}\Lambda(F)\longrightarrow 0$ is exact.
Now we have a pair of adjoint functors between abelian categories
${R{\operatorname{\mathsf{--Mod}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R{\operatorname{\mathsf{--Contra}}},\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbb{L}_{0}\Lambda}$
where the right adjoint functor $R{\operatorname{\mathsf{--Contra}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is the identity inclusion. The right adjoint functor is exact, while the object $F\in R{\operatorname{\mathsf{--Mod}}}$ is adjusted to the left adjoint functor $\mathbb{L}_{0}\Lambda$ in the sense stated in the previous paragraph. By [26, Lemma 1.7(e)], the desired $\operatorname{Ext}^{1}$-adjunction isomorphism follows. Here we are also using the natural $R$-module isomorphism $\mathbb{L}_{0}\Lambda(F)\simeq\Lambda(F)$, which holds by the same results from [14] or by Corollary 2.5.7 above. ∎
Lemma 2.9.3.
Let $R$ be an adic topological ring, $\mathfrak{R}$ be the adic completion of $R$, and $\mathfrak{P}$ be a quotseparated contramodule $R$-module. Then the following seven conditions are equivalent:
  1. (1)
    the functor $\operatorname{Hom}_{R}({-},\mathfrak{P})$ takes short exact sequences of very flat quotseparated contramodule $R$-modules to short exact sequences of abelian groups/$R$-modules/quotseparated contramodule $R$-modules;
  2. (2)
    one has $\operatorname{Ext}^{1}_{R{\operatorname{\mathsf{--Contra}}}}(\mathfrak{F},\mathfrak{P})=0$ for all very flat quotseparated contramodule $R$-modules $\mathfrak{F}$ (where the Ext group is computed in the abelian category $R{\operatorname{\mathsf{--Contra}}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$);
  3. (3)
    one has $\operatorname{Ext}^{1}_{R{\operatorname{\mathsf{--Contra}}}}(\Lambda(R[r^{-1}]),\mathfrak{P})=0$ for all elements $r\in R$;
  4. (4)
    $\mathfrak{P}$ is a contraadjusted $R$-module;
  5. (5)
    $\mathfrak{P}$ is a contraadjusted $\mathfrak{R}$-module;
  6. (6)
    for every ideal $I\subset R$, the $R/I$-module $\mathfrak{P}/I\mathfrak{P}$ is contraadjusted;
  7. (7)
    for every ideal $I\subset\mathfrak{R}$, the $\mathfrak{R}/I$-module $\mathfrak{P}/I\mathfrak{P}$ is contraadjusted.
Furthermore, if $\mathfrak{P}$ is a separated contramodule $R$-module, then conditions (1–7) are also equivalent to the following two conditions:
  1. (8)
    there exists an ideal of definition $I\subset R$ for which $\mathfrak{P}/I\mathfrak{P}$ is a contraadjusted $R/I$-module;
  2. (9)
    for every open ideal $I\subset R$, the $R/I$-module $\mathfrak{P}/I\mathfrak{P}$ is contraadjusted.
Proof.
(1) $\Longleftrightarrow$ (2) Follows from the facts that there are enough very flat quotseparated contramodules in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ and the full subcategory of very flat quotseparated contramodules is closed under kernels of epimorphisms.
(3) $\Longleftrightarrow$ (4) Follows immediately from Lemma 2.9.2 (applied to the flat $R$-module $F=R[r^{-1}]$).
(1) $\Longleftrightarrow$ (4) This is a special case of [17, first assertion of Corollary E.4.8]. In the case of a Noetherian ring $R$, one can also refer to [17, Corollary D.3.5(c,e)].
(1) $\Longleftrightarrow$ (5) This is a particular case of (1) $\Longleftrightarrow$ (4).
(6) $\Longrightarrow$ (4) or (7) $\Longrightarrow$ (5) Take $I=0$.
(4) $\Longrightarrow$ (6) or (5) $\Longrightarrow$ (7) This is [17, Lemma 1.7.7(b)].
(6) $\Longrightarrow$ (9) and (7) $\Longrightarrow$ (9) $\Longrightarrow$ (8) Obvious.
(8) $\Longrightarrow$ (9) This is [17, Lemma 1.7.10(a)].
(9) $\Longrightarrow$ (2) This is a special case of [17, Lemma E.4.3]. ∎
A quotseparated contramodule $R$-module $\mathfrak{P}$ is said to be contraadjusted if it satisfies any one of the equivalent conditions (1–7) of Lemma 2.9.3. The full subcategory of contraadjusted quotseparated contramodule $R$-modules is closed under extensions, quotients, and infinite products in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$.
Proposition 2.9.4.
Let $R$ be an adic topological ring and $\mathfrak{M}$ be a quotseparated contramodule $R$-module. Then
(a) there exists a short exact sequence of quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{F}\longrightarrow\mathfrak{M}\longrightarrow 0$ with a very flat quotseparated contramodule $R$-module $\mathfrak{F}$ and a contraadjusted quotseparated contramodule $R$-module $\mathfrak{P}$;
(b) there exists a short exact sequence of quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{F}\longrightarrow 0$ with a contraadjusted quotseparated contramodule $R$-module $\mathfrak{P}$ and a very flat quotseparated contramodule $R$-module $\mathfrak{F}$.
Proof.
Part (a) is a special case of [17, Corollary E.4.5]. For a Noetherian ring $R$, one can also refer to [17, Corollary D.3.5(b,e)]. Part (b) is a special case of [17, Corollary E.4.6]. For a Noetherian ring $R$, one can also refer to [17, Corollary D.3.5(a,e)]. Another reference is [33, Example 7.12(3)]. ∎
Corollary 2.9.5.
Let $R$ be an adic topological ring and $\mathfrak{F}$ be a quotseparated contramodule $R$-module. Then the following conditions are equivalent:
  1. (1)
    the functor $\operatorname{Hom}_{R}(\mathfrak{F},{-})$ takes short exact sequences of contraadjusted quotseparated contramodule $R$-modules to short exact sequences of abelian groups/$R$-modules/quotseparated contramodule $R$-modules;
  2. (2)
    one has $\operatorname{Ext}^{1}_{R{\operatorname{\mathsf{--Contra}}}}(\mathfrak{F},\mathfrak{P})=0$ for all contraadjusted quotseparated contramodule $R$-modules $\mathfrak{P}$;
  3. (3)
    $\mathfrak{F}$ is a very flat quotseparated contramodule $R$-module.
Proof.
(1) $\Longleftrightarrow$ (2) Follows from the facts that there are enough contraadjusted quotseparated contramodules in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ (by Proposition 2.9.4(b)) and the full subcategory of contraadjusted quotseparated contramodules is closed under cokernels of monomorphisms.
(3) $\Longrightarrow$ (2) Holds by Lemma 2.9.3(2).
(2) $\Longrightarrow$ (3) This a standard argument. By Proposition 2.9.4(a), there exists a short exact sequence $0\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{G}\longrightarrow\mathfrak{F}\longrightarrow 0$ in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ such that $\mathfrak{G}$ is very flat and $\mathfrak{P}$ is contraadjusted. By (2), it follows that $\mathfrak{F}$ is a direct summand of $\mathfrak{G}$. It remains to make the obvious observation that the class of very flat quotseparated contramodule $R$-modules is closed under direct summands. (For a Noetherian ring $R$, the reference to [17, Corollary D.3.5(d,e)] is also applicable.) ∎
The following lemma is the very flat counterpart of Corollary 2.6.4.
Lemma 2.9.6.
Let $f\colon R\longrightarrow S$ be a continuous homomorphism of adic topological rings. Then, for any very flat quotseparated contramodule $R$-module $\mathfrak{F}$, the quotseparated contramodule $S$-module $f^{\sharp}(\mathfrak{F})$ is very flat. For any very flat contramodule $R$-module $\mathfrak{F}$, the contramodule $S$-module $f^{\#}(\mathfrak{F})$ is very flat.
Proof.
The argument is similar to the proof of Corollary 2.6.4 and based on Lemma 2.6.3(b–c). The point is that, for any homomorphism of (abstract, nontopological) commutative rings $\bar{f}\colon\overline{R}\longrightarrow\overline{S}$, the functor $\bar{f}^{*}\colon\overline{R}{\operatorname{\mathsf{--Mod}}}\longrightarrow\overline{S}{\operatorname{\mathsf{--Mod}}}$ takes very flat $R$-modules to very flat $S$-modules [17, Lemma 1.2.2(b)]. ∎
Lemma 2.9.7.
Let $f\colon R\longrightarrow S$ be a continuous homomorphism of adic topological rings. Then, for any contraadjusted quotseparated contramodule $S$-module $\mathfrak{P}$, the quotseparated contramodule $R$-module $f_{\sharp}(\mathfrak{P})$ is contraadjusted.
Proof.
Using the criterion of Lemma 2.9.3(4), the question reduces to the case of a homomorphism of abstract (nontopological) commutative rings, where we can refer to [17, Lemma 1.2.2(a)]. Alternatively, an argument proceeding entirely within the realm of quotseparated contramodules is possible, using the criterion of Lemma 2.9.3(2) and the results of Corollary 2.6.5 and Lemma 2.9.6, and similar to the proof of Lemma 2.10.4 sketched below. ∎
The next lemma is a very flat complement to Lemma 2.1.8.
Lemma 2.9.8.
Let $f\colon R\longrightarrow S$ be a tight continuous map of adic topological rings. Then the following three conditions are equivalent:
  1. (1)
    for every ideal of definition $I\subset R$, the induced ring map $\bar{f}\colon R/I\longrightarrow S/Sf(I)$ is very flat;
  2. (2)
    there exists an ideal of definition $I\subset R$ such that the induced ring map $\bar{f}\colon R/I^{n}\longrightarrow S/Sf(I^{n})$ is very flat for every $n\geq 1$;
  3. (3)
    $f$ is flat (as a map of adic topological rings) and there exists an ideal of definition $I\subset R$ such that the induced ring map $\bar{f}\colon R/I\longrightarrow S/Sf(I)$ is very flat.
Proof.
All the assertions follow from [17, Lemma 1.7.10(b)]. ∎
We will say that a tight continuous homomorphism of adic topological rings $f\colon R\longrightarrow S$ is very flat if it satisfies the equivalent conditions of Lemma 2.9.8. Any formal open immersion of adic topological rings in very flat. Using the criterion of Lemma 2.9.8(1), one can easily check that the composition of any two very flat tight continuous ring maps of adic topological rings is very flat.
The following lemma is the very flat version of Lemma 2.6.9(c–d).
Lemma 2.9.9.
Let $f\colon R\longrightarrow S$ be a very flat tight continuous map of adic topological rings. Then
(a) the functor $f_{\#}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ takes very flat contramodule $S$-modules to very flat contramodule $R$-modules;
(b) the functor $f_{\sharp}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ takes very flat quotseparated contramodule $S$-modules to very flat quotseparated contramodule $R$-modules.
Proof.
The argument is similar to the proof of Lemma 2.6.9(c–d). The point is that, for any very flat homomorphism of (abstract, nontopological) commutative rings $\bar{f}\colon\overline{R}\longrightarrow\overline{S}$, the functor $\bar{f}_{*}\colon\overline{S}{\operatorname{\mathsf{--Mod}}}\longrightarrow\overline{R}{\operatorname{\mathsf{--Mod}}}$ takes very flat $S$-modules to very flat $R$-modules [17, Lemma 1.2.3(b)], [34, Lemma 9.3(a)], [28, Lemma 6.1]. ∎
The next lemma is the very flat version of Lemma 2.8.12.
Lemma 2.9.10.
Let $f_{\alpha}\colon R\longrightarrow S_{\alpha}$ be a formal open covering of an adic topological ring $R$ (in the sense of Section 2.8). Then a quotseparated contramodule $R$-module $\mathfrak{F}$ is very flat if and only if the quotseparated contramodule $S_{\alpha}$-module $f_{\alpha}^{\sharp}(\mathfrak{F})$ is very flat for every index $\alpha$. A contramodule $R$-module $\mathfrak{F}$ is very flat if and only if the contramodule $S_{\alpha}$-module $f_{\alpha}^{\#}(\mathfrak{F})$ is very flat for every index $\alpha$.
Proof.
The proof is similar to that of Lemma 2.8.12 and based on [17, Lemma 1.2.6(a)] together with Lemma 2.6.3(b–c) above. ∎
Lemma 2.9.11.
Let $R$ be an adic topological ring with an ideal of definition $I\subset R$, and let $F$ be an $R$-module such that the $R/I^{n}$-module $F/I^{n}F$ is very flat for all $n\geq 1$. Let $0\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}\longrightarrow 0$ be a short exact sequence of quotseparated contramodule $R$-modules with a contraadjusted quotseparated contramodule $R$-module $\mathfrak{P}$. Then the short sequence of quotseparated contramodule $R$-modules $0\longrightarrow\operatorname{Hom}_{R}(F,\mathfrak{P})\longrightarrow\operatorname{Hom}_{R}(F,\mathfrak{M})\longrightarrow\operatorname{Hom}_{R}(F,\mathfrak{N})\longrightarrow 0$ is exact.
Proof.
First of all, the Hom $R$-modules in question are quotseparated contramodules by Lemma 2.4.2(c). In the special case when $F$ is a flat $R$-module, the exactness of the short sequence in question follows from the fact that $\operatorname{Ext}_{R}^{1}(F,\mathfrak{P})=0$, see [17, Corollary E.4.8]. In the general case, we notice that $\operatorname{Hom}_{R}(F,\mathfrak{N})\simeq\operatorname{Hom}_{R}(\Lambda(F),\mathfrak{N})$ and similarly for $\mathfrak{M}$ and $\mathfrak{P}$, by Corollary 2.5.7. By Lemma 2.5.3,  $\Lambda(F)$ is a very flat (quot)separated contramodule $R$-module under our current assumptions. Now the exactness of the short sequence of $\operatorname{Hom}_{R}(F,{-})$ follows from the fact that $\operatorname{Ext}_{R{\operatorname{\mathsf{--Contra}}}}^{1}(\Lambda(F),\mathfrak{P})=0$ by Lemma 2.9.3(2). ∎
2.10. Cotorsion contramodules
Once again, we start with a reminder of the case of an abstract associative ring $R$. A left $R$-module $P$ is said to be cotorsion (in the sense of Enochs [5]) if $\operatorname{Ext}^{1}_{R}(F,P)=0$ for all flat left $R$-modules $P$. The full subcategory of cotorsion $R$-modules is closed under extensions, cokernels of monomorphisms, and infinite products in $R{\operatorname{\mathsf{--Mod}}}$.
Lemma 2.10.1.
Let $R$ be an adic topological ring and $\mathfrak{P}$ be a quotseparated contramodule $R$-module. Then the following three conditions are equivalent:
  1. (1)
    the functor $\operatorname{Hom}_{R}({-},\mathfrak{P})$ takes short exact sequences of flat quotseparated contramodule $R$-modules to short exact sequences of abelian groups/$R$-modules/quotseparated contramodule $R$-modules;
  2. (2)
    one has $\operatorname{Ext}^{1}_{R{\operatorname{\mathsf{--Contra}}}}(\mathfrak{F},\mathfrak{P})=0$ for all flat quotseparated contramodule $R$-modules $\mathfrak{F}$;
  3. (3)
    one has $\operatorname{Ext}^{n}_{R{\operatorname{\mathsf{--Contra}}}}(\mathfrak{F},\mathfrak{P})=0$ for all flat quotseparated contramodule $R$-modules $\mathfrak{F}$ and all integers $n\geq 1$.
Proof.
All the equivalences follow from the facts that there are enough flat quotseparated contramodules in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ and the full subcategory of flat quotseparated contramodules is closed under kernels of epimorphisms (see Lemma 2.5.9(a)). ∎
Let $R$ be an adic topological ring. A quotseparated contramodule $R$-module $\mathfrak{P}$ is said to be cotorsion if it satisfies the equivalent conditions of Lemma 2.10.1. The full subcategory of cotorsion quotseparated contramodule $R$-modules is closed under extensions, cokernels of monomorphisms, and infinite products in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$.
Proposition 2.10.2.
Let $R$ be an adic topological ring and $\mathfrak{M}$ be a quotseparated contramodule $R$-module. Then
(a) there exists a short exact sequence of quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{F}\longrightarrow\mathfrak{M}\longrightarrow 0$ with a flat quotseparated contramodule $R$-module $\mathfrak{F}$ and a cotorsion quotseparated contramodule $R$-module $\mathfrak{P}$;
(b) there exists a short exact sequence of quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{F}\longrightarrow 0$ with a cotorsion quotseparated contramodule $R$-module $\mathfrak{P}$ and a flat quotseparated contramodule $R$-module $\mathfrak{F}$.
Proof.
This is a special case of [33, Corollary 7.8]. For a Noetherian ring $R$, one can also refer to [17, Corollary D.2.8(a–b)]. ∎
Corollary 2.10.3.
Let $R$ be an adic topological ring and $\mathfrak{F}$ be a quotseparated contramodule $R$-module. Then the following conditions are equivalent:
  1. (1)
    the functor $\operatorname{Hom}_{R}(\mathfrak{F},{-})$ takes short exact sequences of cotorsion quotseparated contramodule $R$-modules to short exact sequences of abelian groups/$R$-modules/quotseparated contramodule $R$-modules;
  2. (2)
    one has $\operatorname{Ext}^{1}_{R{\operatorname{\mathsf{--Contra}}}}(\mathfrak{F},\mathfrak{P})=0$ for all cotorsion quotseparated contramodule $R$-modules $\mathfrak{P}$;
  3. (3)
    one has $\operatorname{Ext}^{n}_{R{\operatorname{\mathsf{--Contra}}}}(\mathfrak{F},\mathfrak{P})=0$ for all cotorsion quotseparated contramodule $R$-modules $\mathfrak{F}$ and all integers $n\geq 1$;
  4. (4)
    $\mathfrak{F}$ is a flat quotseparated contramodule $R$-module.
Proof.
The argument is similar to the proof of Corollary 2.9.5 and based on Proposition 2.10.2 together with Lemma 2.10.1(2–3). For a Noetherian ring $R$, one can also refer to [17, Corollary D.2.8(d)]. ∎
Lemma 2.10.4.
Let $f\colon R\longrightarrow S$ be a continuous homomorphism of adic topological rings. Then, for any cotorsion quotseparated contramodule $S$-module $\mathfrak{P}$, the quotseparated contramodule $R$-module $f_{\sharp}(\mathfrak{P})$ is cotorsion.
Proof.
This is a special case of [32, Corollary 6.3]. The assertion is provable using the criterion of Lemma 2.10.1(2), by an argument based on the $\operatorname{Ext}^{1}$-adjunction lemma from [26, Lemma 1.7(e)] together with the results of Corollaries 2.6.4 and 2.6.5. ∎
The following lemma is a cotorsion version of Lemma 2.9.11.
Lemma 2.10.5.
Let $R$ be an adic topological ring with an ideal of definition $I\subset R$, and let $F$ be an $R$-module such that the $R/I^{n}$-module $F/I^{n}F$ is flat for all $n\geq 1$. Let $0\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}\longrightarrow 0$ be a short exact sequence of quotseparated contramodule $R$-modules with a cotorsion quotseparated contramodule $R$-module $\mathfrak{P}$. Then the short sequence of quotseparated contramodule $R$-modules $0\longrightarrow\operatorname{Hom}_{R}(F,\mathfrak{P})\longrightarrow\operatorname{Hom}_{R}(F,\mathfrak{M})\longrightarrow\operatorname{Hom}_{R}(F,\mathfrak{N})\longrightarrow 0$ is exact.
Proof.
The argument is similar to the proof of Lemma 2.9.11. One notices that $\mathbb{L}_{0}\Lambda(F)\simeq\Lambda(F)$ is a flat (quot)separated contramodule $R$-module and uses the fact that $\operatorname{Ext}^{1}_{R{\operatorname{\mathsf{--Contra}}}}(\Lambda(F),\mathfrak{P})=0$ by Lemma 2.10.1(2). ∎
Lemma 2.10.6.
Let $R$ be a Noetherian commutative ring endowed with an adic topology, $\mathfrak{R}$ be the adic completion of $R$, and $\mathfrak{P}$ be a contramodule $R$-module. Then the following conditions are equivalent:
  1. (1)
    $\mathfrak{P}$ is a cotorsion (quotseparated) contramodule $R$-module;
  2. (2)
    $\mathfrak{P}$ is a cotorsion $R$-module;
  3. (3)
    $\mathfrak{P}$ is a cotorsion $\mathfrak{R}$-module.
Proof.
Recall that a contramodule $R$-module $\mathfrak{F}$ is flat if and only if it is flat as an $R$-module (or as an $\mathfrak{R}$-module), by Lemma 2.5.10. Furthermore, the inclusion of abelian categories $R{\operatorname{\mathsf{--Contra}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ induces isomorphisms of all the Ext groups/modules, by Proposition 2.5.12. In view of these observations, the equivalence (1) $\Longleftrightarrow$ (2) is provided by [17, Corollary D.2.8(c)]. The equivalence (1) $\Longleftrightarrow$ (3) is a particular case of (1) $\Longleftrightarrow$ (2). ∎
2.11. The coextension of scalars is quotseparated
The aim of this section is to formulate and prove quotseparated versions of Lemmas 2.6.6(b) and 2.6.7(b). Our exposition is based on the ideas and techniques from the new preprint [32].
For any adic topological ring $R$ with an ideal of definition $I\subset R$ and any $R$-module $M$, we will denote the kernel of the $I$-adic completion map $\lambda_{I,M}\colon M\longrightarrow\Lambda_{I}(M)$ by $\Omega_{R}(M)=\bigcap_{n\geq 1}I^{n}M\subset M$.
Lemma 2.11.1.
Let $R$ and $S$ be adic topological rings, and let $f\colon R\longrightarrow S$ be a tight continuous ring map. Then an $S$-module $\mathfrak{Q}$ is separated and complete if and only if the $R$-module $f_{*}(\mathfrak{Q})$ is separated and complete. Moreover, for any $S$-module NN, one has $\Omega_{R}(N)=\Omega_{S}(N)\subset N$. Denoting by $I\subset R$ and $J\subset S$ any ideals of definition in $R$ and $S$, we also have a natural isomorphism of $R$-modules $\Lambda_{I}(N)\simeq\Lambda_{J}(N)$ forming a commutative triangular diagram with the completion maps $\lambda_{I,N}\colon N\longrightarrow\Lambda_{I}(N)$ and $\lambda_{J,N}\colon N\longrightarrow\Lambda_{J}(N)$.
Proof.
All the assertions follow immediately from Lemma 2.1.6(4). ∎
Let $f\colon R\longrightarrow S$ be a tight continuous map of adic topological rings. With Lemma 2.11.1 in mind, we will write simply $\Omega(N)$ instead of $\Omega_{R}(N)=\Omega_{S}(N)$, and $\Lambda(N)$ instead of $\Lambda_{I}(N)\simeq\Lambda_{J}(N)$, for any $S$-module NN. The completion map $N\longrightarrow\Lambda(N)$ will be denoted, as usual, by $\lambda_{N}$.
Theorem 2.11.2.
Let $R$ and $S$ be adic topological rings, and let $f\colon R\longrightarrow S$ be a flat tight continuous ring map. Then an $S$-module $\mathfrak{Q}$ is a quotseparated contramodule if and only if the $R$-module $f_{*}(\mathfrak{Q})$ is a quotseparated contramodule.
Proof.
This is our version of [32, Theorem 9.5]. The “only if” assertion holds by Lemma 2.6.1(d); the nontrivial implication is the “if”.
So let $\mathfrak{Q}$ be an $S$-module whose underlying $R$-module is a quotseparated contramodule. Then $\mathfrak{Q}$ is a contramodule $S$-module by Lemma 2.6.6(b). We need to pove that $\mathfrak{Q}$ is a quotseparated contramodule $S$-module.
Any contramodule $R$-module is complete, as mentioned in Section 2.4. Hence the completion map $\lambda_{\mathfrak{Q}}\colon\mathfrak{Q}\longrightarrow\Lambda(\mathfrak{Q})$ is surjective, and we have a short exact sequence of $S$-modules $0\longrightarrow\Omega(\mathfrak{Q})\longrightarrow\mathfrak{Q}\longrightarrow\Lambda(\mathfrak{Q})\longrightarrow 0$.
Pick a flat quotseparated contramodule $S$-module $\mathfrak{F}$ together with a surjective $S$-module map $\mathfrak{F}\longrightarrow\Lambda(\mathfrak{Q})$. For example, it suffices to take a free $S$-module $F$ together with a surjective $S$-module morphism $F\longrightarrow\Lambda(\mathfrak{Q})$, and put $\mathfrak{F}=\Lambda(F)$. Consider the pullback diagram in the category of $S$-modules
[…]
Here the $\mathfrak{Q}_{1}$ is the pullback of the pair of surjective maps $\lambda_{\mathfrak{Q}}\colon\mathfrak{Q}\longrightarrow\Lambda(\mathfrak{Q})$ adn $\mathfrak{F}\longrightarrow\Lambda(\mathfrak{Q})$. Both the top and bottom rows in (3) are short exact sequences. Since the underlying $R$-modules of $\mathfrak{Q}$,  $\Lambda(\mathfrak{Q})$, and $\mathfrak{F}$ are quotseparated contramodule $R$-modules, it follows that the underlying $R$-module of $\mathfrak{Q}_{1}$ is a quotseparated contramodule $R$-module, too.
By Lemma 2.6.9(d), the underlying $R$-module of $\mathfrak{F}$ is a flat quotseparated contramodule $R$-module. By Lemma 2.5.6,  $\mathfrak{F}$ is a separated contramodule $R$-module (as well as a separated contramodule $S$-module). Applying the functor $\Lambda$ and the natural transformation $\lambda$ to the surjective $S$-module map $\mathfrak{Q}_{1}\longrightarrow\mathfrak{F}$, we obtain a factorization of that map as the composition of surjective $S$-module maps
${\mathfrak{Q}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Lambda(\mathfrak{Q}_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathfrak{F}.}$
It follows, in view of (3), that the surjective $S$-module map $\mathfrak{Q}_{1}\longrightarrow\mathfrak{Q}$ induces an injective $S$-module map $\Omega(\mathfrak{Q}_{1})\longrightarrow\Omega(\mathfrak{Q})$.
Let us prove that the map $\Omega(\mathfrak{Q}_{1})\longrightarrow\Omega(\mathfrak{Q})$ cannot be an isomorphism if $\Omega(\mathfrak{Q})\neq 0$. Indeed, if the map $\Omega(\mathfrak{Q}_{1})\longrightarrow\Omega(\mathfrak{Q})$ is an isomorphism, then $\Lambda(\mathfrak{Q}_{1})\simeq\mathfrak{F}$. So $\Lambda(\mathfrak{Q}_{1})$ is a flat (quot)separated contramodule $R$-module. Since $\mathfrak{Q}_{1}$ is a quotseparated contramodule $R$-module, Corollary 2.5.8 is applicable, and it follows that $\Omega(\mathfrak{Q}_{1})=0$. Hence $\Omega(\mathfrak{Q})=0$.
Now we proceed by transfinite induction, iterating the construction above at the successor steps, and passing to the projective limit at the limit steps. Let $\aleph$ be a cardinal greater than the cardinality of $\mathfrak{Q}$. Viewing $\aleph$ as an ordinal, we will construct a projective system of $S$-modules $\mathfrak{Q}_{i}$,  $0\leq i<\aleph$, and surjective $S$-module morphisms $\mathfrak{Q}_{j}\longrightarrow\mathfrak{Q}_{i}$ for all ordinals $0\leq i\leq j<\aleph$. The underlying $R$-module of $\mathfrak{Q}_{i}$ will be a quotseparated contramodule $R$-module for every $0\leq i<\aleph$. The induced maps $\Omega(\mathfrak{Q}_{j})\longrightarrow\mathfrak{Q}_{i}$ will be injective for all $0\leq i\leq j<\aleph$, and the map $\Omega(\mathfrak{Q}_{i+1})\longrightarrow\Omega(\mathfrak{Q}_{i})$ will be a proper embedding (i. e., not an isomorphism) whenever $\Omega(\mathfrak{Q}_{i})\neq 0$.
[…]
After the construction is finished, we observe that a set $\mathfrak{Q}$ of the cardinality less than $\aleph$ cannot have a descending chain of subsets of length $\aleph$. Hence there exists an ordinal $k<\aleph$ for which the map $\Omega(\mathfrak{Q}_{k+1})\longrightarrow\Omega(\mathfrak{Q}_{k})$ is an isomorphism. As we have shown above, it follows that $\Omega(\mathfrak{Q}_{k})=0$.
Thus $\mathfrak{Q}_{k}$ is an $S$-module whose underlying $R$-module is separated and complete. By Lemma 2.11.1, it follows that $\mathfrak{Q}_{k}$ is separated and complete as an $S$-module. Thus the contramodule $S$-module $\mathfrak{Q}$ is a quotient $S$-module of a separated and complete $S$-module $\mathfrak{Q}_{k}$, so $\mathfrak{Q}$ is a quotseparated contramodule $S$-module. ∎
Remark 2.11.3.
Notice that the assertion of Theorem 2.11.2 with the flatness condition on $f$ replaced by the tightness condition (i. e., the obvious version of Lemma 2.6.6(b) for quotseparated contramodules) is not true. Indeed, let $S$ be an adic topological ring with an ideal of definition $J$ such that there exists a nonquotseparated contramodule $S$-module $\mathfrak{Q}$ (see Section 2.4). Let $s_{1}$, …, $s_{m}\in J$ be a finite set of generators of the ideal $J$. Consider the ring of polynomials $R=\mathbb{Z}[x_{1},\dotsc,x_{m}]$ in $m$ variables over the ring of integers $\mathbb{Z}$, and let $f\colon R\longrightarrow S$ be the ring homomorphism taking $x_{j}$ to $s_{j}$ for all $1\leq j\leq m$. Let $I=(x_{1},\dotsc,x_{m})\subset R$ be the ideal generated by the elements $x_{1}$, …, $x_{m}$. Endow $R$ with the $I$-adic topology. Then $f\colon R\longrightarrow S$ is a tight continuous ring map by Lemma 2.1.6(4), and $f_{*}(\mathfrak{Q})$ is a contramodule $R$-module by Lemma 2.6.1(b). Since the ring $R$ is Noetherian, all contramodule $R$-modules are quotseparated; so $f_{*}(\mathfrak{Q})$ is a quotseparated contramodule $R$-module. But $\mathfrak{Q}$ is not a quotseparated contramodule $S$-module.
Lemma 2.11.4.
Let $R$ and $S$ be adic topological rings, and let $f\colon R\longrightarrow S$ be a tight continuous ring map. Then, for any separated contramodule $R$-module $\mathfrak{P}$, the $S$-module $f^{!}(\mathfrak{P})=\operatorname{Hom}_{R}(S,\mathfrak{P})$ is also a separated contramodule.
First proof.
By Lemma 2.4.2(b), the underlying $R$-module of $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is separated and complete. By Lemma 2.11.1, it follows that the $S$-module $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is separated and complete as well. ∎
Second proof.
[…]
$\operatorname{Hom}_{R}(S,\mathfrak{P}/I^{n}\mathfrak{P})\simeq\operatorname{Hom}_{R/I^{n}}(S/Sf(I^{n}),\>\mathfrak{P}/I^{n}\mathfrak{P})=\operatorname{Hom}_{R/I^{n}}(S/J^{n},\mathfrak{P}/I^{n}\mathfrak{P}).$
Finally, $\operatorname{Hom}_{R/I^{n}}(S/J^{n},\mathfrak{P}/I^{n}\mathfrak{P})$ is an $S/J^{n}$-module. So the $S$-module $\operatorname{Hom}_{R}(S,{-})$ is a projective limit of $S/J^{n}$-modules. It remains to point out that all $S/J^{n}$-modules are separated contramodule $S$-modules, and the full subcategory of separated and the full subcategory of separated contramodule $S$-modules is closed under all limits in $S{\operatorname{\mathsf{--Mod}}}$. The latter assertion holds since the full subcategory of contramodule $S$-modules is closed under limits in $S{\operatorname{\mathsf{--Mod}}}$, while the full subcategory of separated $S$-modules is closed under submodules and infinite products. ∎
Proposition 2.11.5.
Let $R$ and $S$ be adic topological rings, and let $f\colon R\longrightarrow S$ be a flat tight continuous ring map. Then, for any quotseparated contramodule $R$-module $\mathfrak{P}$, the $S$-module $f^{!}(\mathfrak{P})=\operatorname{Hom}_{R}(S,\mathfrak{P})$ is also a quotseparated contramodule.
First proof.
By Lemma 2.4.2(c), the underlying $R$-module of $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is a quotseparated contramodule $R$-module. By Theorem 2.11.2, it follows that $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is a quotseparated contramodule $S$-module. ∎
Second proof.
The following argument is a special case of [32, proof of Theorem 10.4]. The $S$-module $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is a contramodule by Lemma 2.6.7(b); we only have to prove that it is a quotseparated contramodule. By Proposition 2.10.2, there exists a short exact sequence of quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{C}\longrightarrow\mathfrak{F}\longrightarrow\mathfrak{P}\longrightarrow 0$ with a flat quotseparated contramodule $R$-module $\mathfrak{F}$ and a cotorsion quotseparated contramodule $R$-module $\mathfrak{C}$. By Lemma 2.10.5 applied to the $R$-module $F=S$, the short sequence of $S$-modules $0\longrightarrow\operatorname{Hom}_{R}(S,\mathfrak{C})\longrightarrow\operatorname{Hom}_{R}(S,\mathfrak{F})\longrightarrow\operatorname{Hom}_{R}(S,\mathfrak{P})\longrightarrow 0$ is exact. Now the contramodule $R$-module $\mathfrak{F}$ is separated by Lemma 2.5.6, hence the contramodule $S$-module $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is separated by Lemma 2.11.4. Thus the contramodule $S$-module $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is a quotient $S$-module of a separated contramodule $S$-module $\operatorname{Hom}_{R}(S,\mathfrak{F})$, as desired. ∎
Now we can finish the discussion of right adjoint functors to the contramodule restriction of scalars that was started in Section 2.6. For any flat tight continuous map of adic topological rings $f\colon R\longrightarrow S$, Proposition 2.11.5 claims that the functor of coextension of scalars $f^{!}\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow S{\operatorname{\mathsf{--Mod}}}$ restricts to a functor
$f^{!}\colon R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muS{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}},$
which we will also call the coextension of scalars. The functor $f^{!}$ is right adjoint to the functor of restriction of scalars $f_{\sharp}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$.
2.12. Colocalization of contraadjustedness and colocality of cotorsion
We continue to use the notation $R{\operatorname{\mathsf{--Contra}}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$.
Corollary 2.12.1.
Let $f\colon R\longrightarrow S$ be a very flat tight continuous map of adic topological rings, and let $0\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}\longrightarrow 0$ be a short exact sequence of quotseparated contramodule $R$-modules with a contraadjusted quotseparated contramodule $R$-module $\mathfrak{P}$. Then the short sequence of quotseparated contramodule $S$-modules $0\longrightarrow f^{!}\mathfrak{P}\longrightarrow f^{!}\mathfrak{M}\longrightarrow f^{!}\mathfrak{N}\longrightarrow 0$ is exact.
Proof.
This is the particular case of Lemma 2.9.11 for the $R$-module $F=S$. ∎
Corollary 2.12.2.
Let $f\colon R\longrightarrow S$ be a flat tight continuous map of adic topological rings, and let $0\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}\longrightarrow 0$ be a short exact sequence of quotseparated contramodule $R$-modules with a cotorsion quotseparated contramodule $R$-module $\mathfrak{P}$. Then the short sequence of quotseparated contramodule $S$-modules $0\longrightarrow f^{!}\mathfrak{P}\longrightarrow f^{!}\mathfrak{M}\longrightarrow f^{!}\mathfrak{N}\longrightarrow 0$ is exact.
Proof.
This is the particular case of Lemma 2.10.5 for the $R$-module $F=S$. ∎
Proposition 2.12.3.
Let $f\colon R\longrightarrow S$ be a very flat tight continuous map of adic topological rings, and let $\mathfrak{P}$ be a contraadjusted quotseparated contramodule $R$-module. Then the quotseparated contramodule $S$-module $f^{!}(\mathfrak{P})=\operatorname{Hom}_{R}(S,\mathfrak{P})$ is contraadjusted.
Proof.
We have a pair of adjoint functors between abelian categories
${R{\operatorname{\mathsf{--Contra}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{!}}$${S{\operatorname{\mathsf{--Contra}}}.\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{\sharp}}$
The left adjoint functor $f_{\sharp}$ is exact. The object $\mathfrak{P}\in R{\operatorname{\mathsf{--Contra}}}$ is adjusted to the right adjoint functor $f^{!}$ in the following sense. By Lemma 2.9.11 applied to the $R$-module $F=S$, the short sequence of contramodule $S$-modules $0\longrightarrow f^{!}\mathfrak{P}\longrightarrow f^{!}\mathfrak{M}\longrightarrow f^{!}\mathfrak{N}\longrightarrow 0$ is exact for any short exact sequence of quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}\longrightarrow 0$. By [26, Lemma 1.7(b)], for any quotseparated contramodule $R$-module $\mathfrak{P}$ and any quotseparated contramodule $S$-module $\mathfrak{G}$, we have a natural injective map of abelian group
$\operatorname{Ext}^{1}_{S{\operatorname{\mathsf{--Contra}}}}(\mathfrak{G},f^{!}\mathfrak{P})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\operatorname{Ext}^{1}_{R{\operatorname{\mathsf{--Contra}}}}(f_{\sharp}\mathfrak{G},\mathfrak{P}),$
which is an isomorphism when $\mathfrak{P}$ is contraadjusted by [26, Lemma 1.7(e)]. Now if the quotseparated contramodule $S$-module $\mathfrak{G}$ is very flat, then so is the quotseparated contramodule $R$-module $f_{\sharp}\mathfrak{G}$, by Lemma 2.9.9(b). Using the criterion of Lemma 2.9.3(2), one concludes that the functor $f^{!}$ takes contraadjusted quotseparated contramodule $R$-modules to contraadjusted quotseparated contramodule $S$-modules. ∎
Proposition 2.12.4.
Let $f\colon R\longrightarrow S$ be a flat tight continuous map of adic topological rings, and let $\mathfrak{P}$ be a cotorsion quotseparated contramodule $R$-module. Then the quotseparated contramodule $S$-module $f^{!}(\mathfrak{P})=\operatorname{Hom}_{R}(S,\mathfrak{P})$ is cotorsion.
Proof.
This is the cotorsion version of Proposition 2.12.3, and the proof is very similar. One needs to use Lemma 2.10.5, Lemma 2.6.9(d), and the criterion of Lemma 2.10.1(2). ∎
We recall the notation $_{I}M=\operatorname{Hom}_{R}(R/I,M)$ for the submodule of all elements in an $R$-module $M$ annihilated by an ideal $I\subset R$.
Lemma 2.12.5.
Let $I$ be a finitely generated ideal in a commutative ring $R$. In this context:
(a) Let $\mathcal{M}$ be an $I$-torsion $R$-module. Assume that $_{I}\mathcal{M}=0$. Then $\mathcal{M}=0$.
(b) Let $\mathfrak{P}$ be an $I$-contramodule $R$-module. Assume that $\mathfrak{P}=I\mathfrak{P}$. Then $\mathfrak{P}=0$.
Proof.
Part (a) is obvious. Part (b) is a very basic example of an assertion from the class of results known as “contramodule Nakayama lemmas”. See [20, Lemma 4.2 and Remark 4.3], which are applicable by [20, Theorems 3.3 and 4.1]. A simple argument by induction on the number of generators of $I$, using the passage from $\mathfrak{P}$ to $\mathfrak{P}/s\mathfrak{P}$ (where $s\in I$ is one of the generators), is also possible. ∎
Lemma 2.12.6.
Let $R$ be an adic topological ring with an ideal of definition $I\subset R$. In this context:
(a) Let $0\longrightarrow\mathcal{K}^{0}\longrightarrow\mathcal{K}^{1}\longrightarrow\mathcal{K}^{2}\longrightarrow\dotsb$ be a bounded below complex of injective torsion $R$-modules. Then the complex $\mathcal{K}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is acyclic if and only if the complex of $R/I$-modules $\operatorname{Hom}_{R}(R/I,\mathcal{K}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic.
(b) Let $\dotsb\longrightarrow\mathfrak{F}_{2}\longrightarrow\mathfrak{F}_{1}\longrightarrow\mathfrak{F}_{0}\longrightarrow 0$ be a bounded above complex of flat quotseparated contramodule $R$-modules. Then the complex $\mathfrak{F}_{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is acyclic if and only if the complex of $R/I$-modules $R/I\otimes_{R}\mathfrak{F}_{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is acyclic.
Proof.
Part (a): the “only if” assertion holds since any acyclic bounded below complex of injective objects (in an abelian/exact category) is contractible. To prove the “if”, let $n\geq 0$ be the minimal integer for which the complex $\mathcal{K}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is not exact at the cohomological degree $n$. Then we have a finite contractible complex of injective torsion $R$-modules $0\longrightarrow\mathcal{K}^{0}\longrightarrow\mathcal{K}^{1}\longrightarrow\dotsb\longrightarrow\mathcal{K}^{n-1}\longrightarrow\mathcal{K}^{n}\longrightarrow\mathcal{J}^{n}\longrightarrow 0$, where $\mathcal{J}^{n}$ is the cokernel of the differential $\mathcal{K}^{n-1}\longrightarrow\mathcal{K}^{n}$. Furthermore, the natural map $\mathcal{J}^{n}\longrightarrow\mathcal{K}^{n+1}$ is not injective. Now $_{I}\mathcal{J}^{n}$ is the cokernel of the map $_{I}\mathcal{K}^{n-1}\longrightarrow{}_{I}\mathcal{K}^{n}$, while Lemma 2.12.5(a) implies that the map $_{I}\mathcal{J}^{n}\longrightarrow{}_{I}\mathcal{K}^{n+1}$ cannot be injective.
Part (b): the “only if” assertion holds by Lemma 2.5.9. To prove the “if”, let $n\geq 0$ be the minimal integer for which the complex $\mathfrak{F}_{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is not exact at the homological degree $n$. Then we have a finite exact complex of flat quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{G}_{n}\longrightarrow\mathfrak{F}_{n}\longrightarrow\mathfrak{F}_{n-1}\longrightarrow\dotsb\longrightarrow\mathfrak{F}_{1}\longrightarrow\mathfrak{F}_{0}\longrightarrow 0$, where $\mathfrak{G}_{n}$ is the kernel of the differential $\mathfrak{F}_{n}\longrightarrow\mathfrak{F}_{n-1}$. Furthermore, the natural map $\mathfrak{F}_{n+1}\longrightarrow\mathfrak{G}_{n}$ is not surjective. Now it follows from Lemma 2.5.9 that $R/I\otimes_{R}\mathfrak{G}_{n}$ is the kernel of the map $R/I\otimes_{R}\mathfrak{F}_{n}\longrightarrow R/I\otimes_{R}\mathfrak{F}_{n-1}$, while Lemma 2.12.5(b) implies that the map $R/I\otimes_{R}\mathfrak{F}_{n+1}\longrightarrow R/I\otimes_{R}\mathfrak{G}_{n}$ cannot be surjective. ∎
In the rest of this section, we work in the following setting. Let $f_{\alpha}\colon R\longrightarrow S$,  $1\leq\alpha\leq N$, be a finite formal open covering of an adic topological ring $R$. For every subsequence of indices $1\leq\alpha_{1}<\dotsb<\alpha_{k}\leq N$, we endow the commutative ring $S_{\alpha_{1},\dotsc,\alpha_{k}}=S_{\alpha_{1}}\otimes_{R}\dotsb\otimes_{R}S_{\alpha_{k}}$ with the tensor product topology, and denote by $f_{\alpha_{1},\dotsc,\alpha_{k}}$ the induced ring map $R\longrightarrow S_{\alpha_{1},\dotsc,\alpha_{k}}$.
Lemma 2.12.7.
Let $\mathfrak{F}$ be a flat quotseparated contramodule $R$-module. Then the cohomological Čech sequence of flat quotseparated contramodule $R$-modules
(4) $0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathfrak{F}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}\mathfrak{F}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}f_{\alpha,\beta}{}_{\sharp}f_{\alpha,\beta}^{\sharp}\mathfrak{F}\\ \mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\dotsb\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muf_{1,\dotsc,N}{}_{\sharp}f_{1,\dotsc,N}^{\sharp}\mathfrak{F}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0$
is exact.
Proof.
First of all, the differentials in the complex (4) are constructed in terms of the adjunction units for the adjoint functors $f^{\sharp}$ and $f_{\sharp}$ involved, as usual for the Čech complexes. Furthermore, since the compositions of formal open immersions are formal open immersions, it follows from Lemma 2.8.13(a) that the maps $f_{\alpha_{1},\dotsc,\alpha_{k}}$ are formal open immersions. In particular, $f_{\alpha_{1},\dotsc,\alpha_{k}}$ are flat tight continuous ring maps. Hence the functors $f_{\alpha_{1},\dotsc,\alpha_{k}}{}_{\sharp}$ take flat quotseparated contramodules to flat quotseparated contramodules by Lemma 2.6.9(d). The functors $f_{\alpha_{1},\dotsc,\alpha_{k}}^{\sharp}$ also take flat quotseparated contramodules to flat quotseparated contramodules by Corollary 2.6.4. Thus (4) is a finite complex of flat quotseparated contramodule $R$-modules.
Now let $I\subset R$ be an ideal of definition. Let $J_{\alpha_{1},\dotsc,\alpha_{k}}\subset S_{\alpha_{1},\dotsc,\alpha_{k}}$ be the ideals spanned by the $f_{\alpha_{1},\dotsc,\alpha_{k}}(I)\subset S_{\alpha_{1},\dotsc,\alpha_{k}}$. Put $\overline{R}=R/I$ and $\overline{S}_{\alpha_{1},\dotsc,\alpha_{k}}=S_{\alpha_{1},\dotsc,\alpha_{k}}/J_{\alpha_{1},\dotsc,\alpha_{k}}$, and let $\bar{f}_{\alpha_{1},\dotsc,\alpha_{k}}\colon\overline{R}\longrightarrow\overline{S}_{\alpha_{1},\dotsc,\alpha_{k}}$ be the induced maps of the quotient rings. Denote also by $\overline{F}$ the $\overline{R}$-module $R/I\otimes_{R}\mathfrak{F}$.
Applying the reduction functor $R/I\otimes_{R}{-}$ to the complex (4), we obtain the complex of flat $\overline{R}$-modules
(5) $0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\overline{F}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}\bar{f}_{\alpha}{}_{*}\bar{f}_{\alpha}^{*}\overline{F}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}\bar{f}_{\alpha,\beta}{}_{*}\bar{f}_{\alpha,\beta}^{*}\overline{F}\\ \mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\dotsb\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bar{f}_{1,\dotsc,N}{}_{*}\bar{f}_{1,\dotsc,N}^{*}\overline{F}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0,$
as one can see from Lemma 2.6.3(c). Now (5) is just the Čech complex for the $\overline{R}$-module $\overline{F}$ with respect to the open covering $\bar{f}_{\alpha}\colon\overline{R}\longrightarrow\overline{S}_{\alpha}$ of the (abstract, nontopological) commutative ring $\overline{R}$. So the complex (5) is exact, and by Lemma 2.12.6(b) it follows that the complex (4) is exact as well. ∎
Lemma 2.12.8.
Let $\mathcal{K}$ be an injective torsion $R$-module. Then the homological Čech complex of injective torsion $R$-modules
(6) $0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muf_{1,\dotsc,N}{}_{\diamond}f_{1,\dotsc,N}^{\diamond}\mathcal{K}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\dotsb\\ \mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}f_{\alpha,\beta}{}_{\diamond}f_{\alpha,\beta}^{\diamond}\mathcal{K}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}f_{\alpha}{}_{\diamond}f_{\alpha}^{\diamond}\mathcal{K}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathcal{K}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0$
is exact (or equivalently, contractible).
Proof.
This lemma is dual-analogous to Lemma 2.12.7. First of all, the differentials in the complex (6) are constructed in terms of the adjunction counits for the adjoint functors $f_{\diamond}$ and $f^{\diamond}$ involved. As mentioned in the proof of Lemma 2.12.7, the maps $f_{\alpha_{1},\dotsc,\alpha_{k}}$ are flat tight continuous ring maps. Hence the functors $f_{\alpha_{1},\dotsc,\alpha_{k}}{}_{\diamond}$ take injective torsion modules to injective torsion modules by Lemma 2.6.9(b). The functors $f_{\alpha_{1},\dotsc,\alpha_{k}}^{\diamond}$ also take injective torsion modules to injective torsion modules, as mentioned in Section 2.6. Thus (6) is a finite complex of injective torsion $R$-modules.
We use the notation from the proof of Lemma 2.12.7. Denote also by $\overline{K}$ the $\overline{R}$-module $\operatorname{Hom}_{R}(R/I,\mathcal{K})$. Following the discussion in Section 2.5, the $R/I$-module $\overline{K}$ is injective. In particular, it follows that $\overline{K}$ is a contraadjusted $\overline{R}$-module.
Applying the functor $\operatorname{Hom}_{R}(R/I,{-})$ to the complex (6), we obtain the complex of injective $\overline{R}$-modules
(7) $0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bar{f}_{1,\dotsc,N}{}_{*}\bar{f}_{1,\dotsc,N}^{!}\overline{K}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\dotsb\\ \mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}f_{\alpha,\beta}{}_{*}f_{\alpha,\beta}^{!}\overline{K}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}f_{\alpha}{}_{*}f_{\alpha}^{!}\overline{K}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\overline{K}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0,$
as one can see from Lemma 2.6.3(a). Now (7) is the homological Čech complex [17, formula (1.3) from Section 1.2] for the contraadjusted $\overline{R}$-module $\overline{K}$ with respect to the open covering $\bar{f}_{\alpha}\colon\overline{R}\longrightarrow\overline{S}_{\alpha}$ of the (abstract, nontopological) commutative ring $\overline{R}$. By [17, Lemma 1.2.6(b)], the complex (7) is exact, and by Lemma 2.12.6(a) above it follows that the complex (6) is exact as well. ∎
Corollary 2.12.9.
The cohomological Čech sequence of very flat quotseparated contramodule $R$-modules
(8) $0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\Lambda(R)\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}\Lambda(S_{\alpha})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}\Lambda(S_{\alpha,\beta})\\ \mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\dotsb\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\Lambda(S_{1,\dotsc,N})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0$
is exact.
Proof.
In the context of Lemma 2.12.7, put $\mathfrak{F}=\Lambda(R)$. Clearly, $\mathfrak{F}$ is a very flat (in fact, free) quotseparated contramodule $R$-module. As explained in the proof of Lemma 2.12.7, the maps $f_{\alpha_{1},\dotsc,\alpha_{k}}$ are formal open immersions; so they are very flat tight continuous ring maps. By Lemma 2.9.9(b), it follows that the complex (4) for $\mathfrak{F}=\Lambda(R)$ is a complex of very flat quotseparated contramodule $R$-modules. Furthermore, since $f_{\alpha_{1},\dotsc,\alpha_{k}}$ are tight continuous ring maps, the topology on $S_{\alpha_{1},\dotsc,\alpha_{k}}$ coincides with the $I$-adic topology (for any ideal of definition $I\subset R$, by Lemma 2.1.6(3)); so the notation $\Lambda(S_{\alpha_{1},\dotsc,\alpha_{k}})$ is unambiguous. Finally, following the discussion in Section 2.6, we have $f_{\alpha_{1},\dotsc,\alpha_{k}}^{\sharp}(\Lambda(R))\simeq\Lambda(S_{\alpha_{1},\dotsc,\alpha_{k}})$; so the complex (8) is the special case of (4) for $\mathfrak{F}=\Lambda(R)$. ∎
Lemma 2.12.10.
Let $\mathcal{M}$ be a torsion $R$-module. Then the cohomological Čech sequence of torsion $R$-modules
(9) $0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathcal{M}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}f_{\alpha}{}_{\diamond}f_{\alpha}^{*}\mathcal{M}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}f_{\alpha,\beta}{}_{\diamond}f_{\alpha,\beta}^{*}\mathcal{M}\\ \mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\dotsb\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muf_{1,\dotsc,N}{}_{\diamond}f_{1,\dotsc,N}^{*}\mathcal{M}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0$
is exact.
Proof.
The differentials in the complex (9) are constructed in terms of the adjunction units for the adjoint functors $f^{*}$ and $f_{\diamond}$ involved. We observe that the same complex (9) can be obtained by applying the functor $\mathcal{M}\otimes_{R}{-}$ to the complex (8). It follows from Lemma 2.5.9 that applying the functor $\mathcal{M}\otimes_{R}{-}$ takes the acyclic complex (8) to an acyclic complex. ∎
Lemma 2.12.11.
Let $\mathfrak{P}$ be a contraadjusted quotseparated contramodule $R$-module. Then the homological Čech sequence of contraadjusted quotseparated contramodule $R$-modules
(10) $0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muf_{1,\dotsc,N}{}_{\sharp}f_{1,\dotsc,N}^{!}\mathfrak{P}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\dotsb\\ \mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}f_{\alpha,\beta}{}_{\sharp}f_{\alpha,\beta}^{!}\mathfrak{P}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}f_{\alpha}{}_{\sharp}f_{\alpha}^{!}\mathfrak{P}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathfrak{P}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0$
is exact.
Proof.
This lemma is dual-analogous to Lemma 2.12.10. First of all, the differentials in the complex (10) are constructed in terms of the adjunction counits for the adjoint functors $f_{\sharp}$ and $f^{!}$ involved. Furthermore, the terms of the complex (10) are contraadjusted quotseparated contramodule $R$-modules by Lemma 2.9.7 and Proposition 2.12.3. Now we observe that the same complex (10) can be obtained by applying the functor $\operatorname{Hom}_{R}({-},\mathfrak{P})$ to the complex (8). The class of very flat quotseparated contramodule $R$-modules is closed under kernels of epimorphisms in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$, so the $R$-modules of cocycles in (8) are very flat quotseparated contramodule $R$-modules. Therefore, it follows from Lemma 2.9.3(1) that applying the functor $\operatorname{Hom}_{R}({-},\mathfrak{P})$ takes the acyclic complex (8) to an acyclic complex. ∎
Corollary 2.12.12.
Let $\mathfrak{P}$ be a contraadjusted quotseparated contramodule $R$-module. Assume that, for every index $\alpha$, the quotseparated contramodule $S_{\alpha}$-module $f_{\alpha}^{!}(\mathfrak{P})$ is cotorsion. Then the quotseparated contramodule $R$-module $\mathfrak{P}$ is cotorsion.
Proof.
By Lemma 2.10.4 and Proposition 2.12.4, if all the quotseparated contramodule $S_{\alpha}$-modules $f_{\alpha}^{!}(\mathfrak{P})$ are cotorsion, then all the quotseparated contramodule $R$-modules $f_{\alpha_{1},\dotsc,\alpha_{k}}{}_{\sharp}f_{\alpha_{1},\dotsc,\alpha_{k}}^{!}\mathfrak{P}$ for $k\geq 1$ are cotorsion, too. Recall that the class of cotorsion quotseparated contramodule $R$-modules is closed under cokernels of monomorphisms, as mentioned in Section 2.10. Going from the leftmost to the rightmost end of the complex (10) and arguing by induction, one proves that the quotseparated contramodule $R$-module $\mathfrak{P}$ is cotorsion as well. ∎
2.13. Colocality of exactness
Let $R$ be an adic topological ring. We will denote by $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}\subset R{\operatorname{\mathsf{--Contra}}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ the full subcategory of contraadjusted quotseparated contramodule $R$-modules. Similarly, the notation $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}\subset R{\operatorname{\mathsf{--Contra}}}$ stands for the full subcategory of cotorsion quotseparated contramodule $R$-modules.
As mentioned in Sections 2.9 and 2.10, the full subcategories $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ and $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ are closed under extensions, cokernels of monomorphisms, and infinite products in $R{\operatorname{\mathsf{--Contra}}}$. (The full subcategory $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ is even closed under quotients.) Thus these full subcategories inherit exact category structures from the abelian exact structure on $R{\operatorname{\mathsf{--Contra}}}$. We will always view $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ and $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ as exact categories with the exact structures inherited from (the abelian exact structure on) $R{\operatorname{\mathsf{--Contra}}}$.
We will also denote by $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}\subset R{\operatorname{\mathsf{--Contra}}}$ the full subcategory of flat quotseparated contramodule $R$-modules, and by $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{vfl}}\subset R{\operatorname{\mathsf{--Contra}}}$ the full subcategory of very flat quotseparated contramodule $R$-modules. According to Lemma 2.5.9(a) and the discussion in Section 2.9, the full subcategories $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ and $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{vfl}}$ are closed under extensions and kernels of epimorphisms in $R{\operatorname{\mathsf{--Contra}}}$. Thus these full subcategories also inherit exact category structures from the abelian exact structure on $R{\operatorname{\mathsf{--Contra}}}$. We will always view $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ and $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{vfl}}$ as exact categories with the exact structures inherited from $R{\operatorname{\mathsf{--Contra}}}$.
For any abstract (nontopological) commutative ring $R$, the similar notation $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cta}}\subset R{\operatorname{\mathsf{--Mod}}}$ and $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{vfl}}\subset R{\operatorname{\mathsf{--Mod}}}$ stands for the exact categories of contraadjusted and very flat $R$-modules. For an associative ring $R$, we denote by $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}\subset R{\operatorname{\mathsf{--Mod}}}$ and $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{fl}}\subset R{\operatorname{\mathsf{--Mod}}}$ the exact subcategories of cotorsion and flat left $R$-modules in the abelian category $R{\operatorname{\mathsf{--Mod}}}$.
Lemma 2.13.1.
Let $R$ be an adic topological ring and $\mathcal{K}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of injective torsion $R$-modules. Then the following three conditions are equivalent:
  1. (1)
    the complex $\mathcal{K}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contractible;
  2. (2)
    there exists an ideal of definition $I\subset R$ such that, for every $n\geq 1$, the complex of injective $R/I^{n}$-modules $_{I^{n}}\mathcal{K}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contractible;
  3. (3)
    for every open ideal $I\subset R$, the complex of injective $R/I$-modules $_{I}\mathcal{K}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contractible.
Proof.
(1) $\Longrightarrow$ (3) $\Longrightarrow$ (2) Obvious.
[…]
For a generalization, see [25, Lemma 4.21], which is applicable in view of [25, Section 2.4(6)]. ∎
Lemma 2.13.2.
Let $R$ be an adic topological ring and $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of flat quotseparated contramodule $R$-modules. Then the following three conditions are equivalent:
  1. (1)
    the complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in the exact category $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$;
  2. (2)
    there exists an ideal of definition $I\subset R$ such that, for every $n\geq 1$, the complex of flat $R/I^{n}$-modules $R/I^{n}\otimes_{R}\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in the exact category $R/I^{n}{\operatorname{\mathsf{--Mod}}}_{\mathsf{fl}}$;
  3. (3)
    for every open ideal $I\subset R$, the complex of flat $R/I$-modules $R/I\otimes_{R}\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in the exact category $R/I{\operatorname{\mathsf{--Mod}}}_{\mathsf{fl}}$.
Proof.
(1) $\Longrightarrow$ (3) Follows from Lemma 2.5.9(b).
(3) $\Longrightarrow$ (2) Obvious.
[…]
For a generalization, see [25, Lemma 4.13], which is applicable in view of [25, Example 3.8(2)]. ∎
Lemma 2.13.3.
Let $R$ be an adic topological ring and $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of very flat quotseparated contramodule $R$-modules. Then the following three conditions are equivalent:
  1. (1)
    the complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in the exact category $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{vfl}}$;
  2. (2)
    there exists an ideal of definition $I\subset R$ such that, for every $n\geq 1$, the complex of very flat $R/I^{n}$-modules $R/I^{n}\otimes_{R}\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in the exact category $R/I^{n}{\operatorname{\mathsf{--Mod}}}_{\mathsf{vfl}}$;
  3. (3)
    for every open ideal $I\subset R$, the complex of very flat $R/I$-modules $R/I\otimes_{R}\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in the exact category $R/I{\operatorname{\mathsf{--Mod}}}_{\mathsf{vfl}}$.
Proof.
Similar to the proof of Lemma 2.13.2. ∎
Lemma 2.13.4.
Let $R$ be an adic topological ring with an ideal of definition $I\subset R$. In this setting:
(a) Let $g\colon\mathcal{M}\longrightarrow\mathcal{N}$ be a morphism of torsion $R$-modules. Then the map $g$ is injective if and only if the induced map $\bar{g}\colon{}_{I}\mathcal{M}\longrightarrow{}_{I}\mathcal{N}$ is injective.
(b) Let $g\colon\mathfrak{P}\longrightarrow\mathfrak{Q}$ be a morphism of contramodule $R$-modules. Then the map $g$ is surjective if and only if the induced map $\bar{g}\colon\mathfrak{P}/I\mathfrak{P}\longrightarrow\mathfrak{Q}/I\mathfrak{Q}$ is surjective.
Proof.
Follows immediately from Lemma 2.12.5. ∎
Now let $f_{\alpha}\colon R\longrightarrow S$,  $1\leq\alpha\leq N$, be a finite formal open covering of an adic topological ring $R$. As in Section 2.12, we put $S_{\alpha_{1},\dotsc,\alpha_{k}}=S_{\alpha_{1}}\otimes_{R}\dotsb\otimes_{R}S_{\alpha_{k}}$ for all $1\leq\alpha_{1}<\dotsb<\alpha_{k}\leq N$, endow the rings $S_{\alpha_{1},\dotsc,\alpha_{k}}$ with the tensor product topologies, and denote the induced ring maps by $f_{\alpha_{1},\dotsc,\alpha_{k}}\colon R\longrightarrow S_{\alpha_{1},\dotsc,\alpha_{k}}$.
Corollary 2.13.5.
(a) A morphism of contramodule $R$-modules $\mathfrak{P}\longrightarrow\mathfrak{Q}$ is an epimorphism in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ if and only if the induced morphism of contramodule $S_{\alpha}$-modules $f_{\alpha}^{\#}(\mathfrak{P})\longrightarrow f_{\alpha}^{\#}(\mathfrak{Q})$ is an epimorphism in $S_{\alpha}{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ for every $\alpha$.
(b) A morphism of quotseparated contramodule $R$-modules $\mathfrak{P}\longrightarrow\mathfrak{Q}$ is an epimorphism in $R{\operatorname{\mathsf{--Contra}}}$ if and only if the induced morphism of quotseparated contramodule $S_{\alpha}$-modules $f_{\alpha}^{\sharp}(\mathfrak{P})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{Q})$ is an epimorphism in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}$ for every $\alpha$.
Proof.
Part (b): the “only if” assertion holds because the functor $f^{\sharp}$ is right exact (as a left adjoint functor) for any continuous homomorphism of adic topological rings $f\colon R\longrightarrow S$. To prove the “if”, let $I\subset R$ be an ideal of definition in $R$. By Lemma 2.13.4(b), it suffices to show that the map $\mathfrak{P}/I\mathfrak{P}\longrightarrow\mathfrak{Q}/I\mathfrak{Q}$ is surjective. Put $J_{\alpha}=S_{\alpha}f_{\alpha}(I)\subset S_{\alpha}$. Then the collection of ring homomorphisms $R/I\longrightarrow S_{\alpha}/J_{\alpha}$ is an open covering. So it is sufficient to check that the map $S_{\alpha}/J_{\alpha}\otimes_{R/I}\mathfrak{P}/I\mathfrak{P}\longrightarrow S_{\alpha}/J_{\alpha}\otimes_{R/I}\mathfrak{Q}/I\mathfrak{Q}$ is surjective for every $\alpha$. For this purpose, we notice that the map $f_{\alpha}^{\sharp}(\mathfrak{P})/J_{\alpha}f_{\alpha}^{\sharp}(\mathfrak{P})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{Q})/J_{\alpha}f_{\alpha}^{\sharp}(\mathfrak{Q})$ is surjective, and refer to Lemma 2.6.3(c). The proof of part (a) is similar and based on Lemma 2.6.3(b). ∎
Lemma 2.13.6.
(a) A short sequence of flat quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{F}\longrightarrow\mathfrak{G}\longrightarrow\mathfrak{H}\longrightarrow 0$ is exact if and only if the short sequence of flat quotseparated contramodule $S_{\alpha}$-modules $0\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{F})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{G})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{H})\longrightarrow 0$ is exact for every $\alpha$.
(b) A morphism of flat quotseparated contramodule $R$-modules $\mathfrak{F}\longrightarrow\mathfrak{G}$ is an admissible monomorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ if and only if the induced morphism of flat quotseparated contramodule $S_{\alpha}$-modules $f_{\alpha}^{\sharp}(\mathfrak{F})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{G})$ is an admissible monomorphism in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ for every $\alpha$.
(c) A morphism of flat quotseparated contramodule $R$-modules $\mathfrak{G}\longrightarrow\mathfrak{H}$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ if and only if the induced morphism of flat quotseparated contramodule $S_{\alpha}$-modules $f_{\alpha}^{\sharp}(\mathfrak{G})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{H})$ is an admissible epimorphism in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ for every $\alpha$.
(d) A complex of flat quotseparated contramodule $R$-modules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ if and only if the complex of flat quotseparated contramodule $S_{\alpha}$-modules $f_{\alpha}^{\sharp}(\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is exact in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ for every $\alpha$.
Proof.
Part (a): the “only if” assertion holds by Corollary 2.6.4. To prove the “if”, notice that, by the same corollary, exactness of the short sequences $0\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{F})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{G})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{H})\longrightarrow 0$ implies exactness of the short sequences $0\longrightarrow f_{\alpha_{1},\dotsc,\alpha_{k}}^{\sharp}(\mathfrak{F})\longrightarrow f_{\alpha_{1},\dotsc,\alpha_{k}}^{\sharp}(\mathfrak{G})\longrightarrow f_{\alpha_{1},\dotsc,\alpha_{k}}^{\sharp}(\mathfrak{H})\longrightarrow 0$ for all $k\geq 1$. Now the three-term sequence of flat quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{F}\longrightarrow\mathfrak{G}\longrightarrow\mathfrak{H}\longrightarrow 0$ induces a three-term sequence of finite exact sequences (4) from Lemma 2.12.7. Since the latter three-term sequence is exact on all the terms of the finite exact sequences (4) except perhaps the leftmost ones, it follows that the short sequence of the leftmost terms $0\longrightarrow\mathfrak{F}\longrightarrow\mathfrak{G}\longrightarrow\mathfrak{H}\longrightarrow 0$ is exact, too.
Part (b): the “only if” assertion follows from part (a). To prove the “if”, notice that the map $\mathfrak{F}\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{F})$ is an admissible monomorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ by Lemmas 2.12.7 and 2.5.9(a). If the map $f_{\alpha}^{\sharp}(\mathfrak{F})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{G})$ is an admissible monomorphism in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$, then the map $f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{F})\longrightarrow f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{G})$ is an admissible monomorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ by Lemma 2.6.9(d) for every $\alpha$. Hence the composition $\mathfrak{F}\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{F})\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{G})$ is an admissible monomorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$. The latter composition is equal to the composition $\mathfrak{F}\longrightarrow\mathfrak{G}\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{G})$, and it follows that $\mathfrak{F}\longrightarrow\mathfrak{G}$ is an admissible monomorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ by the dual version of [4, Proposition 7.6].
Part (c): a morphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ is an admissible epimorphism if and only if it is an epimorphism in $R{\operatorname{\mathsf{--Contra}}}$. So the assertion follows from Corollary 2.13.5(b).
Part (d): the “only if” assertion follows from part (a). To prove the “if”, let $I\subset R$ be an ideal of definition in $R$. By Lemma 2.13.2(2), it suffices to show that the complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}/I\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in $R/I{\operatorname{\mathsf{--Mod}}}_{\mathsf{fl}}$. Continuing in the notation from the proof of Corollary 2.13.5, it is sufficient to check that the complex $S_{\alpha}/J_{\alpha}\otimes_{R/I}\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}/I\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in $S_{\alpha}/J_{\alpha}{\operatorname{\mathsf{--Mod}}}_{\mathsf{fl}}$ for every $\alpha$. For this purpose, we notice that the complex $f_{\alpha}^{\sharp}(\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}})/J_{\alpha}f_{\alpha}^{\sharp}(\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic in $S_{\alpha}/J_{\alpha}{\operatorname{\mathsf{--Mod}}}_{\mathsf{fl}}$ and refer to Lemma 2.6.3(c). Alternatively, the argument similar to the proof of part (a) and dual-analogous to the proof of Lemma 2.13.9(c) below also works (see also the proof of Lemma 2.13.10(c)). ∎
Lemma 2.13.7.
(a) A short sequence of very flat quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{F}\longrightarrow\mathfrak{G}\longrightarrow\mathfrak{H}\longrightarrow 0$ is exact if and only if the short sequence of very flat quotseparated contramodule $S_{\alpha}$-modules $0\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{F})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{G})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{H})\longrightarrow 0$ is exact for every $\alpha$.
(b) A morphism of very flat quotseparated contramodule $R$-modules $\mathfrak{F}\longrightarrow\mathfrak{G}$ is an admissible monomorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{vfl}}$ if and only if the induced morphism of very flat quotseparated contramodule $S_{\alpha}$-modules $f_{\alpha}^{\sharp}(\mathfrak{F})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{G})$ is an admissible monomorphism in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}_{\mathsf{vfl}}$ for every $\alpha$.
(c) A morphism of very flat quotseparated contramodule $R$-modules $\mathfrak{G}\longrightarrow\mathfrak{H}$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{vfl}}$ if and only if the induced morphism of very flat quotseparated contramodule $S_{\alpha}$-modules $f_{\alpha}^{\sharp}(\mathfrak{G})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{H})$ is an admissible epimorphism in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}_{\mathsf{vfl}}$ for every $\alpha$.
(d) A complex of very flat quotseparated contramodule $R$-modules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{vfl}}$ if and only if the complex of flat quotseparated contramodule $S_{\alpha}$-modules $f_{\alpha}^{\sharp}(\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is exact in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}_{\mathsf{vfl}}$ for every $\alpha$.
Proof.
Part (a) is a particular case of Lemma 2.13.6(a) (the result of Lemma 2.9.6 is presumed). The proof of part (b) is similar to that of Lemma 2.13.6(b). One needs to use Lemmas 2.9.6 and 2.9.9(b). Part (c) is similar to Lemma 2.13.6(c). The proof of part (d) is similar to that of Lemma 2.13.6(d). One needs to use the criterion of Lemma 2.13.3(2). ∎
Lemma 2.13.8.
(a) A short sequence of torsion $R$-modules $0\longrightarrow\mathcal{L}\longrightarrow\mathcal{M}\longrightarrow\mathcal{N}\longrightarrow 0$ is exact if and only if the short sequence of torsion $S_{\alpha}$-modules $0\longrightarrow f_{\alpha}^{*}(\mathcal{L})\longrightarrow f_{\alpha}^{*}(\mathcal{M})\longrightarrow f_{\alpha}^{*}(\mathcal{N})\longrightarrow 0$ is exact for every $\alpha$.
(b) A morphism of torsion $R$-modules $\mathcal{L}\longrightarrow\mathcal{M}$ is injective if and only if the induced morphism of torsion $S_{\alpha}$-modules $f_{\alpha}^{*}(\mathcal{L})\longrightarrow f_{\alpha}^{*}(\mathcal{M})$ is injective for every $\alpha$.
(c) A morphism of torsion $R$-modules $\mathcal{M}\longrightarrow\mathcal{N}$ is surjective if and only if the induced morphism of torsion $S_{\alpha}$-modules $f_{\alpha}^{*}(\mathcal{M})\longrightarrow f_{\alpha}^{*}(\mathcal{N})$ is surjective for every $\alpha$.
(d) A complex of torsion $R$-modules $\mathcal{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact if and only if the complex of torsion $S_{\alpha}$-modules $f_{\alpha}^{*}(\mathcal{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is exact for every $\alpha$.
Proof.
The categories $R{\operatorname{\mathsf{--Tors}}}$ and $S_{\alpha}{\operatorname{\mathsf{--Tors}}}$ are abelian, and the functors $f_{\alpha}^{*}$ are exact by Lemma 2.6.9(a). Hence, in order to prove all the assertions (a–d), it suffices to check that one has $\mathcal{M}=0$ whenever $f_{\alpha}^{*}(\mathcal{M})=0$ for all $\alpha$ and some torsion $R$-module $\mathcal{M}$. The latter claim follows immediately from Lemma 2.12.10. ∎
Lemma 2.13.9.
(a) A short sequence of injective torsion $R$-modules $0\longrightarrow\mathcal{M}\longrightarrow\mathcal{L}\longrightarrow\mathcal{K}\longrightarrow 0$ is (split) exact if and only if the short sequence of injective torsion $S_{\alpha}$-modules $0\longrightarrow f_{\alpha}^{\diamond}(\mathcal{M})\longrightarrow f_{\alpha}^{\diamond}(\mathcal{L})\longrightarrow f_{\alpha}^{\diamond}(\mathcal{K})\longrightarrow 0$ is (split) exact for every $\alpha$.
(b) A morphism of injective torsion $R$-modules $\mathcal{L}\longrightarrow\mathcal{K}$ is a split epimorphism in $R{\operatorname{\mathsf{--Tors}}}^{\mathsf{inj}}$ if and only if the induced morphism of injective torsion $S_{\alpha}$-modules $f_{\alpha}^{\diamond}(\mathcal{L})\longrightarrow f_{\alpha}^{\diamond}(\mathcal{K})$ is a split epimorphism in $S_{\alpha}{\operatorname{\mathsf{--Tors}}}^{\mathsf{inj}}$ for every $\alpha$.
(c) A complex of injective torsion $R$-modules $\mathcal{K}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contractible if and only if the complex of injective torsion $S_{\alpha}$-modules $f_{\alpha}^{\diamond}(\mathcal{K}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is contractible for every $\alpha$.
Proof.
This lemma is dual-analogous to Lemmas 2.13.6(a,b,d) and 2.13.7(a,b,d). The “only if” assertions in all parts (a–c) are obvious. Let us prove the “if”.
Part (a): notice that (split) exactness of the short sequences $0\longrightarrow f_{\alpha}^{\diamond}(\mathcal{M})\longrightarrow f_{\alpha}^{\diamond}(\mathcal{L})\longrightarrow f_{\alpha}^{\diamond}(\mathcal{K})\longrightarrow 0$ implies (split) exactness of the short sequences $0\longrightarrow f_{\alpha_{1},\dotsc,\alpha_{k}}^{\diamond}(\mathcal{M})\longrightarrow f_{\alpha_{1},\dotsc,\alpha_{k}}^{\diamond}(\mathcal{L})\longrightarrow f_{\alpha_{1},\dotsc,\alpha_{k}}^{\diamond}(\mathcal{K})\longrightarrow 0$ for all $k\geq 1$. Now the three-term sequence of injective torsion $R$-modules $0\longrightarrow\mathcal{M}\longrightarrow\mathcal{L}\longrightarrow\mathcal{K}\longrightarrow 0$ induces a three-term sequence of finite split exact complexes (6) from Lemma 2.12.8. Since the latter three-term sequence is exact on all the terms of the finite exact sequences (6) except perhaps the rightmost ones, it follows that the short sequence of the rightmost terms $0\longrightarrow\mathcal{M}\longrightarrow\mathcal{L}\longrightarrow\mathcal{K}\longrightarrow 0$ is exact, too.
Part (b): notice that the map $\bigoplus_{\alpha}f_{\alpha}{}_{\diamond}f_{\alpha}^{\diamond}(\mathcal{K})\longrightarrow\mathcal{K}$ is a split epimorphism in $R{\operatorname{\mathsf{--Tors}}}^{\mathsf{inj}}$ by Lemma 2.12.8. If the map $f_{\alpha}^{\diamond}(\mathcal{L})\longrightarrow f_{\alpha}^{\diamond}(\mathcal{K})$ is a split epimorphism in $S_{\alpha}{\operatorname{\mathsf{--Tors}}}^{\mathsf{inj}}$, then the map $f_{\alpha}{}_{\diamond}f_{\alpha}^{\diamond}(\mathcal{L})\longrightarrow f_{\alpha}{}_{\diamond}f_{\alpha}^{\diamond}(\mathcal{K})$ is a split epimorphism in $R{\operatorname{\mathsf{--Tors}}}^{\mathsf{inj}}$ for every $\alpha$ (cf. Lemma 2.6.9(b)). Hence the composition $\bigoplus_{\alpha}f_{\alpha}{}_{\diamond}f_{\alpha}^{\diamond}(\mathcal{L})\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\diamond}f_{\alpha}^{\diamond}(\mathcal{K})\longrightarrow\mathcal{K}$ is a split epimorphism in $R{\operatorname{\mathsf{--Tors}}}^{\mathsf{inj}}$. The latter composition is equal to the composition $\bigoplus_{\alpha}f_{\alpha}{}_{\diamond}f_{\alpha}^{\diamond}(\mathcal{L})\longrightarrow\mathcal{L}\longrightarrow\mathcal{K}$, and it follows that $\mathcal{L}\longrightarrow\mathcal{K}$ is a split epimorphism in $R{\operatorname{\mathsf{--Tors}}}^{\mathsf{inj}}$, too.
Part (c): notice that if the complexes $f_{\alpha}^{\diamond}(\mathcal{K}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are contractible, then so are the complexes $f_{\alpha_{1},\dotsc,\alpha_{k}}^{\diamond}(\mathcal{K}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ for all $k\geq 1$. Applying the construction of the finite split exact sequence (6) from Lemma 2.12.8 to every term of the complex $\mathcal{K}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, we obtain a finite termwise split exact sequence of complexes in $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{tors}}$. Since all the complexes appearing in the latter finite termwise split exact sequence, except perhaps the rightmost one, are split exact, it follows that the rightmost complex $\mathcal{K}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is split exact, too. ∎
Lemma 2.13.10.
(a) A short sequence of contraadjusted quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{L}\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}\longrightarrow 0$ is exact if and only if the short sequence of contraadjusted quotseparated contramodule $S_{\alpha}$-modules $0\longrightarrow f_{\alpha}^{!}(\mathfrak{L})\longrightarrow f_{\alpha}^{!}(\mathfrak{M})\longrightarrow f_{\alpha}^{!}(\mathfrak{N})\longrightarrow 0$ is exact for every $\alpha$.
(b) A morphism of contraadjusted quotseparated contramodule $R$-modules $\mathfrak{M}\longrightarrow\mathfrak{N}$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ if and only if the induced morphism of contraadjusted quotseparated contramodule $S_{\alpha}$-modules $f_{\alpha}^{!}(\mathfrak{M})\longrightarrow f_{\alpha}^{!}(\mathfrak{N})$ is an admissible epimorphism in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ for every $\alpha$.
(c) A complex of contraadjusted quotseparated contramodule $R$-modules $\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ if and only if the complex of contraadjusted quotseparated contramodule $S_{\alpha}$-modules $f_{\alpha}^{!}(\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is exact in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ for every $\alpha$.
Proof.
Part (a): the “only if” assertion holds by Corollary 2.12.1 (the result of Proposition 2.12.3 is presumed). To prove the “if”, notice that, by the same corollary, exactness of the short sequences $0\longrightarrow f_{\alpha}^{!}(\mathfrak{L})\longrightarrow f_{\alpha}^{!}(\mathfrak{M})\longrightarrow f_{\alpha}^{!}(\mathfrak{N})\longrightarrow 0$ implies exactness of the short sequences $0\longrightarrow f_{\alpha_{1},\dotsc,\alpha_{k}}^{!}(\mathfrak{L})\longrightarrow f_{\alpha_{1},\dotsc,\alpha_{k}}^{!}(\mathfrak{M})\longrightarrow f_{\alpha_{1},\dotsc,\alpha_{k}}^{!}(\mathfrak{N})\longrightarrow 0$ for all $k\geq 1$. Now the three-term sequence of contraadjusted quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{L}\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}\longrightarrow 0$ induces a three-terms sequence of finite exact sequences (10) from Lemma 2.12.11. Since the latter three-term sequence is exact on all the terms of the finite exact sequences (10) except perhaps the rightmost ones, it follows that the short sequence of the rightmost terms $0\longrightarrow\mathfrak{L}\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}\longrightarrow 0$ is exact, too.
Part (b): the “only if” assertion follow from part (a). To prove the “if”, notice that the map $\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{N})\longrightarrow\mathfrak{N}$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ by Lemma 2.12.11. Here we are using the fact that the full subcategory $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ is closed under cokernels (in fact, even under quotients) in $R{\operatorname{\mathsf{--Contra}}}$. If the map $f_{\alpha}^{!}(\mathfrak{M})\longrightarrow f_{\alpha}^{!}(\mathfrak{N})$ is an admissible epimorphism in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$, then the map $f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{M})\longrightarrow f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{N})$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ by Lemma 2.9.7. Hence the composition $\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{M})\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{N})\longrightarrow\mathfrak{N}$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$. The latter composition is equal to the composition $\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{M})\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}$, and it follows that $\mathfrak{M}\longrightarrow\mathfrak{N}$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ by [4, Proposition 7.6].
Part (c): the “only if” assertion follow from part (a). To prove the “if”, notice that if the complexes $f_{\alpha}^{!}(\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are exact in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$, then the complexes $f_{\alpha_{1},\dotsc,\alpha_{k}}^{!}(\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are exact in $S_{\alpha_{1},\dotsc,\alpha_{k}}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ for all $k\geq 1$ (by the “only if” assertion). Hence the complexes $f_{\alpha_{1},\dotsc,\alpha_{k}}{}_{\sharp}f_{\alpha_{1},\dotsc,\alpha_{k}}^{!}(\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are exact in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ (by Lemma 2.9.7). Applying the construction of the finite exact sequence (10) from Lemma 2.12.11 to every term of the complex $\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, we obtain a finite termwise exact sequence of complexes in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$. Since all the complexes appearing in the latter finite termwise exact sequence, except perhaps the rightmost one, are exact in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$, it follows that the rightmost complex $\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$, too (e. g., by [4, Corollary 3.6(i)]). ∎
Lemma 2.13.11.
(a) A short sequence of cotorsion quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{L}\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}\longrightarrow 0$ is exact if and only if the short sequence of cotorsion quotseparated contramodule $S_{\alpha}$-modules $0\longrightarrow f_{\alpha}^{!}(\mathfrak{L})\longrightarrow f_{\alpha}^{!}(\mathfrak{M})\longrightarrow f_{\alpha}^{!}(\mathfrak{N})\longrightarrow 0$ is exact for every $\alpha$.
(b) A morphism of cotorsion quotseparated contramodule $R$-modules $\mathfrak{M}\longrightarrow\mathfrak{N}$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ if and only if the induced morphism of cotorsion quotseparated contramodule $S_{\alpha}$-modules $f_{\alpha}^{!}(\mathfrak{M})\longrightarrow f_{\alpha}^{!}(\mathfrak{N})$ is an admissible epimorphism in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ for every $\alpha$.
(c) A complex of cotorsion quotseparated contramodule $R$-modules $\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ if and only if the complex of cotorsion quotseparated contramodule $S_{\alpha}$-modules $f_{\alpha}^{!}(\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is exact in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ for every $\alpha$.
Proof.
This is similar to Lemma 2.13.10. The results of Proposition 2.12.4 and Lemma 2.10.4 are relevant. ∎
Notice that the obvious analogues of Lemmas 2.13.9(b), 2.13.10(b), and 2.13.11(b) for admissible (or split) monomorphisms are not true (not even for discrete regular Noetherian commutative rings $R$ and $S_{\alpha}$ of Krull dimension $1$). The counterexample can be discerned from [17, Example 3.2.1].
Proposition 2.13.12.
Let $R$ be an adic topological ring. In this context:
(a) a complex of contraadjusted quotseparated contramodule $R$-modules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ if and only if it is exact in $R{\operatorname{\mathsf{--Contra}}}$;
(b) a complex of cotorsion quotseparated contramodule $R$-modules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ if and only if it is exact in $R{\operatorname{\mathsf{--Contra}}}$.
Proof.
Part (a) is easy: the assertion follows from the fact that the full subcategory $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ is closed under quotients in $R{\operatorname{\mathsf{--Contra}}}$. Part (b) is the cotorsion periodicity theorem for contramodules; see [27, Corollary 12.4]. ∎
3. Contraherent Cosheaves of Contramodules
3.1. The basics
Let $\mathfrak{X}$ be a locally Noetherian formal scheme and $\mathfrak{O}=\mathfrak{O}_{\mathfrak{X}}$ be its structure sheaf. For every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$, the ring of functions $\mathfrak{R}=\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$ is a Noetherian commutative ring endowed with a complete, separated adic topology (in the sense of Section 2.1).
From this point on, by an adic Noetherian ring we will mean a Noetherian commutative ring endowed with adic topology. So the adic Noetherian rings are the adic topological rings $R$ whose underlying abstract ring $R$ is Noetherian.
An ideal $I\subset R$ is said to be an ideal of definition of a topological commutative ring $R$ if the topology of $R$ is the $I$-adic topology. This terminology was discussed already in Section 2.1. The ideals of definition $I\subset R$ are characterized by two properties: for every integer $n\geq 1$, the ideal $I^{n}$ is open in $R$; and the descending sequence $I^{n}$,  $n\geq 1$, of powers of the ideal $I$ is a base of open neighborhoods of zero in $R$.
Closed subschemes of definition $\overline{X}\subset\mathfrak{X}$ correspond bijectively to coherent sheaves of ideals of definition. The latter term means coherent sheaves of ideals $\mathfrak{I}$ on $\mathfrak{X}$ such that, for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$, the ideal $\mathfrak{I}(\mathfrak{U})\subset\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$ is an ideal of definition in the adic Noetherian ring $\mathfrak{R}=\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$. In particular, the minimal closed subscheme of definition $\overline{X}_{\mathrm{mn}}\subset\mathfrak{X}$ corresponds to the quasi-coherent sheaf formed by the maximal ideals of definition $\mathfrak{I}_{\mathrm{mx}}\subset\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$.
Let $\mathfrak{V}\subset\mathfrak{X}$ be another affine open formal subscheme in $\mathfrak{X}$ such that $\mathfrak{V}\subset\mathfrak{U}$. Put $\mathfrak{S}=\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})$. Then we have the restriction homomorphism of topological rings $f\colon\mathfrak{R}\longrightarrow\mathfrak{S}$. For every ideal of definition $\mathfrak{I}\subset\mathfrak{R}$, the extension $\mathfrak{J}=\mathfrak{S}f(\mathfrak{I})$ of the ideal $\mathfrak{I}$ to the ring $\mathfrak{S}$ is an ideal of definition $\mathfrak{J}\subset\mathfrak{S}$. So $f\colon\mathfrak{R}\longrightarrow\mathfrak{S}$ is a tight continuous ring map in the sense of Section 2.1. The discrete quotient ring $\overline{S}=\mathfrak{S}/\mathfrak{J}$ is the ring of regular functions on an affine open subscheme $\operatorname{Spec}\overline{S}$ in the affine scheme $\operatorname{Spec}\overline{R}$, where $\overline{R}=\mathfrak{R}/\mathfrak{I}$. In other words, $f\colon\mathfrak{R}\longrightarrow\mathfrak{S}$ is a formal open immersion of adic topological rings in the sense of Section 2.8.
If $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{V}_{\alpha}$ is an affine open covering of an affine formal scheme $\mathfrak{U}$ and $\mathfrak{I}\subset\mathfrak{R}$ is an ideal of definition in $\mathfrak{R}=\mathfrak{O}(\mathfrak{U})$, then the collection of affine open subschemes $\operatorname{Spec}\overline{S}_{\alpha}$ is an affine open covering of the affine scheme $\operatorname{Spec}\overline{R}$. Here $\overline{R}=\mathfrak{R}/\mathfrak{I}$ and $\overline{S}_{\alpha}=\mathfrak{S}_{\alpha}/\mathfrak{J}_{\alpha}$, where $\mathfrak{J}_{\alpha}=\mathfrak{S}_{\alpha}f_{\alpha}(\mathfrak{I})$ denotes the extension of the ideal $\mathfrak{I}$ to the ring $\mathfrak{S}_{\alpha}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha})$. The notation $f_{\alpha}$ stands for the restriction homomorphism of topological rings $f_{\alpha}\colon\mathfrak{R}\longrightarrow\mathfrak{S}_{\alpha}$. So the collection of ring maps $f_{\alpha}\colon\mathfrak{R}\longrightarrow\mathfrak{S}_{\alpha}$ is a formal open covering of the adic topological ring $\mathfrak{R}$ in the sense of Section 2.8.
Conversely, to any complete, separated adic Noetherian ring $\mathfrak{R}$, the affine Noetherian formal scheme $\operatorname{Spf}\mathfrak{R}$ (the formal spectrum of $\mathfrak{R}$) is naturally assigned. The contravariant functor $\operatorname{Spf}$ establishes an anti-equivalence between the category of complete, separated adic Notherian rings (and continuous ring homomorphisms) and the category of affine Noetherian formal schemes.
Under the anti-equivalence of categories from the previous paragraph, open immersions of affine Noetherian formal schemes $\operatorname{Spf}\mathfrak{S}\longrightarrow\operatorname{Spf}\mathfrak{R}$ correspond precisely to formal open immersions of complete, separated adic Noetherian rings $\mathfrak{R}\longrightarrow\mathfrak{S}$. Affine open coverings of affine Noetherian formal schemes $\operatorname{Spf}\mathfrak{S}_{\alpha}\longrightarrow\operatorname{Spf}\mathfrak{R}$ correspond precisely to formal open coverings of complete, separated adic Noetherian rings (by complete, separated adic Noetherian rings).
The abelian category of coherent sheaves on $\mathfrak{U}=\operatorname{Spf}\mathfrak{R}$ is naturally equivalent to the abelian category of finitely generated $\mathfrak{R}$-modules. The closed subschemes of definition $\overline{U}\subset\mathfrak{U}$ correspond bijectively to the ideals of definition $\mathfrak{I}\subset\mathfrak{R}$; the rule is $\overline{U}=\operatorname{Spec}\mathfrak{R}/\mathfrak{I}$. Arbitrary closed subschemes $\overline{U}^{\prime}\subset\mathfrak{U}$ correspond bijectively to the open ideals $\mathfrak{I}^{\prime}\subset\mathfrak{R}$; once again, the rule is $\overline{U}^{\prime}=\operatorname{Spec}\mathfrak{R}/\mathfrak{I}^{\prime}$.
Given an open covering $\mathbf{W}$ of a locally Noetherian formal scheme $\mathfrak{X}$, an open formal subscheme $\mathfrak{Y}\subset\mathfrak{X}$ is said to be subordinate to $\mathbf{W}$ if there exists an open subscheme $\mathfrak{W}\in\mathbf{W}$ for which $\mathfrak{Y}\subset\mathfrak{W}$. An open covering $\mathfrak{X}=\bigcup_{\alpha}\mathfrak{Y}_{\alpha}$ is said to be subordinate to $\mathbf{W}$ if $\mathfrak{Y}_{\alpha}$ is subordinate to $\mathbf{W}$ for every $\alpha$. More generally, the same terminology is applied to open coverings of arbitrary topological spaces.
3.2. Cosheaves of modules over a ringed space
Let $(\mathfrak{X},\mathfrak{O})$ be a ringed space and $\mathbf{B}$ be a base of open subsets in $\mathfrak{X}$. Recall the definitions of a copresheaf of $\mathfrak{O}$-modules on $\mathbf{B}$ and a cosheaf of $\mathfrak{O}$-modules on $\mathbf{B}$ from [17, Section 2.1] and [29, Section 2.1].
Let us start with (pre)sheaves. The set $\mathbf{B}$ is partially ordered by inclusion, and every poset can be viewed as a category. A presheaf of abelian groups $\mathcal{M}$ on $\mathbf{B}$ is a contravariant functor $\mathcal{M}\colon\mathbf{B}^{\mathsf{op}}\longrightarrow\mathsf{Ab}$ (where $\mathsf{Ab}=\mathbb{Z}{\operatorname{\mathsf{--Mod}}}$ is the category of abelian groups). A presheaf of $\mathfrak{O}$-modules $\mathcal{M}$ on $\mathbf{B}$ is a presheaf of abelian groups such that, for every open subset $\mathfrak{U}\subset\mathfrak{X}$,  $\mathfrak{U}\in\mathbf{B}$, an $\mathfrak{O}(\mathfrak{U})$-module structure is defined on the abelian group $\mathcal{M}(\mathfrak{U})$. It is required that the restriction map $\mathcal{M}(\mathfrak{U})\longrightarrow\mathcal{M}(\mathfrak{V})$ be an $\mathfrak{O}(\mathfrak{U})$-module morphism for every pair of open subsets $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$,  $\mathfrak{V}$, $\mathfrak{U}\in\mathbf{B}$.
A presheaf of abelian groups $\mathcal{M}$ on $\mathbf{B}$ is called a sheaf if the following sheaf axiom is satisfied. Let $\mathfrak{U}\subset\mathfrak{X}$,  $\mathfrak{U}\in\mathbf{B}$ be an open subset, $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{V}_{\alpha}$ be an open covering of $\mathfrak{U}$ by open subsets $\mathfrak{V}_{\alpha}\in\mathbf{B}$, and let $\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}=\bigcup_{\gamma}\mathfrak{W}_{\alpha\beta\gamma}$ be open coverings of the intersections $\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}$ by open subsets $\mathfrak{W}_{\alpha\beta\gamma}\in\mathbf{B}$. Then the short sequence
(11) $0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathcal{M}(\mathfrak{U})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\prod\nolimits_{\alpha}\mathcal{M}(\mathfrak{V}_{\alpha})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\prod\nolimits_{\alpha,\beta,\gamma}\mathcal{M}(\mathfrak{W}_{\alpha\beta\gamma})$
should be left exact. A presheaf of $\mathfrak{O}$-modules $\mathcal{M}$ on $\mathbf{B}$ is called a sheaf of $\mathfrak{O}$-modules if the underlying presheaf of abelian groups of $\mathcal{M}$ is a sheaf of abelian groups on $\mathbf{B}$.
We denote the additive category of sheaves of $\mathfrak{O}$-modules on $\mathbf{B}$ by $(\mathbf{B},\mathfrak{O}){\operatorname{\mathsf{--Sh}}}$. In the case of sheaves of $\mathfrak{O}$-modules defined an all the open subsets of $\mathfrak{X}$ (i. e., when the base of open subsets under consideration consists of all the open subsets in $\mathfrak{X}$), we denote the corresponding category of sheaves by $(\mathfrak{X},\mathfrak{O}){\operatorname{\mathsf{--Sh}}}$.
The main result in connection with the definition of a sheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$ is the following theorem from [11, Section 0.3.2], [13, Lemma Tag 009U], [17, Proposition 2.1.3], [29, Theorem 2.1(a)]: the natural restriction functor $(\mathfrak{X},\mathfrak{O}){\operatorname{\mathsf{--Sh}}}\longrightarrow(\mathbf{B},\mathfrak{O}){\operatorname{\mathsf{--Sh}}}$ is a category equivalence. So any sheaf of $\mathfrak{O}$-modules on $\mathbf{B}$ can be extended to a sheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$ in a unique way.
Now we turn to co(pre)sheaves. A copresheaf of abelian groups $\mathfrak{P}$ on $\mathbf{B}$ is a covariant functor $\mathfrak{P}\colon\mathbf{B}\longrightarrow\mathsf{Ab}$. We denote by $\mathfrak{P}[\mathfrak{U}]$ the abelian group that the copresheaf $\mathfrak{P}$ assigns to an open subset $\mathfrak{U}\subset\mathfrak{X}$,  $\mathfrak{U}\in\mathbf{B}$. The group $\mathfrak{P}[\mathfrak{U}]$ is called the group of cosections of the copresheaf $\mathfrak{P}$ over an open subset $\mathfrak{U}$. The abelian group map $\mathfrak{P}[\mathfrak{V}]\longrightarrow\mathfrak{P}[\mathfrak{U}]$ that the copresheaf $\mathfrak{P}$ assigns to an identity inclusion $\mathfrak{V}\subset\mathfrak{U}$ of open subsets $\mathfrak{V}$, $\mathfrak{U}\in\mathbf{B}$ is called the corestriction map.
A copresheaf of $\mathfrak{O}$-modules $\mathfrak{P}$ on $\mathbf{B}$ is a copresheaf of abelian groups such that, for every open subset $\mathfrak{U}\subset\mathfrak{X}$,  $\mathfrak{U}\in\mathbf{B}$, an $\mathfrak{O}(\mathfrak{U})$-module structure is defined on the abelian group $\mathfrak{P}[\mathfrak{U}]$. It is required that the corestriction map $\mathfrak{P}[\mathfrak{V}]\longrightarrow\mathfrak{P}[\mathfrak{U}]$ be an $\mathfrak{O}(\mathfrak{U})$-module morphism for every pair of open subsets $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$,  $\mathfrak{V}$, $\mathfrak{U}\in\mathbf{B}$.
A copresheaf of abelian groups $\mathfrak{P}$ on $\mathbf{B}$ is called a cosheaf if the following cosheaf axiom is satisfied. As above, let $\mathfrak{U}\subset\mathfrak{X}$,  $\mathfrak{U}\in\mathbf{B}$ be an open subset, $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{V}_{\alpha}$ be an open covering of $\mathfrak{U}$ by open subsets $\mathfrak{V}_{\alpha}\in\mathbf{B}$, and let $\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}=\bigcup_{\gamma}\mathfrak{W}_{\alpha\beta\gamma}$ be open coverings of the intersections $\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}$ by open subsets $\mathfrak{W}_{\alpha\beta\gamma}\in\mathbf{B}$. Then the short sequence
(12) $\bigoplus\nolimits_{\alpha,\beta,\gamma}\mathfrak{P}[\mathfrak{W}_{\alpha\beta\gamma}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha}\mathfrak{P}[\mathfrak{V}_{\alpha}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathfrak{P}[\mathfrak{U}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0$
should be right exact. A copresheaf of $\mathfrak{O}$-modules $\mathfrak{P}$ on $\mathbf{B}$ is called a cosheaf of $\mathfrak{O}$-modules if the underlying copresheaf of abelian groups of $\mathfrak{P}$ is a cosheaf of abelian groups on $\mathbf{B}$.
We denote the additive category of cosheaves of $\mathfrak{O}$-modules on $\mathbf{B}$ by $(\mathbf{B},\mathfrak{O}){\operatorname{\mathsf{--Cosh}}}$. In the case of cosheaves of $\mathfrak{O}$-modules defined an all the open subsets of $\mathfrak{X}$ (i. e., when the base of open subsets under consideration consists of all the open subsets in $\mathfrak{X}$), we denote the corresponding category of cosheaves by $(\mathfrak{X},\mathfrak{O}){\operatorname{\mathsf{--Cosh}}}$.
The main result in connection with the definition of a cosheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$ is the following theorem from [17, Theorem 2.1.2], [29, Theorem 2.1(b)]: the natural restriction functor $(\mathfrak{X},\mathfrak{O}){\operatorname{\mathsf{--Cosh}}}\longrightarrow(\mathbf{B},\mathfrak{O}){\operatorname{\mathsf{--Cosh}}}$ is a category equivalence. So any cosheaf of $\mathfrak{O}$-modules on $\mathbf{B}$ can be extended to a cosheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$ in a unique way.
Remarks 3.2.1.
(1) The sheaf axiom (11) implies, in particular, injectivity of the maps $\mathcal{M}(\mathfrak{U})\longrightarrow\prod_{\alpha}\mathcal{M}(\mathfrak{V}_{\alpha})$. It follows from the latter property that one need not check axiom (11) for all the choices of the open coverings $\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}=\bigcup_{\gamma}\mathfrak{W}_{\alpha\beta\gamma}$ of the intersections $\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}$. It suffices to check (11) for any one given choice of the open coverings of the pairwise intersections of the open subsets $\mathfrak{V}_{\alpha}$.
Dual-analogously, the cosheaf axiom (12) implies, in particular, surjectivity of the maps $\bigoplus_{\alpha}\mathfrak{P}[\mathfrak{V}_{\alpha}]\longrightarrow\mathfrak{P}[\mathfrak{U}]$. It follows from the latter property that it suffices to check axiom (12) for any one given choice of the open coverings $\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}=\bigcup_{\gamma}\mathfrak{W}_{\alpha\beta\gamma}$ of the intersections $\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}$.
(2) Assume that the base of open subsets $\mathbf{B}$ in $\mathfrak{X}$ consists of quasi-compact open subsets, and furthermore, the intersection of any two open subsets from $\mathbf{B}$ that are contained in a third open subset from $\mathbf{B}$ is also quasi-compact. Then, following [17, Remark 2.1.4], it suffices to check axioms (11) and (12) for finite open coverings $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{V}_{\alpha}$ and finite open coverings of the intersections $\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}=\bigcup_{\gamma}\mathfrak{W}_{\alpha\beta\gamma}$.
Let $(\mathfrak{X},\mathfrak{O})$ be a ringed space, $\mathfrak{W}\subset\mathfrak{X}$ be an open subset, $\mathcal{M}$ be a presheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$, and $\mathfrak{P}$ be a copresheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$. Then the notation $\mathcal{M}|_{\mathfrak{W}}$ and $\mathfrak{P}|_{\mathfrak{W}}$ stands for the restrictions of $\mathcal{M}$ and $\mathfrak{P}$ to open subsets $\mathfrak{U}\subset\mathfrak{X}$ contained in $\mathfrak{W}$, i. e., $\mathfrak{U}\subset\mathfrak{W}$. So $\mathcal{M}$ is a presheaf of $\mathfrak{O}|_{\mathfrak{W}}$-modules on $\mathfrak{W}$ and $\mathfrak{P}|_{\mathfrak{W}}$ is a copresheaf of $\mathfrak{O}|_{\mathfrak{W}}$-modules on $\mathfrak{W}$. The restriction of a (co)sheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$ to $\mathfrak{W}$ is a (co)sheaf of $\mathfrak{O}|_{\mathfrak{W}}$-modules on $\mathfrak{W}$.
Given a presheaf of $\mathfrak{O}$-modules $\mathcal{M}$ on a ringed space $(\mathfrak{U},\mathfrak{O})$ and a finite open covering $\mathfrak{U}=\bigcup_{\alpha=1}^{N}\mathfrak{V}_{\alpha}$ of $\mathfrak{U}$, we denote by $C^{\text{\smaller\smaller$\scriptstyle\bullet$}}(\{\mathfrak{V}_{\alpha}\},\mathcal{M})$ the finite cohomological Čech complex of abelian groups (in fact, $\mathfrak{O}(\mathfrak{U})$-modules)
(13) $0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}\mathcal{M}(\mathfrak{V}_{\alpha})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}\mathcal{M}(\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta})\\ \mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\dotsb\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathcal{M}(\mathfrak{V}_{1}\cap\dotsb\cap\mathfrak{V}_{N})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0.$
The sheaf axiom (11) (for finite open coverings $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{V}_{\alpha}$) essentially says that the natural map of $\mathfrak{O}(\mathfrak{U})$-modules $\mathcal{M}(U)\longrightarrow H^{0}C^{\text{\smaller\smaller$\scriptstyle\bullet$}}(\{\mathfrak{V}_{\alpha}\},\mathcal{M})$ is an isomorphism.
[…]
(14) $0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathfrak{P}[\mathfrak{V}_{1}\cap\dotsb\cap\mathfrak{V}_{N}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\dotsb\\ \mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}\mathfrak{P}[\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}\mathfrak{P}[\mathfrak{V}_{\alpha}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0.$
The cosheaf axiom (12) (for finite open coverings $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{V}_{\alpha}$) essentially says that the natural map of $\mathfrak{O}(\mathfrak{U})$-modules $H_{0}C^{\text{\smaller\smaller$\scriptstyle\bullet$}}(\{\mathfrak{V}_{\alpha}\},\mathfrak{P})\longrightarrow\mathfrak{P}[\mathfrak{U}]$ is an isomorphism.
3.3. Cosheaves of contramodules over a formal scheme
For any Noetherian commutative ring $R$ with an adic topology, we denote by $R{\operatorname{\mathsf{--Tors}}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\subset R{\operatorname{\mathsf{--Mod}}}$ the full subcategory of torsion $R$-modules and by $R{\operatorname{\mathsf{--Contra}}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\subset R{\operatorname{\mathsf{--Mod}}}$ the full subcategory of contramodule $R$-modules. All contramodule $R$-modules are quotseparated due to the assumption that the ring $R$ is Noetherian. See the notation and discussion in Sections 2.2, 2.4, and 2.9.
Let $\mathfrak{X}$ be a locally Noetherian formal scheme. A sheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules $\mathcal{M}$ on $\mathfrak{X}$ is said to be a sheaf of torsion $\mathfrak{O}_{\mathfrak{X}}$-modules if, for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$, the $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-module $\mathcal{M}(\mathfrak{U})$ is torsion/discrete as a module over a topological commutative ring with an adic topology. This means that the annihilator of any element of $\mathcal{M}(\mathfrak{U})$ is an open ideal in $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$, or equivalently, for every ideal of definition $\mathfrak{I}\subset\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$ and every element $x\in\mathcal{M}(\mathfrak{U})$ there exists an integer $n\geq 1$ such that $\mathfrak{I}^{n}x=0$ in $\mathcal{M}(\mathfrak{U})$.
The sheaf axiom for $\mathcal{M}$ implies that it suffices to check the torsion condition for affine open subschemes $\mathfrak{U}\subset\mathfrak{X}$ subordinate to a given open covering of $\mathfrak{X}$. Indeed, if $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{V}_{\alpha}$ is a finite affine open covering of an affine formal scheme $\mathfrak{U}$ and $\mathcal{N}$ is a sheaf of $\mathfrak{O}_{\mathfrak{U}}$-modules on $\mathfrak{U}$, then $\mathcal{N}(\mathfrak{U})$ is a torsion module over $\mathfrak{R}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$ whenever $\mathcal{N}(\mathfrak{V}_{\alpha})$ is a torsion module over $\mathfrak{S}_{\alpha}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha})$ for all indices $\alpha$. The latter assertion follows from the facts that the map $\mathcal{N}(\mathfrak{U})\longrightarrow\bigoplus_{\alpha}\mathcal{N}(\mathfrak{V}_{\alpha})$ is injective, the functors of restrictions of scalars $f_{\alpha}{}_{*}\colon\mathfrak{S}_{\alpha}{\operatorname{\mathsf{--Mod}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}$ take $\mathfrak{S}_{\alpha}{\operatorname{\mathsf{--Tors}}}$ into $\mathfrak{R}{\operatorname{\mathsf{--Tors}}}$ (see Section 2.6), and any submodule of a torsion $\mathfrak{R}$-module is torsion.
A cosheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules $\mathfrak{P}$ on $\mathfrak{X}$ is said to be a cosheaf of contramodule $\mathfrak{O}_{\mathfrak{X}}$-modules if, for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$, the $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-module $\mathfrak{P}[\mathfrak{U}]$ is a contramodule over the adic topological ring $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$. We refer to Section 2.4 for the definition of contramodule $R$-modules for an adic topological ring $R$. Let us only point out that, since the commutative ring $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$ is Noetherian by assumption, all the contramodule $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules are quotseparated.
The cosheaf axiom for $\mathfrak{P}$ implies that is suffices to check the contramoduleness condition for affine open subschemes $\mathfrak{U}\subset\mathfrak{X}$ subordinate to a given open covering of $\mathfrak{X}$. Indeed, let $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{V}_{\alpha}$ be a finite affine open covering of an affine formal scheme $\mathfrak{U}$. Let $\mathfrak{Q}$ be a cosheaf of $\mathfrak{O}_{\mathfrak{U}}$-modules on $\mathfrak{U}$. Then the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$-module $\mathfrak{Q}[\mathfrak{U}]$ is a contramodule over the adic topological ring $\mathfrak{R}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$ whenever the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha})$-module $\mathfrak{Q}[\mathfrak{V}_{\alpha}]$ is a contramodule over the adic topological ring $\mathfrak{S}_{\alpha}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha})$ and the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha,\beta})$-module $\mathfrak{Q}[\mathfrak{V}_{\alpha,\beta}]$ is a contramodule over the adic topological ring $\mathfrak{S}_{\alpha,\beta}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha,\beta})$ for all indices $\alpha$ and $\beta$. Here we put $\mathfrak{V}_{\alpha,\beta}=\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}\subset\mathfrak{U}$.
The latter assertion follows from right exactness of the short sequence (12),
$\bigoplus\nolimits_{\alpha,\beta}\mathfrak{Q}[\mathfrak{V}_{\alpha,\beta}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha}\mathfrak{Q}[\mathfrak{V}_{\alpha}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathfrak{Q}[\mathfrak{U}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0,$
together with the facts that the functors of restriction of scalars $\mathfrak{S}_{\alpha}{\operatorname{\mathsf{--Mod}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}$ and $\mathfrak{S}_{\alpha,\beta}{\operatorname{\mathsf{--Mod}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}$ take $\mathfrak{S}_{\alpha}{\operatorname{\mathsf{--Contra}}}$ and $\mathfrak{S}_{\alpha,\beta}{\operatorname{\mathsf{--Contra}}}$ into $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ (see Section 2.6), and that the full subcategory of contramodule $\mathfrak{R}$-modules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ is closed under cokernels in $\mathfrak{R}{\operatorname{\mathsf{--Mod}}}$.
3.4. Quasi-coherent torsion sheaves
Let $\mathfrak{X}$ be a locally Noetherian formal scheme with an open covering $\mathbf{W}$. Denote by $\mathbf{B}$ the base of open subsets in $\mathfrak{X}$ consisting of all the affine open formal subschemes $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$.
Let $\mathcal{M}$ be a presheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules on $\mathbf{B}$. We will say that $\mathcal{M}$ is a quasi-coherent torsion presheaf if the following conditions hold:
  1. (i)
    For every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$, the $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-module $\mathcal{M}(\mathfrak{U})$ is torsion (as a module over the adic topological ring $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$).
  2. (ii)
    For every pair of affine open formal subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$, the $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})$-module map
    $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})}\mathcal{M}(\mathfrak{U})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathcal{M}(\mathfrak{V})$
    corresponding by adjunction to the restriction map $\mathcal{M}(\mathfrak{U})\longrightarrow\mathcal{M}(\mathfrak{V})$ is an isomorphism.
We will call condition (i) is the torsion axiom, while condition (ii) will be called the quasi-coherence axiom. This definition is to be compared with the Enochs–Estrada description of quasi-coherent sheaves on schemes [6, Section 2], [29, Section 2.2].
Lemma 3.4.1.
The torsion and quasi-coherence axioms imply the sheaf axiom: Any quasi-coherent torsion presheaf on $\mathbf{B}$ is a sheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules on $\mathbf{B}$.
Proof.
The assertion follows from exactness of the leftmost fragment of sequence (9) in Lemma 2.12.10 together with Remarks 3.2.1.
In fact, under our assumption of local Noetherianity of $\mathfrak{X}$, one need not use the torsion axiom for this argument, and the quasi-coherence axiom alone is sufficient. Inceed, the finite exact sequence (8) from Corollary 2.12.9 is an exact sequence of flat $R$-modules by Lemma 2.5.10(6) or (7), so it remains exact after taking the tensor product with an arbitrary $R$-module $\mathcal{M}(\mathfrak{U})$ for $R=\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$. ∎
So any quasi-coherent torsion presheaf $\mathcal{M}$ on $\mathbf{B}$ is, in fact, a sheaf on $\mathbf{B}$, and we will speak of $\mathcal{M}$ as a quasi-coherent torsion sheaf on $\mathbf{B}$. By Lemma 3.4.1 and in view of the discussion in Section 3.2, any quasi-coherent torsion sheaf on $\mathbf{B}$ can be extended to a sheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules on the whole formal scheme $\mathfrak{X}$ in a unique way. Moreover, according to the discussion in Section 3.3, the resulting sheaf is a sheaf of torsion $\mathfrak{O}_{\mathfrak{X}}$-modules on $\mathfrak{X}$.
The following proposition establishes the fact that the notion of a quasi-coherent torsion sheaf is local (i. e., does not depend on the open covering $\mathbf{W}$).
Proposition 3.4.2.
Denote by $\mathbf{W}_{\mathfrak{X}}$ the trivial open covering $\{\mathfrak{X}\}$ of the formal scheme $\mathfrak{X}$, and let $\mathbf{B}_{\mathfrak{X}}$ be the base of open subsets in $\mathfrak{X}$ consisting of all affine open formal subschemes of $\mathfrak{X}$. Let $\mathcal{M}$ be a sheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules on $\mathbf{B}_{\mathfrak{X}}$. Assume that the restriction of $\mathcal{M}$ to $\mathbf{B}$ is a quasi-coherent torsion sheaf on $\mathbf{B}$. Then $\mathcal{M}$ is a quasi-coherent torsion sheaf on $\mathbf{B}_{\mathfrak{X}}$.
Proof.
Let $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ be two affine open formal subschemes, and let $\mathfrak{U}=\bigcup_{\alpha=1}^{N}\mathfrak{W}_{\alpha}$ be a finite open covering of $\mathfrak{U}$ by affine open formal subschemes $\mathfrak{W}_{\alpha}$ subordinate to $\mathbf{W}$. The argument from Section 3.3 explains that $\mathcal{M}(\mathfrak{U})$ is a torsion module over $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$ whenever $\mathcal{M}(\mathfrak{W}_{\alpha})$ are torsion modules over $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{W}_{\alpha})$ for all $\alpha$. So the torsion axiom (i) is satisfied for the sheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules $\mathcal{M}$ on $\mathbf{B}_{\mathfrak{X}}$. We need to check the quasi-coherence axiom (ii) for $\mathcal{M}$ with respect to $\mathfrak{V}\subset\mathfrak{U}$.
Notice that $\mathfrak{V}=\bigcup_{\alpha=1}^{N}(\mathfrak{V}\cap\mathfrak{W}_{\alpha})$ is an finite open covering of $\mathfrak{V}$ by affine open formal subschemes subordinate to $\mathbf{W}$. By the sheaf axiom (11), we have a left exact sequence of torsion $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules
(15) $0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathcal{M}(\mathfrak{U})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}\mathcal{M}(\mathfrak{W}_{\alpha})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}\mathcal{M}(\mathfrak{W}_{\alpha}\cap\mathfrak{W}_{\beta})$
and a left exact sequence of torsion $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})$-modules
(16) $0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathcal{M}(\mathfrak{V})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}\mathcal{M}(\mathfrak{V}\cap\mathfrak{W}_{\alpha})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}\mathcal{M}(\mathfrak{V}\cap\mathfrak{W}_{\alpha}\cap\mathfrak{W}_{\beta}).$
Recall that $\mathfrak{O}(\mathfrak{V})$ is a flat contramodule $\mathfrak{O}(\mathfrak{U})$-module by Lemma 2.8.6(3) (and a flat $\mathfrak{O}(\mathfrak{U})$-module by Lemma 2.5.10(6) or (7)). Hence the tensor product functor $\mathfrak{O}(\mathfrak{V})\otimes_{\mathfrak{O}(\mathfrak{U})}{-}$ preserves exactness of the left exact sequence (15), and we obtain a left exact sequence of $\mathfrak{O}(\mathfrak{V})$-modules
(17) $0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})}\mathcal{M}(\mathfrak{U})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})}\mathcal{M}(\mathfrak{W}_{\alpha})\\ \mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})}\mathcal{M}(\mathfrak{W}_{\alpha}\cap\mathfrak{W}_{\beta}).$
The restriction maps in the (pre)sheaf $\mathcal{M}$ induce a natural morphism from the left exact sequence (17) to the left exact sequence (16). It remains to refer to the next Lemma 3.4.3 to the effect that the map from (17) to (16) is an isomorphism on the middle and rightmost terms. Hence the natural map $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})}\mathcal{M}(\mathfrak{U})\longrightarrow\mathcal{M}(\mathfrak{V})$ of the leftmost terms in an isomorphism, too. ∎
Lemma 3.4.3.
Let $\mathcal{M}$ be a quasi-coherent torsion sheaf on $\mathbf{B}$, let $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ be affine open subschemes, and let $\mathfrak{W}\subset\mathfrak{U}$ be an affine open subscheme subordinate to $\mathbf{W}$. Then the restriction map of $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{W})$-modules $\mathcal{M}(\mathfrak{W})\longrightarrow\mathcal{M}(\mathfrak{V}\cap\mathfrak{W})$ induces an isomorphism of $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})$-modules $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})}\mathcal{M}(\mathfrak{W})\longrightarrow\mathcal{M}(\mathfrak{V}\cap\mathfrak{W})$.
Proof.
Both $\mathfrak{W}$ and $\mathfrak{V}\cap\mathfrak{W}$ are affine open formal subschemes in $\mathfrak{X}$ subordinate to $\mathbf{W}$, so we have $\mathfrak{W}\in\mathbf{B}$ and $\mathfrak{V}\cap\mathfrak{W}\in\mathbf{B}$. By the quasi-coherence axiom (ii), the restriction map $\mathcal{M}(\mathfrak{W})\longrightarrow\mathcal{M}(\mathfrak{V}\cap\mathfrak{W})$ induces an isomorphism of $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V}\cap\mathfrak{W})$-modules
$\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V}\cap\mathfrak{W})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{W})}\mathcal{M}(\mathfrak{W})\simeq\mathcal{M}(\mathfrak{V}\cap\mathfrak{W}).$
There is also an obvious isomorphism
$\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})}\mathcal{M}(\mathfrak{W})\simeq\bigl(\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})}\mathfrak{O}_{\mathfrak{X}}(\mathfrak{W})\bigr)\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{W})}\mathcal{M}(\mathfrak{W}).$
It remains to show that the map
$\bigl(\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})}\mathfrak{O}_{\mathfrak{X}}(\mathfrak{W})\bigr)\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{W})}\mathcal{M}(\mathfrak{W})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V}\cap\mathfrak{W})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{W})}\mathcal{M}(\mathfrak{W})$
induced by the multiplication map of rings
(18) $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})}\mathfrak{O}_{\mathfrak{X}}(\mathfrak{W})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V}\cap\mathfrak{W})$
is an isomorphism. Now the map (18) becomes an isomorphism after the adic completion functor $\Lambda$ is applied, while the $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{W})$-module $\mathcal{M}(\mathfrak{W})$ is torsion, and it remains to refer to Corollary 2.5.4. ∎
Finally, let $\mathcal{M}$ be a sheaf of torsion $\mathfrak{O}_{\mathfrak{X}}$-modules on $\mathfrak{X}$. We will say that $\mathcal{M}$ is a quasi-coherent torsion sheaf on $\mathfrak{X}$ if the quasi-coherence axiom (ii) holds for all affine open subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ (or equivalenlty, for all affine open subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$). According to the preceding discussion in this section, the category of quasi-coherent torsion sheaves on $\mathfrak{X}$ is equivalent to the category of quasi-coherent torsion (pre)sheaves on the topology base $\mathbf{B}$ consisting of all affine open formal subschemes of $\mathfrak{X}$ subordinate to $\mathbf{W}$ (for any given open covering $\mathbf{W}$ of the formal scheme $\mathfrak{X}$).
We will denote the category of quasi-coherent torsion sheaves on $\mathfrak{X}$ by $\mathfrak{X}{\operatorname{\mathsf{--Tors}}}$. One can see that $\mathfrak{X}{\operatorname{\mathsf{--Tors}}}$ is a Grothendieck abelian category and, for every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$, the functor of sections $(\mathfrak{U})\colon\mathfrak{X}{\operatorname{\mathsf{--Tors}}}\longrightarrow\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U}){\operatorname{\mathsf{--Tors}}}\subset\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U}){\operatorname{\mathsf{--Mod}}}$ is an exact functor preserving infinite direct sums. Lemma 2.6.9(a) is relevant here.
A short sequence $0\longrightarrow\mathcal{L}\longrightarrow\mathcal{M}\longrightarrow\mathcal{N}\longrightarrow 0$ of quasi-coherent torsion sheaves on $\mathfrak{X}$ is exact in $\mathfrak{X}{\operatorname{\mathsf{--Tors}}}$ if and only if, for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$, the short sequence of torsion $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules $0\longrightarrow\mathcal{L}(\mathfrak{U})\longrightarrow\mathcal{M}(\mathfrak{U})\longrightarrow\mathcal{N}(\mathfrak{U})\longrightarrow 0$ is exact. By Lemma 2.13.8(a), it suffices to check this condition for affine open subschemes $\mathfrak{U}$ belonging to any chosen affine open covering of $\mathfrak{X}$.
Moreover, a complex $\mathcal{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathfrak{X}{\operatorname{\mathsf{--Tors}}}$ is exact if and only if, for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$, the complex of torsion $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules is exact in $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U}){\operatorname{\mathsf{--Tors}}}$. By Lemma 2.13.8(d), it suffices to check this condition for affine open subschemes $\mathfrak{U}$ belonging to any chosen affine open covering of $\mathfrak{X}$.
For any affine Noetherian formal scheme $\mathfrak{U}$, the functor of global sections $\mathcal{M}\longmapsto\mathcal{M}(\mathfrak{U})$ establishes an equivalence between the abelian category of quasi-coherent torsion sheaves $\mathfrak{U}{\operatorname{\mathsf{--Tors}}}$ on $\mathfrak{U}$ and the abelian category $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U}){\operatorname{\mathsf{--Tors}}}$ of torsion modules over the adic Noetherian ring $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$.
3.5. Contraherent cosheaves of contramodules
We keep the notation of Section 3.4. So $\mathfrak{X}$ is a locally Noetherian formal scheme, $\mathbf{W}$ is an open covering of $\mathfrak{X}$, and $\mathbf{B}$ is the base of open subsets in $\mathfrak{X}$ consisting of all the affine open subschemes $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$.
Let $\mathfrak{P}$ be a copresheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules on $\mathbf{B}$. We will say that $\mathfrak{P}$ is a contraherent copresheaf of contramodules (or a contraherent contramodule copresheaf) if the following conditions hold:
  1. (iii)
    For every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$, the $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-module $\mathfrak{P}[\mathfrak{U}]$ is a contramodule (as a module over the adic topological ring $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$).
  2. (iv)
    For every pair of affine open formal subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$, the $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V})$-module map
    $\mathfrak{P}[\mathfrak{V}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})}(\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V}),\mathfrak{P}[\mathfrak{U}])$
    corresponding by adjunction to the corestriction map $\mathfrak{P}[\mathfrak{V}]\longrightarrow\mathfrak{P}[\mathfrak{U}]$ is an isomorphism.
  3. (v)
    For every pair of affine open formal subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$, one has
    $\operatorname{Ext}^{1}_{\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U}){\operatorname{\mathsf{--Contra}}}}(\mathfrak{O}_{\mathfrak{X}}(\mathfrak{V}),\mathfrak{P}[\mathfrak{U}])=0.$
We will call condition (iii) is the contramoduleness axiom, while condition (iv) will be called the contraherence axiom, and condition (v) will be called the contraadjustedness axiom. Cf. the definitions of a (locally) cohereherent cosheaf of modules over a scheme in [17, Sections 2.2 and 3.1] and [29, Section 2.2].
Comparing the conditions of Lemma 2.9.3(2) and (3), one can see that condition (v) holds for a fixed affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$ and all affine open formal subschemes $\mathfrak{V}\subset\mathfrak{U}$ if and only if the contramodule $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-module $\mathfrak{P}[\mathfrak{U}]$ is contraadjusted. So our terminology is consistent.
Lemma 3.5.1.
The contramoduleness, contraherence, and contraadjustedness axioms imply the cosheaf axiom: Any contraherent copresheaf of contramodules on $\mathbf{B}$ is a cosheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules on $\mathbf{B}$.
Proof.
The assertion follows from exactness of the rightmost fragment of sequence (10) from Lemma 2.12.11 together with Remarks 3.2.1. ∎
Thus any contraherent contramodule copresheaf $\mathfrak{P}$ on $\mathbf{B}$ is, in fact, a cosheaf on $\mathbf{B}$, and we will speak of $\mathfrak{P}$ as a contraherent contramodule cosheaf (or a contraherent cosheaf of contramodules) on $\mathbf{B}$. By Lemma 3.5.1 and in view of the discussion in Section 3.2, any contraherent cosheaf of contramodules on $\mathbf{B}$ can be extended to a cosheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules on the whole formal scheme $\mathfrak{X}$ in a unique way. Moreover, according to the discussion in Section 3.3, the resulting cosheaf is a cosheaf of contramodule $\mathfrak{O}_{\mathfrak{X}}$-modules on $\mathfrak{X}$.
The notion of a contraherent cosheaf over a scheme (hence also of a contraherent cosheaf of contramodules over a formal scheme) is not local. See [17, Example 3.2.1] for a counterexample.
More precisely, in the context of a formal scheme $\mathfrak{X}$, for any contraherent cosheaf of contramodules $\mathfrak{P}$ on $\mathbf{B}$ extended to a cosheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules $\mathfrak{P}$ on the whole of $\mathfrak{X}$, both the contramoduleness axiom (iii) and the contraadjustedness axiom (v) hold for all affine open subschemes $\mathfrak{U}\subset\mathfrak{X}$. For the contramoduleness axiom, this was mentioned in the previous paragraph. For the contraadjustedness axiom, this follows from the facts that the restrictions of scalars preserve contraadjustedness (see Lemma 2.9.7), finite direct sums of contraadjusted contramodules are contraadjusted, and any quotient contramodule of a contraadjusted contramodule is contraadjusted (as mentioned in Section 2.9). It is the contraherence axiom (iv) that is not local.
3.6. Locally contraherent cosheaves of contramodules
We still keep the notation of Sections 3.4 and 3.5.
Let $\mathfrak{P}$ be a cosheaf of contramodule $\mathfrak{O}_{\mathfrak{X}}$-modules on $\mathfrak{X}$. We will say that $\mathfrak{P}$ is a $\mathbf{W}$-locally contraherent cosheaf of contramodules (or a $\mathbf{W}$-locally contraherent contramodule cosheaf) on $\mathfrak{X}$ if the contraherence axiom (iv) and the contraadjustedness axiom (v) hold for all affine open formal subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$. If this is the case, then the contraadjustedness axiom actually holds for all affine open formal subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$, as explained in Section 3.5. According to the discussion in Section 3.5, the category of $\mathbf{W}$-locally contraherent cosheaves of contramodules on $\mathfrak{X}$ is equivalent to the category of contraherent co(pre)sheaves on the topology base $\mathbf{B}$ consisting of all affine open subschemes of $\mathfrak{X}$ subordinate to $\mathbf{W}$.
A cosheaf of contramodule $\mathfrak{O}_{\mathfrak{X}}$-modules $\mathfrak{P}$ is said to be a contraherent cosheaf of contramodules (or a contraherent contramodule cosheaf) on $\mathfrak{X}$ if if the contraherence axiom (iv) and the contraadjustedness axiom (v) hold for all affine open formal subschemes $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{X}$. So a contraherent cosheaf of contramodules is the same thing as a $\mathbf{W}_{\mathfrak{X}}$-locally contraherent cosheaf of contramodules (where $\mathbf{W}_{\mathfrak{X}}=\{\mathfrak{X}\}$ is the trivial open covering of $\mathfrak{X}$).
The next proposition is the formal scheme version of the homological criterion of contraherence from [17, Lemma 3.2.2]. The definition of the homological Čech complex $C_{\text{\smaller\smaller$\scriptstyle\bullet$}}(\{\mathfrak{V}_{\alpha}\},\mathfrak{P})$ was given in Section 3.2.
Proposition 3.6.1.
Let $\mathfrak{U}$ be an affine Noetherian formal scheme with an open covering $\mathbf{W}$, and let $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{W}_{\alpha}$ be a finite affine open covering of $\mathfrak{U}$ subordinate to $\mathbf{W}$. Let $\mathfrak{P}$ be a $\mathbf{W}$-locally contraherent cosheaf of contramodules on $\mathfrak{U}$. Then the following conditions are equivalent:
  1. (1)
    $H_{1}C_{\text{\smaller\smaller$\scriptstyle\bullet$}}(\{\mathfrak{W}_{\alpha}\},\mathfrak{P})=0$;
  2. (2)
    $H_{i}C_{\text{\smaller\smaller$\scriptstyle\bullet$}}(\{\mathfrak{W}_{\alpha}\},\mathfrak{P})=0$ for all $i>0$;
  3. (3)
    $\mathfrak{P}$ is a (globally) contraherent cosheaf of contramodules on $\mathfrak{U}$.
Proof.
(2) $\Longrightarrow$ (1) Obvious.
(3) $\Longrightarrow$ (2) The coagumented homological Čech complex $C_{\text{\smaller\smaller$\scriptstyle\bullet$}}(\{\mathfrak{W}_{\alpha}\},\mathfrak{P})\longrightarrow\mathfrak{P}[\mathfrak{U}]$ is acyclic by Lemma 2.12.11 (applied to the contraadjusted contramodule $\mathfrak{P}[\mathfrak{U}]$ over the adic topological ring $R=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$).
(1) $\Longrightarrow$ (3) This is dual-analogous to (but more complicated than) the proof of Proposition 3.4.2. Let $\mathfrak{V}\subset\mathfrak{U}\subset\mathfrak{U}$ be two affine open formal subschemes. We need to check the contraherence axiom (iv) for $\mathfrak{P}$ with respect to $\mathfrak{V}\subset\mathfrak{U}$.
Notice that $\mathfrak{V}=\bigcup_{\alpha=1}^{N}(\mathfrak{V}\cap\mathfrak{W}_{\alpha})$ is an finite open covering of $\mathfrak{V}$ by affine open formal subschemes subordinate to $\mathbf{W}$. By the cosheaf axiom (12) and the assumption (1), we have a four-term right exact sequence of contramodule $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$-modules
(19) $\bigoplus\nolimits_{1\leq\alpha<\beta<\gamma\leq N}\mathfrak{P}[\mathfrak{W}_{\alpha}\cap\mathfrak{W}_{\beta}\cap\mathfrak{W}_{\gamma}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}\mathfrak{P}[\mathfrak{W}_{\alpha}\cap\mathfrak{W}_{\beta}]\\ \mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}\mathfrak{P}[\mathfrak{W}_{\alpha}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathfrak{P}[\mathfrak{U}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0.$
By the cosheaf axiom (12), we also have a right exact sequence of contramodule $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V})$-modules
(20) $\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}\mathfrak{P}[\mathfrak{V}\cap\mathfrak{W}_{\alpha}\cap\mathfrak{W}_{\beta}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}\mathfrak{P}[\mathfrak{V}\cap\mathfrak{W}_{\alpha}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathfrak{P}[\mathfrak{V}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0.$
All the terms of (19) are contraadjusted contramodule $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$-modules by Lemma 2.9.7. Since the class of contraadjusted contramodule $R$-modules is closed under quotients in $R{\operatorname{\mathsf{--Contra}}}$ for any adic topological ring $R$, we obtain a four-term exact sequence of contramodule $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$-modules
(21) $0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathfrak{Q}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}\mathfrak{P}[\mathfrak{W}_{\alpha}\cap\mathfrak{W}_{\beta}]\\ \mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}\mathfrak{P}[\mathfrak{W}_{\alpha}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathfrak{P}[\mathfrak{U}]\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0.$
with a contraadjusted contramodule $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$-module $\mathfrak{Q}$. The contramodule $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$-module $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V})$ is very flat, so applying the functor $\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}),{-})$ to the sequence (21), we obtain a four-term exact sequence of $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V})$-modules
$0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}),\mathfrak{Q})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}),\>\mathfrak{P}[\mathfrak{W}_{\alpha}\cap\mathfrak{W}_{\beta}])\\ \mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}),\mathfrak{P}[\mathfrak{W}_{\alpha}])\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}),\mathfrak{P}[\mathfrak{U}])\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0.$
So, in particular, the sequence
(22) $\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{1\leq\alpha<\beta\leq N}\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}),\>\mathfrak{P}[\mathfrak{W}_{\alpha}\cap\mathfrak{W}_{\beta}])\\ \mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\bigoplus\nolimits_{\alpha=1}^{N}\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}),\mathfrak{P}[\mathfrak{W}_{\alpha}])\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}),\mathfrak{P}[\mathfrak{U}])\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0.$
is right exact.
The corestriction maps in the cosheaf $\mathfrak{P}$ induce a natural morphism from the right exact sequence (20) to the right exact sequence (22). It remains to refer to the next Lemma 3.6.2 to the effect that the map from (20) to (22) is an isomorphism on the middle and leftmost terms. Hence the natural map $\mathfrak{P}[V]\longrightarrow\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}),\mathfrak{P}[\mathfrak{U}])$ on the rightmost terms is an isomorphism, too. ∎
Lemma 3.6.2.
Let $\mathfrak{U}$ be an affine Noetherian formal scheme with an open covering $\mathbf{W}$, let $\mathfrak{P}$ be a $\mathbf{W}$-locally contraherent cosheaf of contramodules on $\mathfrak{U}$, let $\mathfrak{V}\subset\mathfrak{U}$ be an affine open formal subscheme, and let $\mathfrak{W}\subset\mathfrak{U}$ be an affine open formal subscheme subordinate to $\mathbf{W}$. Then the corestriction map of $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})$-modules $\mathfrak{P}[\mathfrak{V}\cap\mathfrak{W}]\longrightarrow\mathfrak{P}[\mathfrak{W}]$ induces an isomorphism of $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V})$-modules $\mathfrak{P}[\mathfrak{V}\cap\mathfrak{W}]\longrightarrow\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}),\mathfrak{P}[\mathfrak{W}])$.
Proof.
This is the dual-analogous version of Lemma 3.4.3. By the contraherence axiom (iv), the restriction map $\mathfrak{P}[\mathfrak{V}\cap\mathfrak{W}]\longrightarrow\mathfrak{P}[\mathfrak{W}]$ induces an isomorphism of $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}\cap\mathfrak{W})$-modules
$\mathfrak{P}[\mathfrak{V}\cap\mathfrak{W}]\simeq\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}\cap\mathfrak{W}),\>\mathfrak{P}[\mathfrak{W}]).$
There is also an obvious isomorphism
$\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}),\mathfrak{P}[\mathfrak{W}])\simeq\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W}),\>\mathfrak{P}[\mathfrak{W}]).$
It remains to show that the map
$\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}\cap\mathfrak{W}),\mathfrak{P}[\mathfrak{W}])\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\operatorname{Hom}_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})}(\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W}),\>\mathfrak{P}[\mathfrak{W}])$
induced by the multiplication map of rings
(23) $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}\cap\mathfrak{W})$
is an isomorphism. Finally, the map (23) becomes an isomorphism after the adic completion functor $\Lambda$ is applied, while the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})$-module $\mathfrak{P}[\mathfrak{W}]$ is a (quotseparated) contramodule, and it remains to refer to Corollary 2.5.7 (whose flatness assumption is obviously satisfied for the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})$-module $F=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V})\otimes_{\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})}\mathfrak{O}_{\mathfrak{U}}(\mathfrak{W})$). ∎
We will denote the category of $\mathbf{W}$-locally contraherent cosheaves of contramodules on $\mathfrak{X}$ by $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$. The category of contraherent cosheaves of contramodules on $\mathfrak{X}$ will be denoted by $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}=\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\{\mathfrak{X}\}}$. A cosheaf of contramodule $\mathfrak{O}_{\mathfrak{X}}$-modules on $\mathfrak{X}$ is said to be locally contraherent if it is $\mathbf{W}$-locally contraherent with respect to some open covering $\mathbf{W}$ of the formal scheme $\mathfrak{X}$. The category of locally contraherent cosheaves of contramodules on $\mathfrak{X}$ will be denoted by $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}=\bigcup_{\mathbf{W}}\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$.
All infinite direct products exists in the category $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ (in particular, in $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}$). The functors of cosections $\mathfrak{P}\longmapsto\mathfrak{P}[\mathfrak{U}]$ over affine open formal subschemes $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$ preserve infinite products. In fact, the condition that $\mathfrak{U}$ is subordinate to $\mathbf{W}$ is not needed here; and the same assertion holds for all quasi-compact, quasi-separated open formal subschemes $\mathfrak{U}\subset\mathfrak{X}$.
A short sequence of $\mathbf{W}$-locally contraherent cosheaves of contramodules $0\longrightarrow\mathfrak{L}\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}\longrightarrow 0$ on $\mathfrak{X}$ is said to be (admissible) exact if the short sequence of (contramodule) $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules $0\longrightarrow\mathfrak{L}[\mathfrak{U}]\longrightarrow\mathfrak{M}[\mathfrak{U}]\longrightarrow\mathfrak{N}[\mathfrak{U}]\longrightarrow 0$ is exact for every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$. By Lemma 2.13.10(a), it suffices to check this condition for affine open subschemes $\mathfrak{U}$ belonging to any chosen affine open covering of $\mathfrak{X}$ subordinate to $\mathbf{W}$.
Hence a short sequence of contraherent cosheaves of contramodues on $\mathfrak{X}$ is exact in $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}$ if and only if it is exact in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$. More generally, let $\mathbf{W}^{\prime}$ be an open covering of $\mathfrak{X}$ subordinate to $\mathbf{W}$. Then it follows from Lemma 2.13.10(a) that a short sequence in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ is exact in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ if and only if it is exact in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}^{\prime}}$. A short sequence is said to be exact in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}$ if it is exact in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ for some open covering $\mathbf{W}$ of $\mathfrak{X}$.
Lemma 3.6.3.
(a) Let $\mathfrak{X}$ be a locally Noetherian formal scheme with an open covering $\mathbf{W}$. Then the category $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ of $\mathbf{W}$-locally contraherent cosheaves of contramodules on $\mathfrak{X}$, endowed with the class of (admissible) short exact sequences specified above, is an exact category (in the sense of Quillen [4]).
(b) Let $\mathfrak{U}$ be an affine Noetherian formal scheme. Then the functor of global cosections $\mathfrak{P}\longmapsto\mathfrak{P}[\mathfrak{U}]$ establishes an equivalence between the exact category $\mathfrak{U}{\operatorname{\mathsf{--Ctrh}}}$ of contraherent cosheaves of contramodules on $\mathfrak{U}$ and the exact category $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U}){\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ of contraadjusted contramodules over the adic Noetherian ring $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$.
Proof.
Both the assertions follow from Corollary 2.12.1 and Proposition 2.12.3. ∎
It follows from Proposition 3.6.1 that the full subcategory $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}$ is closed under extensions in the exact category $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$. More generally, the full subcategory $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ is closed under extensions in the exact category $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}^{\prime}}$.
Furthermore, it follows from Lemma 2.13.10(b) that the full subcategory $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}$ is closed under kernels of admissible epimorphisms in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$. More generally, the full subcategory $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ is closed under kernels of admissible epimorphisms in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}^{\prime}}$. So a morphism in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ is an admissible epimorphism in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ if and only if it is an admissible epimorphism in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}^{\prime}}$. The similar assersions for admissible monomorphisms are not true [17, Example 3.2.1].
3.7. Locally cotorsion locally contraherent cosheaves
A $\mathbf{W}$-locally contraherent cosheaf of contramodules $\mathfrak{P}$ on $\mathfrak{X}$ is said to be locally cotorsion if the contramodule $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-module $\mathfrak{P}[\mathfrak{U}]$ is cotorsion for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$. By Proposition 2.12.4 and Corollary 2.12.12, it suffices to check that condition for affine open subschemes $\mathfrak{U}$ belonging to any chosen affine open covering of $\mathfrak{X}$ subordinate to $\mathbf{W}$.
We will denote the full subcategory of locally cotorsion $\mathbf{W}$-locally contraherent cosheaves of contramodules on $\mathfrak{X}$ by $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}\subset\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$. The category of locally cotorsion contraherent cosheaves of contramodules on $\mathfrak{X}$ will be denoted by $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}^{\mathsf{lct}}=\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\{\mathfrak{X}\}}^{\mathsf{lct}}$. It follows from the previous paragraph that $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}^{\mathsf{lct}}=\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}\cap\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}$ and $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}=\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}\cap\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}^{\prime}}^{\mathsf{lct}}$. The category of locally cotorsion locally contraherent cosheaves of contramodules on $\mathfrak{X}$ will be denoted by $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}^{\mathsf{lct}}=\bigcup_{\mathbf{W}}\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}$.
Clearly, the full subcategory $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}$ is closed under extensions, cokernels of admissible monomorphisms, and infinite direct products in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$. So the full subcategories $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}^{\mathsf{lct}}$,  $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}$, and $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}^{\mathsf{lct}}$ inherit exact category structures from the ambient exact categories $\mathfrak{X}{\operatorname{\mathsf{--Ctrh}}}$,  $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$, and $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}$. A morphism in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}$ is an admissible monomorphism in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}$ if and only if it is an admissible monomorphism in $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$.
Lemma 3.7.1.
Let $\mathfrak{U}$ be an affine Noetherian formal scheme. Then the equivalence of exact categories $\mathfrak{U}{\operatorname{\mathsf{--Ctrh}}}\simeq\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U}){\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ from Lemma 3.6.3(b) restricts to an equivalence $\mathfrak{U}{\operatorname{\mathsf{--Ctrh}}}^{\mathsf{lct}}\simeq\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U}){\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ between the exact category of locally cotorsion contraherent cosheaves of contramodules on $\mathfrak{U}$ and the exact category of cotorsion contramodules over the adic Noetherian ring $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$.
Proof.
Follows from Proposition 2.12.4. ∎
Lemma 3.7.2.
Let $\mathfrak{X}$ be a locally Noetherian formal scheme with an open covering $\mathbf{W}$. In this context:
(a) A complex $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in the exact category of $\mathbf{W}$-locally contraherent cosheaves of contramodules $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}$ is exact if and only if, for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$, the complex of $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}[\mathfrak{U}]$ is exact. It suffices to check this condition for affine open subschemes $\mathfrak{U}\subset\mathfrak{X}$ belonging to any chosen open covering of $\mathfrak{X}$ subordinate to $\mathbf{W}$.
(a) A complex $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in the exact category of locally cotorsion $\mathbf{W}$-locally contraherent cosheaves of contramodules $\mathfrak{X}{\operatorname{\mathsf{--Lcth}}}_{\mathbf{W}}^{\mathsf{lct}}$ is exact if and only if, for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$ subordinate to $\mathbf{W}$, the complex of $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}[\mathfrak{U}]$ is exact. It suffices to check this condition for affine open subschemes $\mathfrak{U}\subset\mathfrak{X}$ belonging to any chosen open covering of $\mathfrak{X}$ subordinate to $\mathbf{W}$.
Proof.
This is similar to [17, Lemma 3.1.2]. The results of Lemmas 2.13.10(c) and 2.13.11(c) are relevant, as well as the results of Proposition 2.13.12. ∎